### v1

# Distributional Random Forests for Complex Survey Designs on Reproducing Kernel Hilbert Spaces

Abstract
We study estimation of the conditional law $P(Y|X=\mathbf{x})$ and continuous functionals $\Psi(P(Y|X=\mathbf{x}))$ when $Y\in\mathcal{Y}$ takes values in a locally compact Polish space, $X\in\mathbb{R}^{p}$, and the observations arise from a complex survey design. We propose a survey-calibrated distributional random forest (SDRF) that incorporates complex-design features via a pseudo-population bootstrap, PSU-level honesty, and a Maximum Mean Discrepancy (MMD) split criterion computed from kernel mean embeddings of Hájek-type (design-weighted) node distributions. We provide a framework for analyzing forest-style estimators under survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions. As far as we are aware, these are the first results on model-free estimation of conditional distributions under survey designs. Simulations under a stratified two-stage cluster design provide finite sample performance and demonstrate the statistical error price of ignoring the survey design. The broad applicability of SDRF is demonstrated using NHANES: We estimate the tolerance regions of the conditional joint distribution of two diabetes biomarkers, illustrating how distributional heterogeneity can support subgroup-specific risk profiling for diabetes mellitus in the U.S. population.
1 Introduction and Motivation
Experimental design is a key driver of successful statistical practice [97]. A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings [71].
Complex survey design [42] methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies [59]. In the finite-population regime [25], a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals. These estimators can exhibit satisfactory statistical properties, such as efficiency, provided that appropriate statistical calibration and inference procedures are used. This design-based framework is particularly compelling when paired with randomized interventions: probability surveys deliver external validity (representativeness of a well-defined target population), while randomized experiments deliver internal validity (causal identification under the design). Together, they support the policy-relevant generalization of causal effects [45].
Large, nationally representative health studies, such as the National Health and Nutrition Examination Survey (NHANES), which monitors health and nutrition in the U.S. population, are canonical examples of the impact of complex survey methodology in practice [68]. When design information is properly incorporated and analyzed with appropriate statistical methods, such data can be used to quantify disease burden, monitor disparities, and inform clinical and public health policy in a cost-effective way, all while ensuring valid population inference.
Gap and motivation: Machine learning and survey data.
Despite the central role of complex survey designs in nationally representative studies, predictive tools for complex surveys remain limited in two main respects. First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods [61]. Second, available procedures often presuppose low-dimensional covariates and focus on smoothing over relatively simple or very smooth function classes. By contrast, the use of modern machine-learning methods that can capture complex structures more effectively in high-dimensional settings such as neural networks [5, 98] or tree-based ensembles [26]-—within the survey context remains relatively scarce [66, 67].
From an application’s perspective, many scientific questions demand distributional targets beyond means or medians [50], including multivariate quantiles, tolerance regions [34], tail probabilities, and other functionals of multivariate conditional laws. In NHANES, for example, multivariate distribution estimates can be used to define tolerance regions that detect outlying biomarker profiles given subject-specific characteristics. Treating complex-survey observations as i.i.d. data discards design information, degrading calibration and invalidating design-based uncertainty quantification, while properly incorporating weights and design features preserves both consistency and representativeness.
Modern machine-learning methods, such as neural networks and regression-tree ensembles, are effective in many high-dimensional regimes and can capture intricate statistical associations that are difficult to model with classical nonparametric techniques. However, they are typically formulated under i.i.d. sampling and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference. A design-aware approach must propagate weights and clustering through splitting and aggregation, respect replicate-weight or bootstrap schemes that mimic the original selection, and produce distributional outputs for vector-valued $Y$, not merely conditional means. These requirements call for distributional learning procedures that integrate survey design into both the objective function and the algorithmic construction.
From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings.
Our paper. We study nonparametric estimation of multivariate conditional distributions under complex survey designs. Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design. We introduce a design-aware distributional random forest estimator $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ that combines survey-weighted tree construction and aggregation with loss functions induced by mean kernel embeddings of probability measures in a separable reproducing kernel Hilbert space. This yields a single forest-based procedure that (i) targets the full conditional law $\mathbb{P}(Y\mid X=\mathbf{x})$ and, as plug-in estimators, a broad class of component-wise and joint functionals such as conditional quantiles, cross-moments, covariance matrices and tail probabilities, and (ii) remains consistent under the joint law induced by the super–population model and the survey design.
On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) [37, 80, 89], a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the forest weighing the survey achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity. We obtain a decomposition of the MMD risk into design, algorithmic, and stochastic terms, prove that each component vanishes under mild moment conditions on inclusion and replicate weights. Our analysis quantifies the impact of the number of trees $B$ and the bootstrap or replicate-weight draws $M$ on node-level predictions and on the variance of the ensemble, showing a reduction in variance of order $O(M^{-1/2})$ while preserving the consistency of the design. We emphasize that, even in the i.i.d. setting, the theory of classical random forests is far from trivial [9, 10] and has required substantial effort to develop since Breiman’s original proposal of random forests [17]. Moreover, many questions remain open, for instance regarding finite-sample guarantees [55] or extensions to dependent data [36].
We complement the theory with an application to NHANES, where we estimate the joint conditional distribution of two biomarkers used in diabetes diagnosis based on individual-specific characteristics, providing distributional estimates that can inform a range of decisions from a healthcare and public-health perspective.
Positioning within the literature.
Our work is based on two strands. On the survey side, foundational contributions make the sampling mechanism explicit through survey weights [96, 84, 40, 77, 30, 82, 81, 39], including the Horvitz–Thompson and Hájek estimators [44, 28], as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression [51, 75, 61, 74, 70, 47]. On the machine-learning side, we exploit the statistical learning properties of random forests [17] to reduce variance estimation and even bias in many regimes via the bagging step [15, 55]. Random forests [17] and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks [31] and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks. However, existing random forest algorithms are largely developed for i.i.d. data   design-aware random forest methodology and theory for complex surveys remain underdeveloped.
From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation [91, 69]. Classical nonparametric smoothers—including local polynomials [48], Nadaraya–Watson estimators [43], and spline-based additive models—have been adapted to complex surveys [14]. However, most of this literature focuses on continuous outcomes [83, 46] and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs [6, 60, 78, 11]. Empirical-process formulations replace the uniform empirical measure with Horvitz–Thompson/Hájek-weighted measures; under suitable weight conditions, the resulting processes inherit many i.i.d. properties. Survey-aware resampling is well–developed [76]: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals. However, existing theoretical developments are not directly tailored to distributional random forests under complex designs, where the interaction between algorithmic randomness, survey weights, and distributional loss functions requires a novel theoretical survey framework. In this paper, we provide a design-aware distributional random forest for survey data, together with a corresponding theoretical framework for its design-based and model-based properties. When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Table S3.
The remainder of the paper is organized as follows. Section 2 formalizes the complex-survey setup and recalls background information on the mean kernel embeddings. Section 3 presents the algorithm and the main theoretical results. Section 4 reports simulations and comparisons with a distributional random forest that ignores the design. Section 5 analyzes NHANES. Section 6 discusses limitations and future work. All proofs of the theoretical results and a notation table (Table S3) are provided in the Supplementary Material.
2 Set-up and Background
Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i=1}^{n_{s}}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly. Here, $Y_{i}\in\mathcal{Y}$ is the outcome, $X_{i}\in\mathbb{R}^{p}$ are covariates and $w_{i}\in\mathbb{R}^{+}$ are survey weights (possibly incorporating nonresponse adjustment and calibration). Let $P^{N}_{Y\mid X=\mathbf{x}}$ denote the conditional distribution of a finite population induced by $D^{N}$. Let $P_{Y\mid X=\mathbf{x}}$ denote the super-population conditional law. Consider a fixed covariate value $\mathbf{x}\in\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and it is continuous at the target. Our objective is to construct an estimator $\widehat{\Psi}_{N}(\mathbf{x})$ that is design-consistent for the finite-population target $\Psi(P^{N}_{Y\mid X=\mathbf{x}})$ and model-consistent for the super-population target $\Psi(P_{Y\mid X=\mathbf{x}})$. We will work on the joint probability space $(\Omega\times\mathcal{S}^{N},\,\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega\times\mathcal{S}^{N}})$. Specifically, denote $P_{\mathcal{S}^{N}|\omega}$ the measure in the design space, we want that for any $\varepsilon>0$,
$P_{S^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
$P_{\Omega\times\mathcal{S}^{N}}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{ (model consistency). }$
2.1 Survey Design
We follow the finite population framework considered by [13, 41, 42]: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$. Let $U^{N}:=\{1,\dots,N\}$, the index set for each $\mathcal{U}^{N}$ and $\mathcal{S}^{N}:=\{S:S\subset U^{N}\}$ the collection of subsets of $U^{N}$. Let $\sigma(\mathcal{S}^{N})$ denote the $\sigma$-algebra generated by $\mathcal{S}^{N}$. We view the finite population $\mathcal{U}^{N}$ as a realization of i.i.d. random vectors $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ in a common probability space $(\Omega,\mathcal{A},P)$, where $Z_{i}\in\mathcal{Z}$ denotes the auxiliary (design) information available in the frame. Let $D^{N}:=\{(X_{i},Y_{i})\}_{i=1}^{N}$ denote the finite-population data and write $Z^{N}:=(Z_{1},\dots,Z_{N})$. For any region ${A}(\mathbf{x})\subset\mathcal{X}$ indexed by some fixed $\mathbf{x}\in\mathcal{X}$, denote the empirical conditional measure induced by ${D}^{N}$ as
${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}I(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
$N(\mathbf{x})=\sum_{i\in[N]}I(X_{i}\in A(\mathbf{x}))$. A sampling scheme can be formalized as a transition function $\bm{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$. From $\bm{p}$, fix any $\omega\in\Omega$, we define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\bm{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$. We can then uniquely define a product probability measure in the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$ as
$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\bm{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
Thus, a random sample $S\sim P_{\mathcal{S}^{N}}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$. Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\bm{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$. Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$, the size of the random sample $S$ that is a random variable. We do not have access to the entire finite population of size $N$, but we work with a sample of size $n_{s}$. This leads to estimators that use weights to construct consistent estimators ([56]). For example, the Hájek (HJ) estimator for the true conditional probability ${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})$ in the design space is
$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x})).$
The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts the form Hájek, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$. We assume the following with respect to the design.
  • On study design
  • (D1)
    Conditionally non-informative design: $\xi^{N}\mathrel{\perp\!\!\!\perp}Y^{N}\mid(X^{N},Z^{N})$.
  • (D2)
    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, with some $\kappa>4$.
  • (D3)
    Random Sampling: There exist constants $0<\underline{\lambda}\leq\bar{\lambda}<\infty$ such that for any $N\geq 1$, $\sup_{i\in[N]}\pi_{i}\leq\bar{\lambda}n/N$, $\inf_{i\in[N]}\pi_{i}\geq\underline{\lambda}n/N$, where $n=\mathbb{E}_{P_{{\mathcal{S}_{N}}\mid\omega}}(n_{s}(\omega))$ the expected sample size.
  • (D4)
    Controlled second order: For any $N\geq 1$, $\max_{1\leq i\leq j\leq N}|\pi_{ij}-\pi_{i}\pi_{j}|\leq C_{1}/n_{s,N}$ a.s.
Remark.
Assumptions (D1)–(D4), or minor variants thereof, are standard regularity conditions in the survey-sampling literature. (D2)–(D4) imply the design-based law of large numbers and central limit theorem statements, summarized in Lemma 2.1. These probabilistic bounds will be invoked repeatedly to establish the consistency of $\widehat{\Psi}_{N}(\mathbf{x})$. (D1) allows us to separate the super–population randomness from the design randomness and to treat the design expectations conditionally on $(X^{N},Z^{N})$. (D2) imposes a stable sampling fraction and rules out designs with a decreasing sample size. (D3) assumes that the inclusion probabilities of first order are uniformly of order $n/N$ and bounded away from 0 and $1$. This prevents extreme survey weights. (D4) is common [12, 11, 19] and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.
Lemma 2.1.
Under (D2)-(D4), we have that w.r.t the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D1)–(D4). Two–stage designs that satisfy a counter-part of (D1)–(D4) at each of the two stages satisfy also (D1)–(D4) at the subject level (Lemma A.2, Supplementary Material).
2.2 Embedding of Measures
Kernel mean embeddings (KMEs) [38, 85] provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) [1]. Since these representations are defined purely through expectations of the kernel evaluations, all operations on distributions are reduced to linear operations in the RKHS. As a result, questions about probability measures–such as whether two random variables share the same distribution–can be reformulated as questions about distances between vectors and efficiently evaluated via the kernel trick. The distance between two embedded distributions in the RKHS is called the maximum mean discrepancy (MMD). It is a fully nonparametric way to compare distributions, and for characteristic kernels it is equal to zero only when the two distributions are the same. This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., [86, 87, 37, 80, 33, 67].
From the perspective of constructing random forests for distributional targets, incorporating KMEs into the algorithm offers several practical advantages: (i) It provides a distribution-level target, enabling random forests to discriminate conditional laws via an MMD-based splitting criterion. (ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size [32]. (iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data [62]). As a result, modifications to a standard random forest algorithm are minimal, even for outcomes in abstract spaces. (iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to both finite-population and super-population conditional laws under standard regularity conditions, allowing us to analyze a broad class of sampling designs within a single metric framework.
We now introduce the corresponding concepts and definitions from a mathematical perspective.
Let $\mathbf{k}:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ be a bounded positive definite kernel with RKHS $(\mathcal{H},\langle\cdot,\cdot\rangle_{\mathcal{H}})$ and the associated canonical feature map $\Phi_{\mathbf{k}}(y):=\mathbf{k}(y,\cdot)\in\mathcal{H}$. For any $P\in\mathcal{M}_{b}(\mathcal{Y})$ in the space of finite signed Radon measures, define the kernel mean embedding as the mapping
$\mu_{\mathbf{k}}:\mathcal{M}_{b}(\mathcal{Y})\mapsto\mathcal{H},\hskip 5.69046pt{P}\mapsto\int_{\mathcal{Y}}\mathbf{k}(y,\cdot)d{P}(y)\in\mathcal{H}.$
so that by the reproducing property, for any $f\in\mathcal{H}$, $P\in\mathcal{M}_{b}(\mathcal{Y})$, $\mu_{\mathbf{k}}(P)=\int_{\mathcal{Y}}\Phi_{\mathbf{k}}(y)dP(y)=\mathbb{E}_{Y\sim P}[\Phi_{\mathbf{k}}(Y)]\in\mathcal{H},$ a mean element in the RKHS. This yields a natural distance in the RKHS as $d_{\mathbf{k}}(P,\;Q)=\|\mu_{\mathbf{k}}(P)-\mu_{\mathbf{k}}(Q)\|_{\mathcal{H}},\;P,\;Q\in\mathcal{M}_{b}(\mathcal{Y})$. We call $\mathbf{k}$ characteristic if $\mu_{\mathbf{k}}$ is injective. In that case, $d_{\mathbf{k}}(P,\;Q)=0$ if and only if $P=Q$, so $d_{\mathbf{k}}$ is a valid statistical distance for the unique identification of the probability measures.
For our consistency arguments, we require, in addition, that $d_{\mathbf{k}}$ metrizes weak convergence on $\mathcal{M}_{b}(\mathcal{Y})$. That is, for any sequence $P_{1},\dots\in\mathcal{M}_{b}(\mathcal{Y})$, $P_{N}\rightsquigarrow P\iff d_{\mathbf{k}}(P_{N},\;P)\to 0$, as $N\to\infty$. We call kernels that satisfy this property universal (initially proposed in [88]). This property concerns the approximation ability of the KME sample estimate to the true target as $N\to\infty$. We place the following assumptions:
  • On Space $\mathcal{Y}$ and Kernel $\mathbf{k}$
  • (S1)
    $\mathcal{Y}$ is a locally compact Polish space.
  • (K1)
    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of [86]).
  • (K2)
    $\mathbf{k}$ is $c_{0}$-universal. Meaning that the RKHS $\mathcal{H}$ induced by $\bm{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t. the uniform norm (Definition 2, Theorem 1 in [20]).
  • (K3)
    (Optional) When $\mathcal{Y}$ carries a compatible abelian topological group structure so that translations are defined and continuous, $\mathbf{k}$ is translation-invariant.
Remark.
Assumptions (S1)–(K2) ensure that $\mu_{\mathbf{k}}(P)$ is well-defined for $P\in\mathcal{M}_{b}(\mathcal{Y})$ and that $d_{\mathbf{k}}$ is both discriminative (characteristic) and topologically compatible with weak convergence (metrizing). Assumption (K3) is not needed for the statistical theory, but when satisfied, admits a simple finite-dimensional approximation of the MMD through Fourier-feature representation for fast computation. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ ([38]); More details are provided in Section A.10 of the Supplementary Material.
Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding; see Section A.10.
3 SDRF Algorithm and Theoretical Result
Random forests are machine learning methods that combine a CART-style [27] tree ensemble with a bagging step [15], and can also be viewed as kernel-type estimators (see, e.g. [54, 72]). Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems [17], and several works have documented their strong empirical performance [26, 31]. More recently, conditional distributional versions of random forests based on kernel mean embeddings have been proposed for i.i.d. data [21]. Our proposed method, SDRF, constructs a kernel estimator $\widehat{\omega}_{\mathbf{x}}(\cdot)$ that directly targets conditional distributions and appropriately incorporates the survey design. Given the estimated neighborhood and weights (which together define $\widehat{\omega}_{\mathbf{x}}(\cdot)$), we can directly construct an estimator of the conditional distribution $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$, and hence a plug-in estimator for $\Psi(\mathbf{x})$, namely $\Psi(\mathbf{x})$: $\Psi(\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x}))$.
We introduce additional notation to help clarify SDRF. Denote $B$ the number of trees, $T_{b}$ the bb–th tree fitted on re-sampled data $D^{*}_{b}=\{(X_{i},Y_{i},1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$, where $n^{*}_{b,i}\geq 0$ are resampling multipliers. Let $q\in(0,1)$ denote the PSU-level selection probability for inclusion in the splitting subsample $D^{*}_{b,\mathrm{split}}$, and let $\mathcal{I}(D^{*}_{b,\mathrm{split}})\subseteq[n_{s}]$ be the set of subject indices contained in $D^{*}_{b,\mathrm{split}}$. Denote $L(\cdot)$ the oracle partition estimated by SDRF that uses the full finite population. This reduces to fitting an RF on i.i.d. data. Denote $L^{*}(\cdot)$ the partition estimated using training data $D^{*}_{train}$ that involves re-sampled weights $n^{*}$. $L(\cdot)$ and $L^{*}(\cdot)$ map any given $\mathbf{x}$ to a region in $\mathcal{X}$. The complete SDRF implementations details are summarized in Algorithm 1. In the following subsections, we discuss in further detail key components, especially those that concern adaptations to design, and introduce their associated theoretical properties.
Algorithm 1 Survey Distributional Random Forest (SDRF)
[…]
0: Random forest ensemble $\mathcal{T}=\{T_{1},\dots,T_{B}\}$
[…]
2:for $b=1,\dots,B$ do
3:  Draw resample $D^{*}_{b}=\{(X_{i},Y_{i},1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$ respecting the design.
4:  Draw $\forall k(i)$, $I(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
5:  Split at the PSU level $D^{*}_{b}=D^{*}_{b,split}\bigcup D^{*}_{b,est}$.
6:  Train decision tree $T_{b}$ on $D^{*}_{b,split}=\{(X_{i},Y_{i},\;1/\pi_{i}\times n^{*}_{b,i}\times 1/q)\}$ by:
7:  Randomly select mtrymtry features.
8:  Choose the split that maximizes $M^{*}(\mathrm{split};D^{*}_{\mathrm{split}})=\frac{N^{*}_{L}N^{*}_{R}}{N^{*2}_{Pa}}\|\mu_{\mathbf{k}}(\hat{P}^{*}_{L})-\mu_{\mathbf{k}}(\hat{P}^{*}_{R})\|^{2}_{\mathcal{H}}$ until reaching maximum depth, minimum node size, or maximum weight ratio.
9:  $\mathcal{T}\leftarrow\mathcal{T}\;\bigcup\;\{T_{b}\}$.
10:end for
11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
12: For $\Psi(\cdot)$, obtain plug-in estimator $\Psi(\widehat{P}^{N}_{Y\mid X=\mathbf{x}})=\Psi(\sum_{i=1}^{n_{s}}\widehat{\omega}_{i}(\mathbf{x})\;\mathbf{k}(Y_{i},\,\cdot))$.
3.1 Bagging and Honesty
Resampling with Design
Bagging is a key ingredient of random forests [15], but its study is primarily conducted in an i.i.d. setting. In the i.i.d. case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators [16, 18].
Under a survey design, however, Efron’s classical i.i.d. bootstrap [29] can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
We adopt a design–aware bootstrap (BS) method in line with [95] as the default resampling method for SDRF. It generates multipliers $n^{*}$ that determine the additional up– or down– weighting applied to the sampled distribution, and satisfy properties (R1)–(R3) as detailed in Lemma 3.1.
Lemma 3.1.
For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section A.5.1 of the Supplementary Material in the spirit of [95] (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
  1. On resample multipliers $n^{*}$
  2. (R1)
    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains known cluster and strata structure. Each stratum is re–sampled separately, and if the design has two stages, the second stage resampling multipliers are independent across distinct PSUs conditional on first–stage resamples.
  3. (R2)
    $\mathbb{E}(n_{i}^{*}\mid\xi,Z^{N})=1\text{, and }\sup_{i}\mathbb{E}\bigl|n_{i}^{*}\bigr|^{2+\epsilon}<\infty\text{ for some }\epsilon>0.$
  4. (R3)
    $\exists C<\infty$ a constant such that $\sum_{i\neq j}\bigl|Cov(n_{i}^{*},n_{j}^{*}\mid\xi,Z^{N})\bigr|\;\leq\;C\sum_{i}\operatorname{Var}(n_{i}^{*}\mid\xi,Z^{N})\,\,a.s.$ for all $N$.
Remark.
(R1) ensures that the multipliers are generated in a design-adapted and symmetric way and prevents the resampling step from introducing extraneous dependence on unobserved design features. (R2) enforces centering at the obtained sample and stable fluctuations, which are common for BS multipliers. (R3) is a weak dependence requirement across units. It is convenient for deriving concentration bounds while allowing for the existence of dependency.
The method by [95] belongs to the class of pseudo-population-based resampling methods. To summarize: For each re-sample, we first use design information such as $\{w_{i}\}_{i\in[n_{s}]}$ to re-construct via the multinomial distribution a pseudo finite population, with expected size equal to $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$; then we re-apply the sampling design $\mathbf{p}$ on the re-constructed finite population to obtain re-samples, thereby respecting the design structure. We refer interested readers to [63, 23] for a review of common resampling methods under design.
Honesty with Design
For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ ([3, 93, 4]). With i.i.d. data, any disjoint partition of the observations yields honesty by construction. However, data from design carries possible dependency across subjects, so separation must be made at the PSU level to ensure the maximum level of independence conditioning on $\xi$. This motivates the introduction of $q\in(0,\,1)$. Alternatively, if one has access to data obtained from a continuous survey under the same design and is willing to assume that $P_{Y|X}$ stays unchanged between two consecutive time points, it is also possible to directly use one sample as $D_{est}$ and the other as $D_{split}$.
In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, [49]) decreases as depth increases. As we discuss later, the performance of the MMD–based split criterion hinges on the effective sample size of $D_{split}$. In this case, one might consider spending the PSUs gradually over the tree growing process. See Algorithm 2, Supplementary Material for details.
3.2 Splitting Rule
The splitting criterion of an RF algorithm guides the formation of tree structure by determining how the feature space should be partitioned into increasingly homogeneous child nodes. Consider any node where a split is possible (has not met the stopping criteria). Denote $P_{a}\subseteq\mathcal{X}$ the feature space of that node. A candidate split can be parametrized as $\theta=(j,\;t)\in\Pi$, $j=1,\dots,p$, $t\in Range(X_{j})\subseteq[a_{j},\;b_{j}]$. This partitions $P_{a}$ into child regions $C_{L}(\theta):=\{x\in P_{a}:x_{j}\leq t\}$ and $C_{R}(\theta):=\{x\in P_{a}:\;x_{j}>t\}$. With RF considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via Hájek–type estimator. We find the split by maximizing the MMD between the KMEs of the distribution estimates between the two child regions. Formally, $\hat{\theta}=\arg\max_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)$, where
$\begin{split}M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},\end{split}$ (1)
with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}I(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$. When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if exists. We first provide a theorem justifying the choice of this split criterion.
Theorem 3.2 (Local design/model consistency of the MMD split).
Fix a parent region $P_{a}\subset\mathcal{X}$ at depth $d$ and consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which defines child regions $C_{L}(\theta)$ and $C_{R}(\theta)$. Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score in (1), and let $M^{N}(\theta)$ be its finite-population analog. Suppose (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3. Suppose further $\mathcal{X}\subset\mathbb{R}^{p}$ is bounded. Then, conditional on the realized finite population,
$\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M^{N}(\theta)\bigr|\xrightarrow{p_{d}}0.$
If, in addition, the finite population is an i.i.d. realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}$ and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta_{0}$ under the super-population model).
Remark.
Theorem 3.2 provides a local result: for a fixed parent node $P_{a}$, the empirical MMD criterion $M^{*}_{n_{s}}(\theta)$ uniformly approximates its finite-population oracle counterpart $M^{N}(\theta)$ and also the super-population score $M(\theta)$. The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique ([90]). The condition $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$ makes explicit that random feature subsampling (via mtrymtry) does not distort the split choice asymptotically, provided the candidate set becomes sufficiently dense around $\theta_{0}$ in expectation. Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to that if using the full finite population, which is in particular assumption (B2) of Theorem 3.4 required for consistency in MMD metric.
One might wonder that, since resamples are determined by non–negative multipliers $n^{*}$, if one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample. This can help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees. Proposition 3.3 below provides a detailed characterization of its consequence.
Proposition 3.3 (Effect of averaging $M$ resampling draws on split-score stability).
Fix a node $P_{a}$ and a candidate split $\theta$ that induces children $C_{L},C_{R}$ with $N_{P_{a}}:=|\{i\leq N:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\leq N:X_{i}\in C_{\ell}\}|$. Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M_{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ its design-weighted sample analogue, and $M_{n_{s}}^{*(M)}(\theta)$ the resampled score computed using averaged multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$ (equivalently $\bar{\delta}_{i}:=\bar{n}_{i}^{*}-1$). Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$, which are satisfied by the methods described in Section A.5.1, then
$\bigl|M_{n_{s}}^{*(M)}(\theta)-M_{N}(\theta)\bigr|=O_{p}\!\left(N_{P_{a}}^{-1/2}\right)\;+\;O_{p}\!\left(\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\right),$
where $n_{\mathrm{eff},\ell}$ is the design-weighted effective sample size ([49]) in child $\ell$ induced by $\theta$.
Remark.
Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M_{N}(\theta)|=O_{p}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\big)$.
The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\\ n_{\mathrm{eff},\ell}^{-3/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains. At deep nodes (small $n_{\mathrm{eff},\ell}$), this threshold is reached quickly: for instance, if both children have $n_{\mathrm{eff},\ell}\approx 100$ and $N_{P_{a}}\approx 200$, then $\tilde{M}\approx 8$ suffices. Beyond this point, split stability is governed by the effective sample sizes—controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions qq—rather than by the number of resamples.
Moreover, the marginal gain from increasing $M$ deteriorates rapidly: doubling from $M$ to $2M$ reduces the resampling component by only $2^{-1/2}\approx 0.71$, and more generally $(1+\Delta/M)^{-1/2}\approx 1-\Delta/(2M)$ when $\Delta\ll M$. Thus, in typical regimes where $M$ moderately exceeds $\tilde{M}$, averaging offers limited benefit and computational resources are better invested in increasing $n_{\mathrm{eff},\ell}$.
3.3 Prediction and Aggregation
We now establish the main result of the paper, namely the consistency of the proposed SDRF plug-in estimator $\widehat{\Psi}_{N}(\mathbf{x})$. The argument proceeds in two steps. First, Theorem 3.4 shows that, under the additional algorithmic conditions (A1)–(A3), (B1)–(B2) detailed below, the MMD between the SDRF estimate of the conditional distribution and the corresponding target conditional distribution converges to zero in $P_{\mathcal{S}^{N}\mid\omega}$-probability. Second, Corollary 3.5 shows that this distributional convergence implies the desired consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ for $\Psi(\mathbf{x})$ in both the model- and design-based senses. The implication follows because SDRF is constructed to directly separate conditional distributions, and because $\Psi$ is continuous as a mapping on $\mathcal{M}_{b}(\mathcal{Y})$.
Theorem 3.4.
Suppose (S1) for the space $\mathcal{Y}$, (D1)–(D4) for the survey design $\mathbf{p}$, (R1)–(R3) for the resample multiplier $n^{*}$, and (K1)–(K3) for the reproducing kernel $\mathbf{k}$. Suppose for the true KME of the conditional measure in the model–space, $\exists L_{lip}>0$ such that for any $\mathbf{x},\mathbf{x}^{\prime}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, $\left\|\mathbb{E}_{P}[\mathbf{k}(Y,\cdot)\mid X=\mathbf{x}]-\mathbb{E}_{P}[\mathbf{k}(Y,\cdot)\mid X=\mathbf{x}^{\prime}]\right\|_{\mathcal{H}}\leq L_{lip}\|\mathbf{x}-\mathbf{x}^{\prime}\|$. Suppose also the following:
  • Algorithm regularization
  • (A1)
    Shape regularity: For a.s. $L(\mathbf{x})$ for any $\mathbf{x}$ in the sampled covariate space satisfying $diam(L(\mathbf{x}))<k_{N}$,  $\exists\;0<c_{1}\leq 1$ s.t. $\bar{B}_{r_{N}}(\mathbf{x})\subseteq L(\mathbf{x})$, $r_{N}=c_{1}k_{N}$.
  • (A2)
    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ the ball centered at $\mathbf{x}$ with radius $r$.
  • (A3)
    For any $L^{*}$ found by $D_{split}$, when populating the values with $D_{est}$, all leaves receive at least $n_{L^{*},min}$ subjects. In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}I(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
  • Algorithmic behavior
  • (B1)
    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ yet $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $iam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2}), dd$ some metric on $\mathcal{X}^{p}$.
  • (B2)
    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(x)\,\Delta\,L(x)\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the estimate from SDRF. Then $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0$ as $N,B\to\infty$.
Remark.
Regularization–type conditions (A1)–(A3) can be controlled via adaptive stopping criteria. (A1) imposes a geometric control on the shape of $L$ so that they do not be stretched. (A2) requires a local mass control for the local empirical density to be well-behaved. (A3) ensures the local density estimate stays bounded. (B1)-(B3) formalize how the partitions produced by the algorithm behave as the finite population size grows. (B1) is a shrinking diameter assumption that mirrors the adaptive concentration result proved for i.i.d. forests ([94]). (B2) requires that the thickness of the error band be negligible relative to the leaf’s own diameter. In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated if using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate. Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(x)\triangle L(x)$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(x))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively. Among all the assumptions, (B2) is the hardest to directly control. Thus we provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
Corollary 3.5.
Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ i.i.d. realizations from $P_{\Omega}(\cdot)$. Suppose $\mathbf{k}$ satisfying (K1)–(K3), and $\mathcal{Y}$ locally compact Polish (satisfying (S1)). Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ an estimate for it. Denote $P_{\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the super population space. Then for any specified $\mathbf{x}\in\mathcal{X}$ and $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),\;d_{\mathbf{k}})\to(\mathcal{V},d)$ a measurable functional continuous at $P^{N}(Y|X=\mathbf{x})$ for a.e. $\omega\in\Omega$, if   $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0,\text{ then }$
$\Psi(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))$ $\to_{p_{d}}\Psi({P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\;\;\text{ (design consistency) },\text{ and }$
[…]
Remark.
Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{S^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law. Many standard summaries satisfy this requirement under mild regularity. For example, when $\mathcal{Y}=\mathbb{R}^{d}$, linear functionals such as conditional means $\Psi(P)=\int y\,dP(y)$ and, more generally, $\Psi_{g}(P)=\int g(y)\,dP(y)$ for $g\in C_{b}(\mathcal{Y})$ are continuous under weak convergence. For distributional functionals, $\Psi_{t}(P)=F_{P}(t)$ and $\Psi_{t}(P)=P(Y>t)$ are continuous at $P$ whenever $P(\{t\})=0$. The quantile functional $\Psi_{q}(P)=F_{P}^{-1}(q)$ is continuous at $P$ whenever the $q$-quantile is unique. Equivalently, that $F_{P}$ has no flat segment at level $q$. The kernel mean element itself and its norm $\Psi(P)=\|\mu_{\mathbf{k}}(P)\|_{\mathcal{H}}$, and the MMD between the true conditional $P$ to a reference distribution $Q$, $d_{\mathbf{k}}(P,Q)$, are also examples.
3.4 Implementation Considerations
We collect practical implementation considerations for SDRF, consolidating several points noted earlier.
  • For the re–sampling, ([22]) showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ pseudo population “and” the re-sample.
  • RBF–type kernel is a common class of kernels that is translation–invariant. Median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice ([35, 64]).
  • Hyperparameters governing the stopping criteria are introduced to help ensure regularity-type conditions (A1)–(A3). For finite samples, it may be preferable to use a moderate number of shallow trees to keep local sample weights well-controlled. Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section A.5.1).
  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step [92] in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d.; described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies. The algorithm is flexible in the choice of $\Psi$, the space of $Y$, as well as the type of design.
4 Simulations
4.1 Data Generation
To assess the finite-sample performance of the proposed survey-calibrated distributional random forest (SDRF), we conduct a simulation study designed to mimic the NHANES case study that follows a complex multistage survey design. Data are generated under a stratified cluster sampling scheme with probability-proportional-to-size (PPS) selection of primary sampling units (PSUs), followed by simple random sampling without replacement (SRSWOR) of secondary units within selected PSUs.
Our goals are twofold:
  1. 1.
    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF [21], an existing distributional random forest that does not incorporate survey weights in the estimation process.
  2. 2.
    Study the effect of the number of trees, the sample size and the number of PSUs per stratum $N_{\mathrm{psu},h}$, on the finite-sample performance.
We begin by defining the generative model in the super–population framework. For each unit $i=1,\dots,N$, let
$Y_{i}=(Y_{i1},Y_{i2})^{\top}\in\mathbb{R}^{2},\qquad X_{i}=(X_{i1},\dots,X_{ip})^{\top}\in\mathbb{R}^{p},\qquad Z_{i}\in\mathbb{R},$
where $Y_{i}$ is the outcome vector, $X_{i}$ the predictor vector and $Z_{i}$ a design variable that influences sampling but is not observed at the unit level in the analysis.
The finite population $\{(Y_{i},X_{i},Z_{i}):i=1,\dots,N\}$ is generated as follows: For each $i$,
$Z_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1),\quad X_{i1}\overset{\mathrm{iid}}{\sim}\text{Uniform}(0,1),$
$B_{i}\overset{\mathrm{iid}}{\sim}\operatorname{Bernoulli}\!\left(\tfrac{1}{2}\right),\quad X_{i2}=2B_{i}-1,$
$X_{ij}\overset{\mathrm{iid}}{\sim}\mathcal{N}(0,1),\quad j=3,\dots,p,$
where all variables in the construction above are mutually independent across both units $i$ and indices $j$, unless otherwise stated. We then define the conditional mean of $Y_{i}$ given $(X_{i},Z_{i})$ by
$Y_{i}\;=\;\mu(X_{i},Z_{i})+\Sigma^{1/2}\,\varepsilon_{i},\qquad\varepsilon_{i}\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mathbf{0},I_{2}),\text{ where }$
$\mu(X_{i},Z_{i})=\begin{pmatrix}\mu_{1}(X_{i},Z_{i})\\[2.0pt] \mu_{2}(X_{i},Z_{i})\end{pmatrix},\quad\Sigma=\begin{pmatrix}1&0.3\\[2.0pt] 0.3&1\end{pmatrix},\text{ and }$
$\mu_{1}(X_{i},Z_{i})=1.5\,X_{i2}+(2+50Z_{i})\,X_{i1},\qquad\mu_{2}(X_{i},Z_{i})=1.5\,X_{i2}-Z_{i}\,(3X_{i2}+1).$
$\{\varepsilon_{i}\}^{N}_{i=1}$ are generated independently of $\{(X_{i},Z_{i})\}$. Equivalently, $Y_{i}\mid(X_{i},Z_{i})\sim\mathcal{N}_{2}\bigl(\mu(X_{i},Z_{i}),\,\Sigma\bigr)$ i.i.d. across i.
We now describe the sampling design. Let $H\geq 2$ be the number of strata, and introduce a stratum indicator
$\mathrm{Strat}_{i}\in\{1,\dots,H\},\qquad i=1,\dots,N.$
Each unit is assigned independently to a stratum with equal probability:
$\bigl(\mathds{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathds{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
For each stratum $h\in\{1,\dots,H\}$, let
$U_{h}=\{i:\mathrm{Strat}_{i}=h\},\qquad N_{h}=|U_{h}|$
denote the set and size of units in stratum $h$. Within stratum $h$, define the membership map $g_{h}:U_{h}\to\{1,\ldots,N_{h}^{\mathrm{PSU}}\}$ by ordering units in $U_{h}$ by $X_{i1}$ (and breaking ties arbitrarily) and grouping the ordered list into $N_{h}^{\mathrm{PSU}}$ consecutive bins of approximately equal size. Write $g_{h}(i)=j$ if unit $i$ is assigned to PSU $j$, and define
$\mathrm{PSU}_{hj}:=\{i\in U_{h}:g_{h}(i)=j\},\qquad j=1,\ldots,N_{h}^{\mathrm{PSU}}.$
We define measure-of-size as a function of $Z$: For each PSU, $\mathrm{PSU}_{hj}$,
$M_{hj}\;=\;8\mathds{1}(Z_{hj}>0)+2\mathds{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.
First-stage sampling selects PSUs within each stratum $h$ with probabilities proportional to size. Fix an expected number $n_{h}^{(1)}$ of PSUs to be selected in stratum $h$, and define the inclusion probabilities of the first-stage
$\pi^{(1)}_{hj}\;=\;\min\!\Bigl(1,\;n^{(1)}_{h}\,\frac{M_{hj}}{\sum_{j^{\prime}=1}^{N_{h}^{\mathrm{PSU}}}M_{hj^{\prime}}}\Bigr),\qquad j=1,\dots,N_{h}^{\mathrm{PSU}}.$
We then select PSUs by unequal-probability systematic sampling,
$\{\mathrm{PSU}_{hj}:\xi^{(1)}_{hj}=1\}\sim\operatorname{UP\text{-}systematic}\!\bigl(\{\pi^{(1)}_{hj}\}_{j=1}^{N_{h}^{\mathrm{PSU}}}\bigr),$
where $\xi^{(1)}_{hj}\in\{0,1\}$ is the indicator that PSU $j$ in stratum $h$ is selected at the first stage.
In the second stage, conditional on $\xi^{(1)}_{hj}=1$, we sample units within PSU $\mathrm{PSU}_{hj}$ by SRSWOR. Specifically, we select
$n^{(2)}_{hj}\;=\;\lceil 0.3\,N_{hj}\rceil$
units without replacement in each selected PSU. Thus, for any unit $i\in\mathrm{PSU}_{hj}$,
$\pi^{(2)}_{hji}\;=\;\Pr(\xi^{(2)}_{hji}=1\mid\xi^{(1)}_{hj}=1)\;=\;\frac{n^{(2)}_{hj}}{N_{hj}}\;=\;0.3,$
where $\xi^{(2)}_{hji}\in\{0,1\}$ is the second-stage selection indicator.
Under this two-stage design, the overall unit inclusion indicator is $\xi_{hji}=\xi^{(1)}_{hj}\,\xi^{(2)}_{hji}$, with unit inclusion probability $\pi_{hji}=\Pr(\xi_{hji}=1)=\pi^{(1)}_{hj}\,\pi^{(2)}_{hji}$, and corresponding survey weights $w_{hji}=\frac{1}{\pi_{hji}}.$ This completes the generation of analysis dataset $D=\{(X_{i},Y_{i},w_{hji})\}_{\{i:\xi_{hji}=1\}}$. The sampling scheme considered here satisfies the conditional non-informativeness requirement (D1).
4.2 Computation Details
We run the SDRF and DRF ([21]) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$. For each of the scenario, 200 seeds were run. We let $p=3$ to align with the case study (Section 5). As $N$ increases, we also increase the total number of strata $H$ and the number of PSUs per stratum $N_{\mathrm{psu},h}$. The expected sample size per stratum (and hence the overall sample size) increases with $N$. The sampling mechanism (PPS at stage 1 and SRSWOR at stage 2 with fixed rates) is kept fixed in the values of $N$, which is consistent with the asymptotics of the increase in the population survey.
For analytical evaluation, we report (a) the maximum mean discrepancy (MMD) between the estimated kernel mean embedding and the population kernel mean embedding for the conditional distribution on a hold-out set, and (b) the accuracy in the estimation for the functional
$\Psi(\mathbf{x}=(x_{1},\dots,x_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
in terms of mean squared error (MSE) of $\widehat{\Psi}(\mathbf{x}_{1})$.
Unlike the training data, which are the complex-sample draws produced by the design, the testing data are generated directly from the super-population model; thus no design information (weights or $Z$) is required to evaluate $\Psi(\mathbf{x}_{1})$ in the test set. We focus on $\Psi(\mathbf{x}_{1})$ to enable visual inspection of pointwise performance across the range of $X_{1}$.
For both random-forest estimators, we use a Gaussian kernel on the outcome space with bandwidth set by the median of pairwise Euclidean distances between the observed $Y$’s. The maximum tree depth is fixed at $8$. The minimum node size is set to $\max\!\left\{\,20,\;\left\lceil\sqrt{n_{s}}\,\right\rceil\right\}$, where $n$ denotes the size of the training-sample. We cap effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights. Although additional hyperparameter tuning could further improve performance, we keep the settings minimal because the untuned configurations already reveal the main trends across $N$ and $B$.
We do not compare against alternative survey-calibrated distributional methods beyond SDRF because, to our knowledge, such methods are not currently available in the survey-sampling literature. (The unweighted DRF serves as the i.i.d. baseline and is a logical competitor). All code is implemented under R, version 4.4.2 [73].
4.3 Results
Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P(Y\mid X)$ and its estimate $\widehat{P}(Y\mid X)$ from SDRF. The mean MMD decreases as the finite population size $N$ increases, while increasing the number of trees $B$ primarily reduces the SD (stabilizing forest performance) with comparatively smaller improvements in the mean and diminishing returns beyond moderate $B$. Overall, these finite-sample patterns are consistent with the convergence behavior asserted in Theorem 3.4.
Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}(Y\mid X),\;P^{N}(Y\mid X))$, where $\widehat{P}(Y\mid X)$ is the distribution estimate from SDRF, across finite population size $N$ and number of trees $B$.
NN $B=10$ $B=30$ $B=70$ $B=200$
4,000 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
5,000 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
15,000 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
22,500 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for Survey-DRF (SDRF) and benchmark DRF algorithms as $N$ and the number of trees grow.
NN $N_{\text{PSU},h}$ $\bar{n}_{s}$ Algorithm Number of Trees $B$
10 30 70 200
4,000 80 320 DRF 7.80 7.49 7.32 7.33
SDRF 4.11 3.44 3.29 3.27
5,000 100 400 DRF 8.04 7.55 7.40 7.33
SDRF 3.69 3.16 2.95 2.96
15,000 200 1,200 DRF 7.64 7.45 7.45 7.44
SDRF 2.48 2.11 2.00 1.95
22,500 300 1,800 DRF 7.52 7.37 7.33 7.32
SDRF 2.16 1.83 1.74 1.74
Figure 1: Pointwise mean square error (MSE) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}_{1}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
Figure 2: Pointwise standard error (SD) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
Table 2 and Figures 12 compare SDRF with the i.i.d.-based DRF for estimating the regression functional $\Psi(\mathbf{x}_{1})=\mathbb{E}[Y_{1}\mid X_{1}=\mathbf{x}_{1}]$. Averaged over a uniform grid of 200 points in $[0,1]$, SDRF achieves smaller RMSE than DRF for all finite population sizes $N$ and ensemble sizes $B$ considered (Table 2).
Figure 1 shows the corresponding pointwise MSE: for both methods, MSE decreases with $B$, but DRF exhibits a clear error floor that persists as $N$ and $B$ increase, whereas SDRF’s MSE decreases uniformly towards the truth, including in the upper range of $X_{1}$ where sampling weights are more variable. Figure 2, combined with Figure 1, suggests that the MSE gap is driven primarily by persistent bias in DRF. This is expected as reasoned in Section 1. Also, Figure 2 shows that pointwise variability decreases with $B$ for both methods and remains comparable between SDRF and DRF; notably, SDRF does not exhibit an appreciable variance inflation despite the additional randomness introduced by its design-aware resampling step.
5 Biomedical Public Health Case Study
This section demonstrates the practical relevance of survey-calibrated distributional random forests (SDRF) for performing data analysis tasks in nationally representative studies. We use data from the National Health and Nutrition Examination Survey (NHANES), which employs a multistage, stratified, cluster sampling design to represent the U.S. civilian, noninstitutionalized population: primary sampling units (PSUs) are selected at the county level, followed by smaller geographic areas and then individuals. All analyzes performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx . More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g. split selection and node estimation).
The objective of the statistical analysis is to characterize the conditional joint distribution of two glucose biomarkers used to diagnose and monitor the progression of diabetes, fasting plasma glucose (FPG) and glycated hemoglobin (HbA1c), given demographic and anthropometric covariates: sex, age, and body mass index (BMI). Our primary clinical goal is to quantify how the population distribution of this bivariate outcome varies between these characteristics, to support the detection of diabetes, the evaluation of glycemic-control, and healthcare planning through the trends observed in the subgroups of the population. A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) [2], which ignore specific individual characteristics. For analytical purposes, a total of $N=2732$ individuals with and without diabetes were included in the analysis.
Let us go to define mathematically the problem. Let the clinical outcome be $Y=(Y_{1},Y_{2})^{\top}$, where $Y_{1}$ denotes fasting plasma glucose (FPG) and $Y_{2}$ denotes HbA1c. Let the covariate vector be $X=(X_{1},X_{2},X_{3})^{\top}$, where $X_{1}=\mathrm{Age}$ and $X_{2}=\mathrm{BMI}$ are continuous, and $X_{3}=\mathrm{Sex}\in\{0,1\}$ is binary with $0=\text{male}$ and $1=\text{female}$.
From an empirical perspective, we run SDRF and estimate the conditional distribution of the bivariate result $Y=(\mathrm{FPG},\mathrm{HbA1c})^{\top}$ given covariates $X=\mathbf{x}$. Specifically, we computed the estimated conditional CDF
$\hat{F}(y\mid\mathbf{x})=\hat{P}^{N}_{\mathrm{HT}}\big(Y\leq y\mid X=\mathbf{x}\big),\qquad\text{for all }y\in\mathcal{Y},\ \mathbf{x}\in\mathcal{X},$
where inequality $Y\leq y$ is understood componentwise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
$\mathrm{Age}\in\{30,50,70\},\quad\mathrm{BMI}\in\{20,30,40\},\quad\mathrm{Sex}\in\{0,1\}.$
Let $\hat{\omega}_{i}(x)$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $x\in\mathcal{X}$. Using these weights, we compute the estimated conditional mean and covariance,
$\hat{\mu}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,\bigl(Y_{i}-\hat{\mu}(x)\bigr)\bigl(Y_{i}-\hat{\mu}(x)\bigr)^{\top}.$
We then define a conditional Mahalanobis score (weighted) for any pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ by
$S(x,y)=\bigl(y-\hat{\mu}(x)\bigr)^{\top}\hat{\Sigma}(x)^{-1}\bigl(y-\hat{\mu}(x)\bigr).$
Let $\{(x_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(x_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
In the spirit of [53] (for finite-dimensional Euclidean responses) and [57] (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $x\in\mathcal{X}$ as
$\hat{C}^{\alpha}(x)=\bigl\{\,y\in\mathcal{Y}:\;S(x,y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
Compared to the natural competitor [53], our approach is more flexible: it requires a single model, avoids user-chosen smoothing parameters due to the use of multiple local polynomial regression, and enables valid inference under the complex design of the NHANES survey.
Figures 3 and 4 show tolerance regions (in expectation) for $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$ in the covariate combinations described above (a larger $1-\alpha$ implies a wider region). Figure 3 contains information on age and BMI, while Figure 4 shows gender differences. The FPG variable is plotted on a logarithmic scale. The colored lines indicate the diagnostic thresholds of the American Diabetes Association (ADA): $\mathrm{FPG}\geq 126\penalty 10000\ \mathrm{mg/dL}$ and $\mathrm{HbA1c}\geq 6.5\%$.
We discuss the statistically supported variable by variable. The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observation and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels. With increasing age, both FPG and HbA1c shift upward, indicating poorer glycemic control at older ages; this effect is most pronounced and consistent near the $0.9$ quantile. A higher body mass index (BMI) is moderately associated with higher glucose levels: people with $\mathrm{BMI}>30\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$ show higher values than those with $\mathrm{BMI}\in[20,30]\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$. Between $30$ and $40\penalty 10000\ \mathrm{kg}/\mathrm{m}^{2}$, the tolerance regions change little between the quantile levels. In general, glucose control differs between overweight/obesity individuals and those with a BMI in the normal range.
From these results, two implications follow: (i) glycemic control in the U.S. population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and political interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with [8] . From a modeling and analysis perspective, SDRF provides a principled framework to define reference values for specific subpopulations, reclassify individuals based on biomarker profiles, and forecast public–health risks leveraging the full conditional distribution. In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to less informative inform to perform the decision making.
Figure 3: SDRF–estimated tolerance regions (in expectation) for $\log(\mathrm{FPG})$ vs. $\mathrm{HbA1c}$ at quantile levels $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$. Panels display regions conditional on specific age groups (top two) and BMI ranges (bottom three). Green and gray points indicate observations falling inside and outside the estimated regions, respectively.
Figure 4: SDRF–estimated tolerance regions (in expectation) for $\log(\mathrm{FPG})$ vs. $\mathrm{HbA1c}$ stratified by gender. The contours correspond to $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$. Green and gray points indicate observations falling inside and outside the estimated regions, respectively.
6 Discussion
We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements [95]. To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, filling an important methodological gap. From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs to estimate conditional distributions for multivariate outcomes. Our approach leverages reproducing-kernel Hilbert spaces (RKHS) and distances between kernel mean embeddings to estimate the full conditional distribution. This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach [7, 67, 79, 37].
SDRF framework provides a powerful and flexible tool for biomedical studies conducted under complex survey designs, as illustrated by the NHANES case study reported here, studies that are essential for optimizing public-health decisions. Our machine learning algorithm for estimating the conditional distribution of multivariate outcomes represents a substantial modeling advance, enabling investigators to address clinically relevant questions in these cohorts with direct implications for health policy, as demonstrated in our case study illustration.
Survey data and machine learning offer promising synergies for more efficient estimation in finite sample regimes. Building on recent work, we have the potential to improve calibration, reduce overfitting, and improve algorithmic fairness in population-based studies, while mitigating selection bias and preserving practical causal model interpretations-—especially in nationally representative healthcare studies.
In future work, we will continue to investigate the survey-design aspects of our method. In particular, we will study the upward bias that can arise in ratio-type estimators when the denominator (here, the leaf size or subpopulation size) must be estimated because the true leaf sizes are not available from the finite population or census. Extending SDRF for outcome-dependent sampling is another direction.
From a data-analytic perspective, we will develop algorithms to handle non-Euclidean random objects [58] that arise in digital health applications, and we will design methods to quantify uncertainty (for example, via confidence bands) within our random forest frameworks. It is also of applied interest to extend our algorithm to functional data [24], which can be naturally accommodated by using kernel RKHS embeddings to define split rules. There are recent regression methods for functional data in complex survey designs, but they focus only on linear functional response models and the conditional regression mean operator [52, 65].

### v2

@@
 We propose a survey-calibrated distributional random forest (SDRF) that incorporates complex-design features via a pseudo-population bootstrap, PSU-level honesty, and a Maximum Mean Discrepancy (MMD) split criterion computed from kernel mean embeddings of Hájek-type (design-weighted) node distributions.
-We provide a framework for analyzing forest-style estimators under survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions.
+We provide a framework for analyzing forest-style estimators under various survey designs; establish design consistency for the finite-population target and model consistency for the super-population target under explicit conditions on the design, kernel, resampling multipliers, and tree partitions.
 As far as we are aware, these are the first results on model-free estimation of conditional distributions under survey designs.
@@
 1 Introduction and Motivation
-Experimental design is a key driver of successful statistical practice [97].
-A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings [71].
-Complex survey design [42] methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies [59].
-In the finite-population regime [25], a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals.
+Experimental design is a key driver of successful statistical practice yates1964sir.
+A rigorous study design protects and makes transparent the assumptions that link the data to scientific conclusions, thus improving the reliability and reproducibility of the empirical findings peng2015reproducibility.
+Complex survey design Han_Weller_2021 methods provide a powerful methodology for population-representative inference in a wide range of settings, such as biomedical studies lumley2011complex.
+In the finite-population regime DeMatteis2025, a probability sample with known inclusion probabilities and carefully constructed weights allows design-unbiased (or asymptotically design-unbiased) estimation of population functionals.
 These estimators can exhibit satisfactory statistical properties, such as efficiency, provided that appropriate statistical calibration and inference procedures are used.
 This design-based framework is particularly compelling when paired with randomized interventions: probability surveys deliver external validity (representativeness of a well-defined target population), while randomized experiments deliver internal validity (causal identification under the design).
-Together, they support the policy-relevant generalization of causal effects [45].
+Together, they support the policy-relevant generalization of causal effects cinelli2025challengesstatisticsdozenchallenges; imbens2015causal.
 Large, nationally representative health studies, such as the National Health and Nutrition Examination Survey (NHANES), which monitors health and nutrition in the U.S.
-population, are canonical examples of the impact of complex survey methodology in practice [68].
+population, are canonical examples of the impact of complex survey methodology in practice mozumdar2011persistent.
 When design information is properly incorporated and analyzed with appropriate statistical methods, such data can be used to quantify disease burden, monitor disparities, and inform clinical and public health policy in a cost-effective way, all while ensuring valid population inference.
@@
 Despite the central role of complex survey designs in nationally representative studies, predictive tools for complex surveys remain limited in two main respects.
-First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods [61].
+First, most existing methods target simple functionals such as conditional means or medians, rather than the full conditional distribution, and are typically based on linear models or classical nonparametric smoothers such as Nadaraya–Watson estimators and spline-based methods lumley2017fitting.
 Second, available procedures often presuppose low-dimensional covariates and focus on smoothing over relatively simple or very smooth function classes.
-By contrast, the use of modern machine-learning methods that can capture complex structures more effectively in high-dimensional settings such as neural networks [5, 98] or tree-based ensembles [26]-—within the survey context remains relatively scarce [66, 67].
-From an application’s perspective, many scientific questions demand distributional targets beyond means or medians [50], including multivariate quantiles, tolerance regions [34], tail probabilities, and other functionals of multivariate conditional laws.
+By contrast, modern machine-learning methods that can capture complex structure in high-dimensional settings—such as neural networks BauerKohler2019DeepLearningCurseDimensionality; zammit2022deep and tree-based ensembles diaz2006gene—remain relatively uncommon in survey applications matabuena2025screeningdiabetesmellitusus; matabuena2025kernel.
+In scientific applications, many questions demand distributional targets beyond means or medians kneib2023rage, including multivariate quantiles, tolerance regions fraser1953nonparametric, tail probabilities, and other functionals of multivariate conditional laws.
 In NHANES, for example, multivariate distribution estimates can be used to define tolerance regions that detect outlying biomarker profiles given subject-specific characteristics.
@@
 However, they are typically formulated under i.i.d.
-sampling and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference.
+sampling assumptions and, when naively applied to complex surveys, can distort node-wise conditional distributions, bias risk estimates, and compromise inference.
 A design-aware approach must propagate weights and clustering through splitting and aggregation, respect replicate-weight or bootstrap schemes that mimic the original selection, and produce distributional outputs for vector-valued $Y$, not merely conditional means.
 These requirements call for distributional learning procedures that integrate survey design into both the objective function and the algorithmic construction.
-From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings.
-Our paper.
+From a machine-learning perspective, integrating predictive algorithms with complex survey designs creates opportunities to increase the statistical efficiency of estimators, mitigate selection bias, reduce overfitting through design-guided regularization, improve subgroup fairness, and achieve proper predictive calibration in inference tasks in applied settings
+Our paper: Overview.
 We study nonparametric estimation of multivariate conditional distributions under complex survey designs.
-Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design.
+Let $(Y,X)$ denote a pair consisting of a vector-valued response $Y\in\mathcal{Y}$ and covariates $X\in\mathcal{X}\subseteq\mathbb{R}^{p}$ observed through a probability sample drawn under a complex survey design.
 We introduce a design-aware distributional random forest estimator $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ that combines survey-weighted tree construction and aggregation with loss functions induced by mean kernel embeddings of probability measures in a separable reproducing kernel Hilbert space.
 This yields a single forest-based procedure that (i) targets the full conditional law $\mathbb{P}(Y\mid X=\mathbf{x})$ and, as plug-in estimators, a broad class of component-wise and joint functionals such as conditional quantiles, cross-moments, covariance matrices and tail probabilities, and (ii) remains consistent under the joint law induced by the super–population model and the survey design.
-On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) [37, 80, 89], a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the forest weighing the survey achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity.
+By “full conditional law” we mean the entire probability distribution $\mathbb{P}(Y\mid X=\mathbf{x})$, not only summary functionals such as conditional means $\mathbb{E}[Y\mid X=\mathbf{x}]$ or quantiles.
+When $Y$ is multivariate, we estimate the joint conditional distribution, preserving dependence structure among components.
+On the theoretical side, we formulate the estimation problem in terms of the maximum mean discrepancy (MMD) gretton2012kernel; sejdinovic_equivalence_2013; szekely2003statistics, a family of statistical distances that metrizes weak convergence in suitable classes of probability measures, and show that the proposed design-aware forest achieves design and model consistency; that is, the MMD between $\widehat{\mathbb{P}}_{Y\mid X=\mathbf{x}}$ and $\mathbb{P}_{Y\mid X=\mathbf{x}}$ converges uniformly to zero over $x$ at rates governed by tree-partition and design complexity.
 We obtain a decomposition of the MMD risk into design, algorithmic, and stochastic terms, prove that each component vanishes under mild moment conditions on inclusion and replicate weights.
-Our analysis quantifies the impact of the number of trees $B$ and the bootstrap or replicate-weight draws $M$ on node-level predictions and on the variance of the ensemble, showing a reduction in variance of order $O(M^{-1/2})$ while preserving the consistency of the design.
+Crucially, we show that PSU-level sample splitting combined with pseudopopulation bootstrap preserves both honesty and design-based consistency, despite the dependency structure induced by the complex sampling mechanism.
 We emphasize that, even in the i.i.d.
-setting, the theory of classical random forests is far from trivial [9, 10] and has required substantial effort to develop since Breiman’s original proposal of random forests [17].
-Moreover, many questions remain open, for instance regarding finite-sample guarantees [55] or extensions to dependent data [36].
+setting, the theory of classical random forests is far from trivial biau2008consistency; biau2016random and has required substantial effort to develop since Breiman’s original proposal of random forests breiman2001random.
+Moreover, many questions remain open, for instance regarding finite-sample guarantees JMLR:v26:24-0255 or extensions to longitudinal data goehry2020random@book.
 We complement the theory with an application to NHANES, where we estimate the joint conditional distribution of two biomarkers used in diabetes diagnosis based on individual-specific characteristics, providing distributional estimates that can inform a range of decisions from a healthcare and public-health perspective.
@@
 Our work is based on two strands.
-On the survey side, foundational contributions make the sampling mechanism explicit through survey weights [96, 84, 40, 77, 30, 82, 81, 39], including the Horvitz–Thompson and Hájek estimators [44, 28], as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression [51, 75, 61, 74, 70, 47].
-On the machine-learning side, we exploit the statistical learning properties of random forests [17] to reduce variance estimation and even bias in many regimes via the bagging step [15, 55].
-Random forests [17] and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks [31] and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks.
+On the survey side, foundational contributions make the sampling mechanism explicit through survey weights yates1953sampling; hajek1960limiting; roy1960estimating; ericson1969subjective; shindell1963population; hajek1959optimal, including the Horvitz–Thompson and Hájek estimators horvitz1952generalization; dupavc1975jaroslav, as well as design-aware procedures widely used, for instance, in model checking, hypothesis testing, and regression koch1975strategies; rao1987simple; lumley2017fitting; rabe2006multilevel; natarajan2012extension; jimenez2018estimation.
+On the machine-learning side, we exploit the statistical learning properties of random forests breiman2001random to reduce variance estimation and even bias in many regimes via the bagging step breiman1996bagging; JMLR:v26:24-0255.
+Random forests breiman2001random and their distributional variants (through kernel mean embeddings and MMD) provide flexible nonparametric tools that achieve competitive performance in high-dimensional settings across many benchmarks fernandez2014we and can be used efficiently for two-sample testing, variable selection, and clustering, in addition to predictive tasks.
 However, existing random forest algorithms are largely developed for i.i.d.
 data   design-aware random forest methodology and theory for complex surveys remain underdeveloped.
-From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation [91, 69].
-Classical nonparametric smoothers—including local polynomials [48], Nadaraya–Watson estimators [43], and spline-based additive models—have been adapted to complex surveys [14].
-However, most of this literature focuses on continuous outcomes [83, 46] and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
-From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs [6, 60, 78, 11].
+From a regression-model perspective, existing survey methods focus primarily on parametric and semiparametric estimation using (pseudo) probability weights and estimating equations related to M- and Z-estimation van1996m; nan2013general.
+Classical nonparametric smoothers—including local polynomials kikechi2017local, Nadaraya–Watson estimators harms2010kernel, and spline-based additive models—have been adapted to complex surveys breidt2005model.
+However, most of this literature focuses on continuous outcomes signorino2018using; iparragirre2023variable and conditional learning; comparatively less attention has been paid to multivariate outcomes and full conditional learning.
+From a theoretical perspective, design-based central limit theorems are available for a variety of survey designs bellhouse2001central; lumley2012empirical; schochet2022design; boistard2017functional.
 Empirical-process formulations replace the uniform empirical measure with Horvitz–Thompson/Hájek-weighted measures; under suitable weight conditions, the resulting processes inherit many i.i.d.
 properties.
-Survey-aware resampling is well–developed [76]: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals.
+Survey-aware resampling is well–developed rao1988resampling: bootstrap and replicate-weight methods with asymptotic guaranties approximate sampling distributions and can be adapted to distributional functionals.
 However, existing theoretical developments are not directly tailored to distributional random forests under complex designs, where the interaction between algorithmic randomness, survey weights, and distributional loss functions requires a novel theoretical survey framework.
 In this paper, we provide a design-aware distributional random forest for survey data, together with a corresponding theoretical framework for its design-based and model-based properties.
-When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Table S3.
+When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential for statistical consistency, as we illustrate in Section S2 of the Supplementary Material.
 The remainder of the paper is organized as follows.
@@
 2 Set-up and Background
-Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i=1}^{n_{s}}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly.
+For notational convenience, for any $a\in\mathbb{Z}_{>0}$, we will denote $a]:=\{1,\dots,a\}$.
+‘:=’ will be used to emphasize definition.
+Suppose that we observe a sample $D=\{(X_{i},Y_{i},w_{i})\}_{i\in[n_{s}]}$ drawn from a finite population $D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i\in[N]}$ of size $N$ under a sampling design $\mathbf{p}$ that may depend on auxiliary design variables $Z_{i}$, which we will formalize shortly.
 Here, $Y_{i}\in\mathcal{Y}$ is the outcome, $X_{i}\in\mathbb{R}^{p}$ are covariates and $w_{i}\in\mathbb{R}^{+}$ are survey weights (possibly incorporating nonresponse adjustment and calibration).
@@
 Let $P_{Y\mid X=\mathbf{x}}$ denote the super-population conditional law.
-Consider a fixed covariate value $\mathbf{x}\in\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and it is continuous at the target.
+Consider a fixed covariate value $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{p}$ and a measurable functional $\Psi:\mathcal{M}_{b}(\mathcal{Y})\to(\mathcal{V},d_{v})$ taking on some probability measures on $\mathcal{Y}$ in some separable metric space $(\mathcal{V},d_{v})$, and which is continuous at the target.
 Our objective is to construct an estimator $\widehat{\Psi}_{N}(\mathbf{x})$ that is design-consistent for the finite-population target $\Psi(P^{N}_{Y\mid X=\mathbf{x}})$ and model-consistent for the super-population target $\Psi(P_{Y\mid X=\mathbf{x}})$.
 We will work on the joint probability space $(\Omega\times\mathcal{S}^{N},\,\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega\times\mathcal{S}^{N}})$.
-Specifically, denote $P_{\mathcal{S}^{N}|\omega}$ the measure in the design space, we want that for any $\varepsilon>0$,
-$P_{S^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
+Specifically, denoting $P_{\mathcal{S}^{N}|\omega}$ as the measure in the design space, we want that for any $\varepsilon>0$,
+$P_{\mathcal{S}^{N}\mid\omega}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P^{N}_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{$P_{\Omega}$--a.e.\,(design consistency), and }$
 $P_{\Omega\times\mathcal{S}^{N}}\!\left(d_{v}\!\left(\widehat{\Psi}_{N}(\mathbf{x}),\Psi(P_{Y\mid X=\mathbf{x}})\right)>\varepsilon\right)\to 0\quad\text{ (model consistency). }$
 2.1 Survey Design
-We follow the finite population framework considered by [13, 41, 42]: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$.
+We follow the finite population framework considered by boistard_functional_2017; Han_Weller_2019; Han_Weller_2021: Consider a sequence of finite populations $\{\mathcal{U}^{N}\}_{N\geq 1}$.
 Let $U^{N}:=\{1,\dots,N\}$, the index set for each $\mathcal{U}^{N}$ and $\mathcal{S}^{N}:=\{S:S\subset U^{N}\}$ the collection of subsets of $U^{N}$.
@@
 For any region ${A}(\mathbf{x})\subset\mathcal{X}$ indexed by some fixed $\mathbf{x}\in\mathcal{X}$, denote the empirical conditional measure induced by ${D}^{N}$ as
-${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}I(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
-$N(\mathbf{x})=\sum_{i\in[N]}I(X_{i}\in A(\mathbf{x}))$.
-A sampling scheme can be formalized as a transition function $\bm{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$.
-From $\bm{p}$, fix any $\omega\in\Omega$, we define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\bm{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$.
+${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in{A}(\mathbf{x}))\;\delta_{Y_{i}},\text{ where $\delta_{Y_{i}}$ the Dirac measure,}$
+$N(\mathbf{x})=\sum_{i\in[N]}\mathbbm{1}(X_{i}\in A(\mathbf{x}))$.
+A sampling scheme can be formalized as a transition function $\mathbf{p}:\Omega\times\sigma(\mathcal{S}^{N})\mapsto[0,1]$.
+Given $\mathbf{p}$, fix any $\omega\in\Omega$ and define a probability measure in the design space $(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}))$ as $P_{\mathcal{S}^{N}}(A):=\sum_{S\in A}\mathbf{p}({Z}^{N}(\omega),S)$, $\forall A\in\sigma(\mathcal{S}^{N})$.
 We can then uniquely define a product probability measure in the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$ as
-$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\bm{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
-Thus, a random sample $S\sim P_{\mathcal{S}^{N}}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$.
-Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\bm{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$.
-Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$, the size of the random sample $S$ that is a random variable.
+$P_{\Omega\times\mathcal{S}^{N}}(E\times S):=\int_{E}\mathbf{p}(\omega,\{S\}){P}(d\omega),\hskip 8.5359pt\forall E\times\{S\}\text{ in simple rectangles in }\mathcal{A}\otimes\sigma(\mathcal{S}^{N}).$
+Thus, a random sample $S\sim P_{\mathcal{S}^{N}|\omega}$ can be obtained by ${\xi}^{N}=(\xi_{1},\dots,\xi_{N})$, a random vector on $\mathcal{S}^{N}$, where $\xi_{i}:=\mathbbm{1}_{(i\in S)}\in\{0,1\}$.
+Define first-order inclusion probabilities $\pi_{i}(\omega):=\mathbb{E}_{\mathcal{S}}(\xi_{i}|Z(\omega))=\sum_{\{S:i\in S\}}\mathbf{p}(S,{Z}^{N}(\omega))$, and second-order inclusion probabilities $\pi_{ij}(\omega):=\mathbb{E}(\xi_{i}\xi_{j}|Z(\omega))$.
+Denote $n_{s}:=card(S)=\sum_{i=1}^{N}\xi_{i}$ as the size of the random sample $S$.
 We do not have access to the entire finite population of size $N$, but we work with a sample of size $n_{s}$.
-This leads to estimators that use weights to construct consistent estimators ([56]).
+This leads to estimators that use weights to construct consistent estimators (Lohr2021-cm).
 For example, the Hájek (HJ) estimator for the true conditional probability ${P}^{N}_{Y|X\in A(\mathbf{x})}(\mathbf{x})$ in the design space is
-$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}I(X_{i}\in A(\mathbf{x})).$
-The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts the form Hájek, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$.
+$\widehat{P}^{N}_{HJ,Y|X\in A(\mathbf{x})}(\mathbf{x})=\frac{1}{\widehat{N}(\mathbf{x})}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x}))\;\delta_{Y_{i}},\quad\widehat{N}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x})).$
+The SDRF estimator, which we denote as $\widehat{\Psi}_{N}(\mathbf{x})$, adopts a Hájek-like form, but with critical modifications from design-compatible resampling and an adaptive data-driven partition $L^{*}(\mathbf{x})$ in place of $A(\mathbf{x})$.
 We assume the following with respect to the design.
@@
   • (D2)
-    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, with some $\kappa>4$.
+    Stable sampling fraction: $n_{s}/N\rightarrow_{p_{d}}f$ for some $f\in(0,1]$, and $\exists\delta\in(0,\;f)$, $C<\infty$, such that for $N$ large enough, $P_{\mathcal{S}^{N}|\omega}(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$, for some $\kappa>4$.
   • (D3)
@@
 This prevents extreme survey weights.
-(D4) is common [12, 11, 19] and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.
+(D4) is common Boistard2012-gc; boistard2017functional; Cardot2010-cn and a mild second order condition that controls the dependence among the inclusion indicators $\{\xi_{i}\}_{i=1}^{N}$ to be of order $1/n_{s}$.
 Lemma 2.1.
-Under (D2)-(D4), we have that w.r.t the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
-Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D1)–(D4).
-Two–stage designs that satisfy a counter-part of (D1)–(D4) at each of the two stages satisfy also (D1)–(D4) at the subject level (Lemma A.2, Supplementary Material).
+Under (D2)-(D4), we have that with respect to the design law $P_{\mathcal{S}^{N}|\omega}$, $\frac{1}{N}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\xrightarrow{p_{d}}0$, and $\frac{1}{\sqrt{N}}\sum_{i=1}^{N}\!\left(\frac{\xi_{i}}{\pi_{i}}-1\right)=O_{p_{d}}(1).$
+Simple random sampling without replacement (SRSWOR), Poisson (Bernoulli) sampling, SRSWOR, PPS with replacement with MOS bounded above and away from zero, and their stratified versions are all examples that satisfy (D2)–(D4) (Lemma S1, Supplementary Material).
+Two–stage designs that satisfy a counterpart of (D2)–(D4) at each of the two stages satisfy also (D2)–(D4) at the subject level (Lemma S2, Supplementary Material).
 2.2 Embedding of Measures
-Kernel mean embeddings (KMEs) [38, 85] provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) [1].
+Kernel mean embeddings (KMEs) Gretton_2007; Smola_2007 provide a flexible and conceptually simple way to represent probability distributions as single points in a reproducing kernel Hilbert space (RKHS) Aronszajn1950.
 Since these representations are defined purely through expectations of the kernel evaluations, all operations on distributions are reduced to linear operations in the RKHS.
@@
 It is a fully nonparametric way to compare distributions, and for characteristic kernels it is equal to zero only when the two distributions are the same.
-This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., [86, 87, 37, 80, 33, 67].
+This framework supports a wide range of kernel based methods for two-sample testing, independence testing, clustering, biclustering and related problems; see, e.g., Sriperumbudur2010-xr; sripernumbudur_hilbert_2010; gretton2012kernel; sejdinovic_equivalence_2013; francca2020kernel; matabuena2025kernel.
 From the perspective of constructing random forests for distributional targets, incorporating KMEs into the algorithm offers several practical advantages: (i) It provides a distribution-level target, enabling random forests to discriminate conditional laws via an MMD-based splitting criterion.
-(ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size [32].
-(iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data [62]).
+(ii) Empirical MMD can be computationally advantageous relative to alternative statistical distances, such as bounded Lipschitz or Wasserstein distances, for example, when $\mathcal{Y}=\mathbb{R}^{d}$ and $d$ is of moderate size Fournier2015-mu.
+(iii) From an implementation standpoint, using MMD is simple: it only requires changing how distances between random objects in the ambient space are computed (e.g., in a general Hilbert space or, more generally, in a metric space of negative type, as in the case of spherical data lyons2020strong).
 As a result, modifications to a standard random forest algorithm are minimal, even for outcomes in abstract spaces.
-(iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to both finite-population and super-population conditional laws under standard regularity conditions, allowing us to analyze a broad class of sampling designs within a single metric framework.
+(iv) Most importantly, because the KME is defined directly on probability measures, it applies equally to finite-population and super-population conditional laws under standard regularity conditions, allowing for the analysis of a broad class of sampling designs within a unified metric framework.
 We now introduce the corresponding concepts and definitions from a mathematical perspective.
@@
 That is, for any sequence $P_{1},\dots\in\mathcal{M}_{b}(\mathcal{Y})$, $P_{N}\rightsquigarrow P\iff d_{\mathbf{k}}(P_{N},\;P)\to 0$, as $N\to\infty$.
-We call kernels that satisfy this property universal (initially proposed in [88]).
+We call kernels that satisfy this property universal (initially proposed in Steinwart2001-ja).
 This property concerns the approximation ability of the KME sample estimate to the true target as $N\to\infty$.
-We place the following assumptions:
+We require the following assumptions
 On Space $\mathcal{Y}$ and Kernel $\mathbf{k}$
@@
   • (K1)
-    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of [86]).
+    $\mathbf{k}$ is uniformly bounded, positive definite, and $C_{0}$-continuous, where $C_{0}$-continuous means $\mathbf{k}(\cdot,y)\in C_{0}(\mathcal{Y})$ for all $y\in\mathcal{Y}$, where $C_{0}(\mathcal{Y})$ denotes the Banach space of bounded continuous functions vanishing at infinity (Section 2 of Sriperumbudur2010-xr).
   • (K2)
     $\mathbf{k}$ is $c_{0}$-universal.
-Meaning that the RKHS $\mathcal{H}$ induced by $\bm{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t.
-the uniform norm (Definition 2, Theorem 1 in [20]).
+Meaning that the RKHS $\mathcal{H}$ induced by $\mathbf{k}$ satisfying (K1) is dense in $C_{0}(\mathcal{Y})$ defined above w.r.t.
+the uniform norm (Definition 2, Theorem 1 in Carmeli2010-ps).
   • (K3)
@@
 Assumption (K3) is not needed for the statistical theory, but when satisfied, admits a simple finite-dimensional approximation of the MMD through Fourier-feature representation for fast computation.
-For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ ([38]); More details are provided in Section A.10 of the Supplementary Material.
-Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding; see Section A.10.
+For example, when $\mathcal{Y}=\mathbb{R}^{d}$, the rate of $d_{\mathbf{k}}$ can be independent of $d$ Gretton_2007; More details are provided in Section S10, Supplementary Material.
+Assumptions (S1) and (K1)–(K3) cover many common outcome types, including $\mathcal{Y}=\mathbb{R}^{d}$, graph representations in $0,1]^{m(m-1)/2}$ or $\mathbb{R}^{m\times m}_{\mathrm{sym}}$, torus-valued data $\mathbb{T}^{d}$ (via $\theta\mapsto(\cos\theta,\sin\theta)$), and the simplex $\Delta^{p-1}$, by equipping $\mathcal{Y}$ with standard kernels such as the Gaussian RBF $\mathbf{k}(y,y^{\prime})=\exp\{-\|y-y^{\prime}\|^{2}/(2\sigma^{2})\}$ with $\sigma>0,$ or its restriction after embedding.
 3 SDRF Algorithm and Theoretical Result
-Random forests are machine learning methods that combine a CART-style [27] tree ensemble with a bagging step [15], and can also be viewed as kernel-type estimators (see, e.g.
-[54, 72]).
-Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems [17], and several works have documented their strong empirical performance [26, 31].
+Random forests are machine learning methods that combine a CART-style donoho1997cart tree ensemble with a bagging step breiman1996bagging, and can also be viewed as kernel-type estimators (see, e.g.
+lin2006random; qiu2024random).
+Traditional random forests were designed to estimate the conditional mean regression function and prediction in classification problems breiman2001random, and several works have documented their strong empirical performance diaz2006gene; fernandez2014we.
 More recently, conditional distributional versions of random forests based on kernel mean embeddings have been proposed for i.i.d.
-data [21].
+data cevid_michel_meinshausen_bühlmann_2022.
 Our proposed method, SDRF, constructs a kernel estimator $\widehat{\omega}_{\mathbf{x}}(\cdot)$ that directly targets conditional distributions and appropriately incorporates the survey design.
@@
 3:  Draw resample $D^{*}_{b}=\{(X_{i},Y_{i},1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$ respecting the design.
-4:  Draw $\forall k(i)$, $I(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
+4:  Draw $\forall k(i)$, $\mathbbm{1}(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$.
 5:  Split at the PSU level $D^{*}_{b}=D^{*}_{b,split}\bigcup D^{*}_{b,est}$.
@@
 10:end for
-11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,I(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
+11:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$, $\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}$.
 12: For $\Psi(\cdot)$, obtain plug-in estimator $\Psi(\widehat{P}^{N}_{Y\mid X=\mathbf{x}})=\Psi(\sum_{i=1}^{n_{s}}\widehat{\omega}_{i}(\mathbf{x})\;\mathbf{k}(Y_{i},\,\cdot))$.
 3.1 Bagging and Honesty
-Resampling with Design
-Bagging is a key ingredient of random forests [15], but its study is primarily conducted in an i.i.d.
+Bagging is a key ingredient of random forests breiman1996bagging, but its study is primarily conducted in an i.i.d.
 setting.
 In the i.i.d.
-case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators [16, 18].
+case, it is well established that bagging introduces diversity across trees and allows the ensemble to approximate complex regression operators breiman1996bagging; Buhlmann2002-vp.
 Under a survey design, however, Efron’s classical i.i.d.
-bootstrap [29] can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
-We adopt a design–aware bootstrap (BS) method in line with [95] as the default resampling method for SDRF.
+bootstrap Efron1993-eg can distort the dependence structure and unequal-probability features induced by the design, and thus may fail to preserve the survey-weighted empirical measure in a controlled way.
+We adopt a design–aware bootstrap (BS) method in line with wang2022bootstrap as the default resampling method for SDRF.
 It generates multipliers $n^{*}$ that determine the additional up– or down–weighting applied to the sampled distribution, and satisfy properties (R1)–(R3) as detailed in Lemma 3.1.
 Lemma 3.1.
-For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section A.5.1 of the Supplementary Material in the spirit of [95] (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
-  1. On resample multipliers $n^{*}$
+For single–stage designs including Poisson, SRSWOR, and PPS-with-replacement designs, and two–stage clustering designs with bounded cluster sample-size, possibly with stratification, the pseudo-population bootstrap described in Section S5.1 of the Supplement in the spirit of wang2022bootstrap (with minor modification) generates unit (or SSU–level) multipliers $\{n_{i}^{*}\}_{i\in[n_{s}]}$ satisfying:
+On resample multipliers $n^{*}$
   1. (R1)
-    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains known cluster and strata structure.
+    The conditional law of $n^{*}_{i}$ depends on the design only through $\xi^{N}$ and $Z^{N}$ that contains the known cluster and strata structure.
 Each stratum is re–sampled separately, and if the design has two stages, the second stage resampling multipliers are independent across distinct PSUs conditional on first–stage resamples.
@@
 It is convenient for deriving concentration bounds while allowing for the existence of dependency.
-The method by [95] belongs to the class of pseudo-population-based resampling methods.
+The method by wang2022bootstrap belongs to the class of pseudo-population-based resampling methods.
 To summarize: For each re-sample, we first use design information such as $\{w_{i}\}_{i\in[n_{s}]}$ to re-construct via the multinomial distribution a pseudo finite population, with expected size equal to $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$; then we re-apply the sampling design $\mathbf{p}$ on the re-constructed finite population to obtain re-samples, thereby respecting the design structure.
-We refer interested readers to [63, 23] for a review of common resampling methods under design.
-Honesty with Design
-For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ ([3, 93, 4]).
+We refer interested readers to Mashreghi2016-nh; Conti_2022 for a review of common resampling methods under design.
+For random forests algorithms, honesty refers to $D_{est}\mathrel{\perp\!\!\!\perp}D_{split}$, where $D_{split}$ is used for construction of the tree structure $L$ and $D_{est}$ is used for populating the leaves with values to form the kernel estimate $\widehat{\omega}(\mathbf{x})$ Athey2016-hs; Wager2018-lu; Athey2019-xy.
 With i.i.d.
@@
 Alternatively, if one has access to data obtained from a continuous survey under the same design and is willing to assume that $P_{Y|X}$ stays unchanged between two consecutive time points, it is also possible to directly use one sample as $D_{est}$ and the other as $D_{split}$.
-In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, [49]) decreases as depth increases.
+In practice, when the number of sampled PSUs is small, the effective sample size (for example, $n_{\text{eff}}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$ for single–stage designs, Kish1995-aq) decreases as depth increases.
 As we discuss later, the performance of the MMD–based split criterion hinges on the effective sample size of $D_{split}$.
@@
 This partitions $P_{a}$ into child regions $C_{L}(\theta):=\{x\in P_{a}:x_{j}\leq t\}$ and $C_{R}(\theta):=\{x\in P_{a}:\;x_{j}>t\}$.
-With RF considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via Hájek–type estimator.
+With RF being considered as an adaptive kernel estimation algorithm, given any fixed region $L$, we can estimate the conditional probability via a Hájek–type estimator.
 We find the split by maximizing the MMD between the KMEs of the distribution estimates between the two child regions.
 Formally, $\hat{\theta}=\arg\max_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)$, where
-$\begin{split}M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},\end{split}$ (1)
-with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}I(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$.
-When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if exists.
+$M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\,\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in[n_{s}]:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}},$
+with $\hat{N}_{\ell}=\sum_{i\in[n_{s}]}\mathbbm{1}(X_{i}\in C_{\ell})\frac{n^{*}_{b,i}}{q\times\pi_{i}}$, $\ell\in\{L,R\}$.
+When under stratified two–stage designs, notice that the same form for $M^{*}_{n_{s}}$ holds, except that we replace the sum over subjects by sum $\sum_{h,k}\sum_{i\in\mathcal{U}_{hk}}$ over SSUs $(h,k)$ and then over subjects within SSUs, if they exist.
 We first provide a theorem justifying the choice of this split criterion.
 Theorem 3.2 (Local design/model consistency of the MMD split).
-Fix a parent region $P_{a}\subset\mathcal{X}$ at depth $d$ and consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which defines child regions $C_{L}(\theta)$ and $C_{R}(\theta)$.
-Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score in (1), and let $M^{N}(\theta)$ be its finite-population analog.
-Suppose (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3.
-Suppose further $\mathcal{X}\subset\mathbb{R}^{p}$ is bounded.
+For a parent region $P_{a}\subseteq\mathcal{X}$ at depth $d$ generated by the tree-growing process, consider axis-aligned potential splits $\theta=(j,t)\in\Pi:=\{(j,t):j\in[p],\ t\in[a_{j},b_{j}]\}$ which define child regions $C_{L}(\theta)$ and $C_{R}(\theta)$.
+Let $M^{*}_{n_{s}}(\theta)$ denote the (design-resampled) split score, and let $M^{N}(\theta)$ be its finite-population analog.
+Assume (D1)–(D4), (R1)–(R3), (S1), (K1)–(K2), and (A2)–(A3) described in Section 3.3.
+Suppose further that $\mathcal{X}\subseteq\mathbb{R}^{p}$ is bounded.
 Then, conditional on the realized finite population,
@@
 If, in addition, the finite population is an i.i.d.
-realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}$ and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta_{0}$ under the super-population model).
+realization from a super-population law $P$ and $M(\theta)$ denotes the corresponding population score, then $\sup_{\theta\in\Pi}\bigl|M^{*}_{n_{s}}(\theta)-M(\theta)\bigr|\xrightarrow{p_{\otimes}}0.$ Moreover, if $M^{N}(\theta)$ has a unique maximizer $\theta_{0}\in\Pi$, $M(\theta)$ has a unique maximizer $\theta^{\dagger}\in\Pi$, and the splits are chosen over a random set of candidates $\Pi(U)$ induced by mtrymtry features such that $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$, then $M^{N}$ satisfies the modulus of continuity condition with $d(\theta,\theta^{\prime}):=d((j,t),(j^{\prime},t^{\prime}))=1(j\neq j^{\prime})+|t-t^{\prime}|$, and any maximizer $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$ satisfies $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$ (and hence also $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta^{\dagger}$ under the super-population model).
 Remark.
 Theorem 3.2 provides a local result: for a fixed parent node $P_{a}$, the empirical MMD criterion $M^{*}_{n_{s}}(\theta)$ uniformly approximates its finite-population oracle counterpart $M^{N}(\theta)$ and also the super-population score $M(\theta)$.
-The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique ([90]).
+The uniformity over $\theta\in\Pi$ is essential for split selection, as it permits an argmax continuous-mapping argument and yields consistency of the selected split $\hat{\theta}_{n_{s}}$ whenever the oracle maximizer $\theta_{0}$ is unique van-der-Vaart2023-ur.
 The condition $\mathbb{E}_{U}[\mathrm{dist}(\theta_{0},\Pi(U))]\to 0$ makes explicit that random feature subsampling (via mtrymtry) does not distort the split choice asymptotically, provided the candidate set becomes sufficiently dense around $\theta_{0}$ in expectation.
-Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to that if using the full finite population, which is in particular assumption (B2) of Theorem 3.4 required for consistency in MMD metric.
-One might wonder that, since resamples are determined by non–negative multipliers $n^{*}$, if one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample.
-This can help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees.
-Proposition 3.3 below provides a detailed characterization of its consequence.
+Although extending from a fixed node to an entire data-adaptive tree requires additional control of the growing process, this theorem sheds light on why one might believe the partitions found via resampled data to be close to those obtained from using the full finite population, which is in particular assumption (B2) of Theorem 3.4 which is required for consistency in the MMD metric.
+This raises the question that, since resamples are determined by non–negative multipliers $n^{*}$, can one can use an averaged multiplier from $M$ (conditional) independent draws to form one re-sample?
+This could help stabilize the resample, especially when the original sampling weights vary greatly, but can also reduce diversity across trees.
+Proposition 3.3 below provides an affirmative answer to this question.
 Proposition 3.3 (Effect of averaging $M$ resampling draws on split-score stability).
-Fix a node $P_{a}$ and a candidate split $\theta$ that induces children $C_{L},C_{R}$ with $N_{P_{a}}:=|\{i\leq N:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\leq N:X_{i}\in C_{\ell}\}|$.
-Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M_{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ its design-weighted sample analogue, and $M_{n_{s}}^{*(M)}(\theta)$ the resampled score computed using averaged multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$ (equivalently $\bar{\delta}_{i}:=\bar{n}_{i}^{*}-1$).
-Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$, which are satisfied by the methods described in Section A.5.1, then
-$\bigl|M_{n_{s}}^{*(M)}(\theta)-M_{N}(\theta)\bigr|=O_{p}\!\left(N_{P_{a}}^{-1/2}\right)\;+\;O_{p}\!\left(\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\right),$
-where $n_{\mathrm{eff},\ell}$ is the design-weighted effective sample size ([49]) in child $\ell$ induced by $\theta$.
+Fix a node region $P_{a}\subseteq\mathcal{X}$ and a candidate split $\theta=(j,t)$ that induces children $C_{L},C_{R}$.
+Define ${N}_{P_{a}}:=|\{i\in[N]:X_{i}\in P_{a}\}|$, $N_{\ell}:=|\{i\in[N]:X_{i}\in C_{\ell}\}|$.
+Suppose there exists $\alpha\in(0,1/2]$ such that $\frac{N_{L}}{N_{P_{a}}},\ \frac{N_{R}}{N_{P_{a}}}\in[\alpha,1-\alpha].$ Let $M^{N}(\theta)$ be the finite-population split score, $M_{n_{s}}(\theta)$ be the design-weighted sample score, and $M_{n_{s}}^{*(M)}(\theta)$ be the score using averaged bootstrap multipliers $\bar{n}_{i}^{*}:=M^{-1}\sum_{b=1}^{M}n_{b,i}^{*}$.
+Under (D1)–(D4), (R1)–(R3), and (K1)–(K2), if the multiplier increments $\delta_{b,i}:=n_{b,i}^{*}-1$ are conditionally sub-gamma and negatively dependent given $\xi$ (satisfied by the re-sampling methods described in Section S5.1, Supplementary Material, by Lemma S6, Supplementary Material), then
+$\bigl|M_{n_{s}}^{*(M)}(\theta)-M^{N}(\theta)\bigr|=O_{p_{d}}\!\left(N_{P_{a}}^{-1/2}\;+\;\frac{1}{\sqrt{M}}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-1/2}\right),$
+where $n_{\mathrm{eff},\ell}=\frac{(\sum_{i\in S\bigcap C_{\ell}}\pi_{i}^{-1})^{2}}{\sum_{i\in S\bigcap C_{\ell}}\pi_{i}^{-2}}$ is the design-weighted effective sample size (Kish1995-aq) in child $\ell$.
 Remark.
-Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M_{N}(\theta)|=O_{p}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-3/2}\big)$.
-The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\\ n_{\mathrm{eff},\ell}^{-3/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains.
+Proposition 3.3 separates two sources of split-score fluctuation: an intrinsic design term $|M_{n_{s}}(\theta)-M^{N}(\theta)|=O_{p_{d}}(N_{P_{a}}^{-1/2})$ which concerns the sampling noise already present even with fixed $n^{*}$, and a resampling term $|M_{n_{s}}^{*(M)}(\theta)-M_{n_{s}}(\theta)|=O_{p_{n^{*}|\xi}}\!\big(M^{-1/2}\sum_{\ell\in\{L,R\}}n_{\mathrm{eff},\ell}^{-1/2}\big)$.
+The cross-over threshold occurs when these terms balance: if $M\gtrsim\tilde{M}$ where $\tilde{M}\asymp N_{P_{a}}\!\big(\sum_{\ell\in\{L,R\}}\,n_{\mathrm{eff},\ell}^{-1/2}\big)^{2},$ then the design term dominates and increasing $M$ further yields negligible stability gains.
 At deep nodes (small $n_{\mathrm{eff},\ell}$), this threshold is reached quickly: for instance, if both children have $n_{\mathrm{eff},\ell}\approx 100$ and $N_{P_{a}}\approx 200$, then $\tilde{M}\approx 8$ suffices.
-Beyond this point, split stability is governed by the effective sample sizes—controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions qq—rather than by the number of resamples.
+Beyond this point, split stability is governed by the effective sample sizes, controlled via minimum-node-size constraints, shallower depths, or larger subsampling fractions $q$, rather than by the number of resamples.
 Moreover, the marginal gain from increasing $M$ deteriorates rapidly: doubling from $M$ to $2M$ reduces the resampling component by only $2^{-1/2}\approx 0.71$, and more generally $(1+\Delta/M)^{-1/2}\approx 1-\Delta/(2M)$ when $\Delta\ll M$.
@@
   • (A2)
-    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ the ball centered at $\mathbf{x}$ with radius $r$.
+    There exist constants $0<m<M<\infty$ such that for all radius r, of some $L(\mathbf{x})$ for the given $\mathbf{x}$ such that $0<r\leq k_{N}$, $mr^{p}\leq\#\{i:X_{i}\in B_{r}(\mathbf{x})\}/N\leq Mr^{p}$, where $B_{r}(\mathbf{x})$ is the ball centered at $\mathbf{x}$ with radius $r$.
   • (A3)
     For any $L^{*}$ found by $D_{split}$, when populating the values with $D_{est}$, all leaves receive at least $n_{L^{*},min}$ subjects.
-In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}I(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
+In other words, $n_{L^{*}(\mathbf{x})}=\sum_{i}^{N}\xi_{i}\mathbbm{1}(X_{i}\in L^{*}(\mathbf{x}))\geq n_{L^{*},min}>0$.
 Algorithmic behavior
   • (B1)
-    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ yet $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $iam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2}), dd$ some metric on $\mathcal{X}^{p}$.
+    For any $\mathbf{x}$ and some deterministic chosen sequence $k_{N}\to 0$ such that $k^{p}_{N}N\to\infty$ as $N\to\infty$, there exists $t_{N}\to 0$ such that $P_{L}(diam(L(\mathbf{x}))\geq k_{N})\leq t_{N}$, where $diam(L(\mathbf{x}))=\sup_{x_{1},x_{2}\in L(\mathbf{x})}d(x_{1},x_{2})$, and $d$ is some metric on $\mathcal{X}^{p}$.
   • (B2)
-    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(x)\,\Delta\,L(x)\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
-Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the estimate from SDRF.
+    $\sup_{x\in\mathcal{X}_{1}\cup\mathcal{X}_{2}}\mathbb{P}\{X\in L^{*}(\mathbf{x})\,\Delta\,L(\mathbf{x})\}=o_{p}(d_{N})$ for some $d_{N}=o(k_{N}^{p})$, where $\Delta$ denotes symmetric difference: $A\Delta B:=(A\setminus B)\;\cup\;(B\setminus A)$.
+Denote $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ as the SDRF estimate.
 Then $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\to_{p_{d}}0$ as $N,B\to\infty$.
@@
 Regularization–type conditions (A1)–(A3) can be controlled via adaptive stopping criteria.
-(A1) imposes a geometric control on the shape of $L$ so that they do not be stretched.
+(A1) imposes a geometric control on the shape of $L$ to control the stretching.
 (A2) requires a local mass control for the local empirical density to be well-behaved.
-(A3) ensures the local density estimate stays bounded.
+(A3) ensures that the local density estimate stays bounded.
 (B1)-(B3) formalize how the partitions produced by the algorithm behave as the finite population size grows.
 (B1) is a shrinking diameter assumption that mirrors the adaptive concentration result proved for i.i.d.
-forests ([94]).
+forests Wager2015-nc.
 (B2) requires that the thickness of the error band be negligible relative to the leaf’s own diameter.
-In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated if using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate.
-Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(x)\triangle L(x)$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(x))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively.
+In other words, the rate at which the leaf $L^{*}(\cdot)$ function estimated from the bootstrap sample approaches the leaf $L(\cdot)$ estimated when using the full finite population must be faster than the oracle leaf’s own shrinking-volume rate.
+Under the required relationship between $d_{N}$ and $k_{N}$, together with (A1) regularity conditions on the geometry, the symmetric-difference region $L^{*}(\mathbf{x})\triangle L(\mathbf{x})$ has measure $O_{p}(d_{N}/k_{N}^{p})$, which implies that both the tree estimator $\widehat{T}^{*}_{N}(L_{\xi}^{*}(\mathbf{x}))$ and its conditional variance coincide with their oracle counterparts up to $o_{p}(N^{-1/2})$ and $o_{p}(N^{-1})$, respectively.
 Among all the assumptions, (B2) is the hardest to directly control.
-Thus we provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
+Thus we have provided Theorem 3.2 that partially shed light on conditions under which (B2) can be reasonably assumed to hold.
 Corollary 3.5.
-Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ i.i.d.
+Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ are i.i.d.
 realizations from $P_{\Omega}(\cdot)$.
-Suppose $\mathbf{k}$ satisfying (K1)–(K3), and $\mathcal{Y}$ locally compact Polish (satisfying (S1)).
-Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ an estimate for it.
-Denote $P_{\omega}(\mathbf{x})$ the true conditional measure of $Y|X=\mathbf{x}$ in the super population space.
+Suppose $\mathbf{k}$ satisfies (K1)–(K3), and $\mathcal{Y}$ is locally compact Polish (satisfying (S1)).
+Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ is an estimate for it.
+Denote $P_{\Omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the super population space.
 Then for any specified $\mathbf{x}\in\mathcal{X}$ and $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),\;d_{\mathbf{k}})\to(\mathcal{V},d)$ a measurable functional continuous at $P^{N}(Y|X=\mathbf{x})$ for a.e.
@@
 Remark.
-Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{S^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law.
+Corollary 3.5 reduces design– and model–based consistency of $\widehat{\Psi}_{N}(\mathbf{x})$ to control of $\widehat{P}_{\mathcal{S}^{N}\mid\omega}(\mathbf{x})$ in the $d_{\mathbf{k}}$-metric, together with continuity of the functional $\Psi$ at the relevant target law.
 Many standard summaries satisfy this requirement under mild regularity.
@@
 The kernel mean element itself and its norm $\Psi(P)=\|\mu_{\mathbf{k}}(P)\|_{\mathcal{H}}$, and the MMD between the true conditional $P$ to a reference distribution $Q$, $d_{\mathbf{k}}(P,Q)$, are also examples.
+For higher-order continuous functionals such as covariances or variances of kernel embeddings, larger effective sample sizes may be needed to ensure adequate concentration of the estimator $\widehat{P}_{\mathcal{S}^{N}|\omega}$ in the $d_{\mathbf{k}}$-metric.
 3.4 Implementation Considerations
 We collect practical implementation considerations for SDRF, consolidating several points noted earlier.
-  • For the re–sampling, ([22]) showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ pseudo population “and” the re-sample.
+  • For the re–sampling, Conti2020-mc showed that asymptotically, generating one pseudo population and then drawing $B$ samples from it (equivalently, realizing $B$ times $n^{*}$ conditioning on the realized pseudo population) is asymptotically equivalent to realizing independently $B$ times both the pseudo population “and” the re-sample.
   • RBF–type kernel is a common class of kernels that is translation–invariant.
-Median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice ([35, 64]).
+The median-heuristic, i.e., specifying the bandwidth $\sigma>0$ to be the median of pairwise sampled distance, is a commonly adopted, and reasonable choice (Garreau2017-ro; matabuena2023hypothesis).
   • Hyperparameters governing the stopping criteria are introduced to help ensure regularity-type conditions (A1)–(A3).
 For finite samples, it may be preferable to use a moderate number of shallow trees to keep local sample weights well-controlled.
-Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section A.5.1).
-  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step [92] in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
-To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d.; described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies.
+Moreover, when the number of PSUs is large but PSU sizes are small, one may consider omitting second-stage resampling (Section S5.1, Supplementary Material).
+We note that the maximum node–level weight ratio $\max_{i\in\mathcal{I}(L)}\hat{w}_{i}/\min_{i\in\mathcal{I}(L)}\hat{w}_{i}$ where $\mathcal{I}(L)$ the subject index set for node $L$ is the only additional hyperparameter compared to the classical i.i.d.
+random forests for keeping the weights controlled.
+  • Because SDRF yields a kernel-type estimator, one could further post-process $\widehat{P}(Y\mid X=\mathbf{x})$ (or $\widehat{\Psi}(\mathbf{x})$) with an additional kernel-smoothing step verdinelli2021forest in $\mathbf{x}$ to reduce finite-sample variability and stabilize the resulting estimates, particularly when it is reasonable to assume that the underlying target varies smoothly in $\mathbf{x}$.
+To conclude, up to this point, we have illustrated the necessity of an appropriate method when data arise under sampling designs as opposed to being i.i.d., specifically when $Y\not\!\perp\!\!\!\perp Z\mid X$ and $\xi\not\!\perp\!\!\!\perp Z\mid X$ (Section S2, Supplementary Material); described in detail our proposed algorithm as a design-aware solution; proved both types of consistency for the resulting estimator under reasonable assumptions; discussed practical implementation-side considerations; and outlined specific example targets for which this algorithm applies.
 The algorithm is flexible in the choice of $\Psi$, the space of $Y$, as well as the type of design.
@@
   1. 1.
-    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF [21], an existing distributional random forest that does not incorporate survey weights in the estimation process.
+    Quantify gains in statistical accuracy and efficiency as the sample size increases and compare SDRF with DRF cevid_michel_meinshausen_bühlmann_2022, an existing distributional random forest that does not incorporate survey weights in the estimation process.
   2. 2.
@@
 Each unit is assigned independently to a stratum with equal probability:
-$\bigl(\mathds{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathds{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
+$\bigl(\mathbbm{1}\{\mathrm{Strat}_{i}=1\},\ldots,\mathbbm{1}\{\mathrm{Strat}_{i}=H\}\bigr)\sim\operatorname{Multinomial}\!\left(1;\tfrac{1}{H},\ldots,\tfrac{1}{H}\right).$
 For each stratum $h\in\{1,\dots,H\}$, let
@@
 We define measure-of-size as a function of $Z$: For each PSU, $\mathrm{PSU}_{hj}$,
-$M_{hj}\;=\;8\mathds{1}(Z_{hj}>0)+2\mathds{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
-the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.
+$M_{hj}\;=\;8\mathbbm{1}(Z_{hj}>0)+2\mathbbm{1}(\bar{Z}_{hj}\leq 0),\quad\text{where }\bar{Z}_{hj}\;=\;\frac{1}{N_{hj}}\sum_{i\in\mathrm{PSU}_{hj}}Z_{i}$
+is the PSU-level mean of $Z_{i}$, $N_{hj}=|\mathrm{PSU}_{hj}|$ the number of subjects within that PSU.
 First-stage sampling selects PSUs within each stratum $h$ with probabilities proportional to size.
@@
 4.2 Computation Details
-We run the SDRF and DRF ([21]) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$.
+We run the SDRF and DRF (cevid_michel_meinshausen_bühlmann_2022) algorithms with the above data generating process for each of the finite-population size $N\in\{4000,5000,15000,22500\}$ and the number of trees $B\in\{10,30,70,200\}$.
 For each of the scenario, 200 seeds were run.
@@
 For analytical evaluation, we report (a) the maximum mean discrepancy (MMD) between the estimated kernel mean embedding and the population kernel mean embedding for the conditional distribution on a hold-out set, and (b) the accuracy in the estimation for the functional
-$\Psi(\mathbf{x}=(x_{1},\dots,x_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
+$\Psi(\mathbf{x}=(\mathbf{x}_{1},\dots,\mathbf{x}_{p}))\;=\;\mathbb{E}\!\left[\,Y_{1}\mid X=\mathbf{x}_{1}\,\right],$
 in terms of mean squared error (MSE) of $\widehat{\Psi}(\mathbf{x}_{1})$.
@@
 The minimum node size is set to $\max\!\left\{\,20,\;\left\lceil\sqrt{n_{s}}\,\right\rceil\right\}$, where $n$ denotes the size of the training-sample.
-We cap effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights.
+We cap node–level effective sampling weights by $\lambda_{\max}\;=\;5.5\times\frac{\max_{i}w_{i}}{\min_{i}w_{i}}$, where $w_{i}$ are the survey weights.
 Although additional hyperparameter tuning could further improve performance, we keep the settings minimal because the untuned configurations already reveal the main trends across $N$ and $B$.
@@
 baseline and is a logical competitor).
-All code is implemented under R, version 4.4.2 [73].
+All code is implemented under R, version 4.4.2 R_cite.
 4.3 Results
-Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P(Y\mid X)$ and its estimate $\widehat{P}(Y\mid X)$ from SDRF.
+Table 1 reports the mean (SD) of the MMD between the true finite population conditional distribution $P_{S^{N}|\omega}(Y\mid X)$ and its estimate $\widehat{P}_{S^{N}|\omega}(Y\mid X)$ from SDRF.
 The mean MMD decreases as the finite population size $N$ increases, while increasing the number of trees $B$ primarily reduces the SD (stabilizing forest performance) with comparatively smaller improvements in the mean and diminishing returns beyond moderate $B$.
 Overall, these finite-sample patterns are consistent with the convergence behavior asserted in Theorem 3.4.
-Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}(Y\mid X),\;P^{N}(Y\mid X))$, where $\widehat{P}(Y\mid X)$ is the distribution estimate from SDRF, across finite population size $N$ and number of trees $B$.
-NN $B=10$ $B=30$ $B=70$ $B=200$
-4,000 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
-5,000 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
-15,000 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
-22,500 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
-Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for Survey-DRF (SDRF) and benchmark DRF algorithms as $N$ and the number of trees grow.
+Table 1: Mean (SD) of $100\times d_{\mathbf{k}}(\widehat{P}_{S^{N}|\omega}(Y\mid X),\;P_{S^{N}|\omega}(Y\mid X))$ across finite population size NN (hence different average sample sizes $\bar{n}_{s}$) and the number of trees $B$.
+NN $\bar{n}_{s}$ $B=10$ $B=30$ $B=70$ $B=200$
+4,000 320 19.60 (2.52) 18.60 (2.27) 18.30 (2.19) 18.30 (2.10)
+5,000 400 19.30 (2.49) 18.40 (2.14) 18.00 (2.14) 18.10 (2.07)
+15,000 1,200 17.30 (1.86) 16.80 (1.44) 16.70 (1.30) 16.80 (1.23)
+22,500 1,800 17.40 (1.52) 16.80 (1.23) 16.60 (1.16) 16.70 (1.08)
+Table 2: Root mean squared error (RMSE) of the conditional mean estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}]$ averaged over $X_{1}$ for SDRF and benchmark DRF algorithms as finite population size $N$ and the number of trees $B$ grow.
 NN $N_{\text{PSU},h}$ $\bar{n}_{s}$ Algorithm Number of Trees $B$
@@
 SDRF 2.16 1.83 1.74 1.74
-Figure 1: Pointwise mean square error (MSE) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}_{1}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
-Figure 2: Pointwise standard error (SD) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
 Table 2 and Figures 12 compare SDRF with the i.i.d.-based DRF for estimating the regression functional $\Psi(\mathbf{x}_{1})=\mathbb{E}[Y_{1}\mid X_{1}=\mathbf{x}_{1}]$.
@@
 Also, Figure 2 shows that pointwise variability decreases with $B$ for both methods and remains comparable between SDRF and DRF; notably, SDRF does not exhibit an appreciable variance inflation despite the additional randomness introduced by its design-aware resampling step.
+Figure 1: Pointwise mean square error (MSE) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}_{1}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
+Figure 2: Pointwise standard error (SD) of the estimator $\widehat{\mathbb{E}}[Y_{1}|X_{1}=\mathbf{x}]$ from SDRF and DRF. Each point represents an average over 200 seeds.
 5 Biomedical Public Health Case Study
@@
 civilian, noninstitutionalized population: primary sampling units (PSUs) are selected at the county level, followed by smaller geographic areas and then individuals.
-All analyzes performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx .
-More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g.
+All analyses performed incorporate survey weights and design features provided by the Centers for Disease Control and Prevention (CDC) [^1] https://wwwn.cdc.gov/nchs/nhanes/Default.aspx .
+More specifically, in this paper, we focus on the 2011–2012 NHANES cycle and apply an SDRF algorithm that incorporates the sampling mechanism of the NHANES study during bagging/subsampling and throughout tree construction (e.g.
 split selection and node estimation).
@@
 Our primary clinical goal is to quantify how the population distribution of this bivariate outcome varies between these characteristics, to support the detection of diabetes, the evaluation of glycemic-control, and healthcare planning through the trends observed in the subgroups of the population.
-A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) [2], which ignore specific individual characteristics.
+A secondary clinical objective is to evaluate whether the observed heterogeneity motivates individual diabetes diagnostic thresholds for FPG and HbA1c, in contrast to fixed cutoffs proposed by medical associations such as the American Diabetes Association (ADA) american20252, which ignore specific individual characteristics.
 For analytical purposes, a total of $N=2732$ individuals with and without diabetes were included in the analysis.
-Let us go to define mathematically the problem.
+We now articulate the analysis goal mathematically.
 Let the clinical outcome be $Y=(Y_{1},Y_{2})^{\top}$, where $Y_{1}$ denotes fasting plasma glucose (FPG) and $Y_{2}$ denotes HbA1c.
@@
 From an empirical perspective, we run SDRF and estimate the conditional distribution of the bivariate result $Y=(\mathrm{FPG},\mathrm{HbA1c})^{\top}$ given covariates $X=\mathbf{x}$.
-Specifically, we computed the estimated conditional CDF
+Specifically, we compute the estimated conditional CDF
 $\hat{F}(y\mid\mathbf{x})=\hat{P}^{N}_{\mathrm{HT}}\big(Y\leq y\mid X=\mathbf{x}\big),\qquad\text{for all }y\in\mathcal{Y},\ \mathbf{x}\in\mathcal{X},$
-where inequality $Y\leq y$ is understood componentwise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
+where inequality $Y\leq y$ is understood component–wise. For representation purposes, we evaluate $\hat{F}(\cdot\mid\mathbf{x})$ on a grid of covariate combinations:
 $\mathrm{Age}\in\{30,50,70\},\quad\mathrm{BMI}\in\{20,30,40\},\quad\mathrm{Sex}\in\{0,1\}.$
-Let $\hat{\omega}_{i}(x)$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $x\in\mathcal{X}$.
+Let $\hat{\omega}_{i}(\mathbf{x})$ denote the SDRF distribution-regression weights for observation $i\in[N]$ at covariate value $\mathbf{x}\in\mathcal{X}$.
 Using these weights, we compute the estimated conditional mean and covariance,
-$\hat{\mu}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(x)=\sum_{i=1}^{N}\hat{\omega}_{i}(x)\,\bigl(Y_{i}-\hat{\mu}(x)\bigr)\bigl(Y_{i}-\hat{\mu}(x)\bigr)^{\top}.$
-We then define a conditional Mahalanobis score (weighted) for any pair $(x,y)\in\mathcal{X}\times\mathcal{Y}$ by
-$S(x,y)=\bigl(y-\hat{\mu}(x)\bigr)^{\top}\hat{\Sigma}(x)^{-1}\bigl(y-\hat{\mu}(x)\bigr).$
-Let $\{(x_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(x_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
-In the spirit of [53] (for finite-dimensional Euclidean responses) and [57] (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $x\in\mathcal{X}$ as
-$\hat{C}^{\alpha}(x)=\bigl\{\,y\in\mathcal{Y}:\;S(x,y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
-Compared to the natural competitor [53], our approach is more flexible: it requires a single model, avoids user-chosen smoothing parameters due to the use of multiple local polynomial regression, and enables valid inference under the complex design of the NHANES survey.
+$\hat{\mu}(\mathbf{x})=\sum_{i=1}^{N}\hat{\omega}_{i}(\mathbf{x})\,Y_{i}\quad\text{and}\quad\hat{\Sigma}(\mathbf{x})=\sum_{i=1}^{N}\hat{\omega}_{i}(\mathbf{x})\,\bigl(Y_{i}-\hat{\mu}(\mathbf{x})\bigr)\bigl(Y_{i}-\hat{\mu}(\mathbf{x})\bigr)^{\top}.$
+We then define a conditional Mahalanobis score (weighted) for any pair $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$ by
+$S(\mathbf{x},y)=\bigl(y-\hat{\mu}(\mathbf{x})\bigr)^{\top}\hat{\Sigma}(\mathbf{x})^{-1}\bigl(y-\hat{\mu}(\mathbf{x})\bigr).$
+Let $\{(\mathbf{x}_{a},y_{a})\}_{a=1}^{N}$ denote the observed random sample and let $\hat{q}_{1-\alpha}$ be the empirical $(1-\alpha)$ quantile of the sample $\{S(\mathbf{x}_{a},y_{a})\}_{a=1}^{N}$ (using the survey weights).
+In the spirit of lado2021 (for finite-dimensional Euclidean responses) and lugosi2025conformal (for metric space outcomes), we define the tolerance region $(1-\alpha)$ at the covariate value $\mathbf{x}\in\mathcal{X}$ as
+$\hat{C}^{\alpha}(\mathbf{x})=\bigl\{\,y\in\mathcal{Y}:\;S(\mathbf{x},y)\leq\hat{q}_{1-\alpha}\,\bigr\}.$
+A natural competitor to SDRF in this setting is the method of Lado et al.
+(2021) lado2021.
+That approach fits multiple univariate local polynomial regressions to estimate the conditional mean, and then uses residual-based estimates to approximate the conditional covariance and construct a $(1-\alpha)$ tolerance region.
+Because it proceeds marginally via separate univariate fits, it is unlikely to be fully efficient for joint estimation.
+Moreover, the required selection of smoothing parameters is performed without an explicit global criterion, which can lead to suboptimal choices.
+Finally, the method does not incorporate survey weights during estimation.
+In contrast, the SDRF approach is more flexible: it requires only a single model, avoids the user-chosen smoothing parameters required by multiple local polynomial regressions, and enables valid inference under the complex design of the NHANES survey.
 Figures 3 and 4 show tolerance regions (in expectation) for $\alpha\in\{0.10,0.25,0.50,0.75,0.90\}$ in the covariate combinations described above (a larger $1-\alpha$ implies a wider region).
@@
 The colored lines indicate the diagnostic thresholds of the American Diabetes Association (ADA): $\mathrm{FPG}\geq 126\penalty 10000\ \mathrm{mg/dL}$ and $\mathrm{HbA1c}\geq 6.5\%$.
-We discuss the statistically supported variable by variable.
-The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observation and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels.
+We discuss the statistical results variable by variable.
+The estimated joint distributions for women and men are similar in shape; however, women exhibit more extreme observations and a higher prevalence of elevated fasting plasma glucose (FPG), especially at higher $\alpha\in(0,1)$ levels.
 With increasing age, both FPG and HbA1c shift upward, indicating poorer glycemic control at older ages; this effect is most pronounced and consistent near the $0.9$ quantile.
@@
 From these results, two implications follow: (i) glycemic control in the U.S.
-population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and political interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with [8] .
+population varies significantly between demographic and anthropometric individual characteristics, with consequences for public-health planning and potential policy interventions; and (ii) diagnostic cutoff points for diabetes based on FPG and HbA1c may warrant adjustment for at least age and possibly BMI, rather than relying on the same rules for all population in line with Berkovic2025 .
 From a modeling and analysis perspective, SDRF provides a principled framework to define reference values for specific subpopulations, reclassify individuals based on biomarker profiles, and forecast public–health risks leveraging the full conditional distribution.
-In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to less informative inform to perform the decision making.
+In contrast, approaches that model only the conditional mean can miss clinically important features—such as quantiles, dispersion, and tail behavior, leading to lower richness of information for decision making.
 Figure 3: SDRF–estimated tolerance regions (in expectation) for $\log(\mathrm{FPG})$ vs.
@@
 6 Discussion
-We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements [95].
-To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, filling an important methodological gap.
-From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs to estimate conditional distributions for multivariate outcomes.
+We propose a random forest tailored to complex survey designs, in which the bagging step explicitly mimics the sampling design (e.g., strata, primary sampling units, and weights), inspired by recent progress on bootstrap methods for complex designs with rate improvements wang2022bootstrap.
+To the best of our knowledge, random forests and other tree ensembles have not previously been adapted to complex survey data, thus our proposed approach fills an important methodological gap.
+From an estimation and applied data analysis perspective, we introduce what is, to our knowledge, the first general framework for complex designs for estimating conditional distributions for multivariate outcomes.
 Our approach leverages reproducing-kernel Hilbert spaces (RKHS) and distances between kernel mean embeddings to estimate the full conditional distribution.
-This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach [7, 67, 79, 37].
-SDRF framework provides a powerful and flexible tool for biomedical studies conducted under complex survey designs, as illustrated by the NHANES case study reported here, studies that are essential for optimizing public-health decisions.
+This representation enables straightforward derivation of population quantities (e.g., quantiles, ratios) and supports downstream tasks–including two-sample testing, selection of distributional variables, and semi-supervised clustering—via a forest-based kernel mean embedding approach benard2024mmd; matabuena2025kernel; scornet2016random; gretton2012kernel.
+The SDRF framework provides a powerful and flexible tool for research studies conducted under complex survey designs, as illustrated by the NHANES case study reported here.
+Such studies are essential for optimizing public-health decisions.
 Our machine learning algorithm for estimating the conditional distribution of multivariate outcomes represents a substantial modeling advance, enabling investigators to address clinically relevant questions in these cohorts with direct implications for health policy, as demonstrated in our case study illustration.
@@
 In particular, we will study the upward bias that can arise in ratio-type estimators when the denominator (here, the leaf size or subpopulation size) must be estimated because the true leaf sizes are not available from the finite population or census.
-Extending SDRF for outcome-dependent sampling is another direction.
-From a data-analytic perspective, we will develop algorithms to handle non-Euclidean random objects [58] that arise in digital health applications, and we will design methods to quantify uncertainty (for example, via confidence bands) within our random forest frameworks.
-It is also of applied interest to extend our algorithm to functional data [24], which can be naturally accommodated by using kernel RKHS embeddings to define split rules.
-There are recent regression methods for functional data in complex survey designs, but they focus only on linear functional response models and the conditional regression mean operator [52, 65].
+Extending SDRF for outcome-dependent sampling is another potentially interesting direction.
+From a data-analytic perspective, it is important to develop algorithms which can handle non-Euclidean random objects lugosi2025conformalknnpredictiveuncertainty that arise in digital health applications, and we will design methods to quantify uncertainty (for example, via confidence bands) within our random forest framework.
+It is also of applied interest to extend our algorithm to functional data crainiceanu2024functional, which can be naturally accommodated by using kernel RKHS embeddings to define split rules.
+While there are recent regression methods for functional data in complex survey designs, they focus only on linear functional response models and the conditional regression mean operator koffman2025functionscalarregressioncomplex; 10.1093/biostatistics/kxaf013.
+There is a need for further development in these directions.
+Data and Code Availability.
+The NHANES data analyzed in Section 5 are publicly available from the CDC at https://wwwn.cdc.gov/nchs/nhanes/Default.aspx.
+R code implementing SDRF and reproducing all results is available from the authors upon request.
+Supplementary Material
+S1 Notation Table
+Table S1: Summary of notation
+Symbol Description
+Basic variables
+$Z\in\mathcal{Z}\subseteq\mathbb{R}^{q}$ Design variables (used by the sampling design)
+$X\in\mathcal{X}\subseteq\mathbb{R}^{p}$ Covariates / features in Euclidean space
+$Y\in\mathcal{Y}$ Outcome of interest on $\mathcal{Y}$
+$D^{N}=\{(X_{i},Y_{i},Z_{i})\}_{i\in[N]}$ Finite Population
+Probability spaces, measures, and targets
+$(\Omega,\mathcal{A},P_{\Omega})$ Super-population probability space for $(Z,X,Y)$
+$(\mathcal{S}^{N},\sigma(\mathcal{S}^{N}),P_{\mathcal{S}^{N}\mid\omega})$ Design probability space
+$(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\Omega}\otimes P_{\mathcal{S}^{N}\mid\omega})$ Joint probability space
+$P(Y\mid X=\mathbf{x})$ or $\mu_{\mathcal{P}}(\mathbf{x})$ Super-population conditional law given $\mathbf{x}$
+$P_{S_{N}\mid\omega}(Y\mid X=\mathbf{x})$ or $\mu_{N}(\mathbf{x})$ Design conditional law given $\mathbf{x}$
+$\Psi(\cdot)$ A continuous functional of the conditional law that is of interest
+$\widehat{\Psi}_{N}(\cdot)$ Estimator from SDRF
+Design quantities
+$N\in\mathbb{Z}_{>0}$ Number of subjects in the finite population
+$\xi_{i}\in\{0,1\}$ Selection (inclusion) indicator for unit $i$
+$n_{s}=\sum_{i\in[N]}\xi_{i}$ Realized sample size
+$\pi_{i}=\mathbb{E}[\xi_{i}]$ First–order inclusion probability of unit $i$
+$\pi_{ij}=\mathbb{E}[\xi_{i}\xi_{j}]$ Second–order joint inclusion probability
+$\underline{\lambda},\bar{\lambda}$ Positive constants s.t. $\min_{i}\pi_{i}\geq\underline{\lambda}n_{s}/N$, $\max_{i}\pi_{i}\leq\bar{\lambda}n_{s}/N$
+$\mathbf{p}$ Sampling design transition function generating $\{\xi_{i}\}$
+$H$ Number of strata in a stratified design
+Design-aware bagging and honesty
+$n_{i}^{*}$ Resampled multiplier (count) for unit $i$ used to start a tree
+$D_{b}^{*}=(X,Y,n_{b}^{*}/\pi)$ Design-aware resampled dataset for tree $b$
+$\hat{w}_{i}=\frac{1}{\pi_{i}}n_{b,i}^{*}$ Tree-specific weight after resampling
+$\omega_{i}(\mathbf{x})$, $\widehat{\omega}_{i}(\mathbf{x})$ Final forest prediction weights at query $\mathbf{x}$ (averaged over trees)
+$B$ Number of trees in the forest
+$M$ Number of re-sample weights to average over to form data for one tree
+$D_{\text{est}},D_{\text{split}}$ Estimation part vs. splitting part of the sample at a node (honesty)
+Partitions, leaves, and sets
+$L(\cdot)$ (Oracle) partition map or leaf function on $\mathcal{X}$ if $D^{N}$ is used
+$L_{\xi}(\cdot)$ Partition map obtained from the whole sample induced by $\xi$
+$L^{*}_{\xi}(\cdot)$ or $L^{*}(\cdot)$ Partition map obtained from design-aware resampled data
+$\operatorname{diam}(L(\mathbf{x}))$ Diameter of a leaf $L(\mathbf{x})$ under the Euclidean distance
+$C_{L}(\theta),C_{R}(\theta)$ Left/right child region induced by a split $\theta=(j,t)$
+$P_{a}\subseteq\mathcal{X}$ Region on $\mathcal{X}$ (parent node) being split
+$N_{L},N_{R},N_{P_{a}}$ (Finite-population) sizes of left, right, and parent regions
+$\widehat{N}_{L},\widehat{N}_{R},\widehat{N}_{P_{a}}$ Sample / resample analogues used in split score
+$k_{N}\to 0$ Target (upper) diameter of oracle leaves (shrinking with $N$)
+$r_{N}=c_{1}k_{N}$ Radius defining an inner ball inside $L(x)$
+$\overline{B}_{r_{N}}(\mathbf{x})$ Closed balls in $\mathcal{X}$ centered at $\mathbf{x}$
+Kernel, RKHS, and MMD
+$\mathbf{k}(\cdot,\cdot)$ Reproducing kernel, assumed characteristic/universal and bounded
+$\mathcal{H}$ Reproducing kernel Hilbert space (induced by $\mathbf{k}$)
+$\mu_{\mathbf{k}}(P)$ Kernel mean embedding (KME) of some finite signed Radon measures $P$, an element in $\mathcal{H}$
+$d_{\mathbf{k}}(\cdot,\cdot)$ MMD distance between embedded distributions, induced by $\mathbf{k}$
+Convergence and order notation
+$O_{p_{d}}(\cdot),\,o_{p_{d}}(\cdot)$ Usual stochastic order w.r.t $P_{\mathcal{S}^{N}|\omega}$
+$X_{n}=\Theta_{p_{d}}(a_{n})$ $X_{n}=O_{p_{d}}(a_{n})$ and $a_{n}=O_{p_{d}}(|X_{n}|)$ w.r.t $P_{\mathcal{S}^{N}|\omega}$
+$a_{n}\lesssim b_{n}$ There exists $C>0$ with $a_{n}\leq Cb_{n}$ (i.e. $a_{n}=O(b_{n})$)
+$a_{n}\lesssim_{p}b_{n}$ Same statement in probability
+$X_{n}\xrightarrow{p_{d}}X$ Convergence in design probability space $(\mathcal{S},\sigma(\mathcal{S}),P_{\mathcal{S}^{N}|\omega})$
+(Design consistency)
+$X_{n}\xrightarrow{p_{\otimes}}X$ Convergence in the joint probability space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),\,P_{\omega}\otimes P_{\mathcal{S}^{N}|\omega})$
+(Model consistency)
+Hyperparameters for SDRF
+$q\in(0,1)$ Base probability to send a PSU to the “split” part
+$mtry\in\{1,\dots,p\}$ Number of candidate features considered at each split
+$\lambda_{\max}:=\max_{i\in L}\hat{w}_{i}/\min_{i\in L}\hat{w}_{i}$ The maximum within-leaf max-to-min weight ratio
+$n_{L,\min}\in\{1,\dots,p/2\}$ Minimum number of (un-weighted) observations per leaf
+We use $O_{p}$ notation in the usual sense and $X_{n}=\Theta_{p}(a_{n})$ if and only if $X_{n}=O_{p}(a_{n})$ and $a_{n}=O_{p}(|X_{n}|)$.
+We use $a_{n}\lesssim b_{n}$ to denote that there exists $C>0$ such that $a_{n}\leq Cb_{n}$.
+In other words, $a_{n}\lesssim b_{n}\iff a_{n}=O(b_{n})$; and similarly $a_{n}\lesssim_{p}b_{n}\iff a_{n}=O_{p}(b_{n})$.
+We denote the index set $a]=\{1,\dots,a\}$ for any positive integer $a\in\mathbb{Z}_{>0}$.
+We use $X_{n}\xrightarrow{p_{d}}X$ to denote convergence in probability with respect to all sources of randomness conditional on the fixed finite population of size $N$.
+Specifically, $p_{d}$ encompasses randomness from the sampling design (indicators $\{\xi_{i}\}$), bootstrap resampling (multipliers $\{n^{*}_{b,i}\}$), and algorithmic randomness (PSU-level splitting indicators $\{\eta_{k}\}$, random feature selection, and tree construction).
+We use $X_{n}\xrightarrow{p_{\otimes}}X$ to denote convergence in probability with respect to the joint measure $P_{\Omega\times\mathcal{S}^{N}}$ over the product space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}))$.
+This joint measure $p_{\otimes}$ includes all sources of randomness: super-population randomness (generating the finite population $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ from $P_{\Omega}$), sampling design, bootstrap resampling, and algorithmic randomness.
+The “*” superscript emphasizes dependency on resampled weights ${n}^{*}_{b,i}$.
+We write $n_{s}$ or $n_{s,N}$ interchangeably for the (possibly random) sample size, with the subscript $N$ emphasizing dependency on the population size.
+Under (stratified) two–stage cluster designs, let $h\in[H]$ index strata.
+In stratum $h$, there are $K_{h}$ primary sampling units (PSUs) indexed by $k\in[K_{h}]$, and PSU $(h,k)$ contains $M_{hk}$ secondary sampling units (SSUs) indexed by $j\in[M_{hk}]$.
+We therefore index an SSU by the triple $u=(h,k,j)$.
+Let $\xi^{(1)}_{hk}\in\{0,1\}$ denote the stage–1 inclusion indicator for PSU $(h,k)$, and, conditional on $\xi^{(1)}_{hk}=1$, let $\xi^{(2)}_{hkj}\in\{0,1\}$ denote the stage–2 inclusion indicator for SSU $(h,k,j)$ within PSU $(h,k)$.
+The overall SSU inclusion indicator is $\xi_{u}:=\xi^{(1)}_{hk}\xi^{(2)}_{hkj}$.
+Define the corresponding inclusion probabilities $\pi^{(1)}_{hk}:=\Pr(\xi^{(1)}_{hk}=1)$, $\pi^{(2)}_{hkj\mid hk}:=\Pr(\xi^{(2)}_{hkj}=1\mid\xi^{(1)}_{hk}=1)$, and $\pi_{u}:=\Pr(\xi_{u}=1)=\pi^{(1)}_{hk}\pi^{(2)}_{hkj\mid hk}$.
+For distinct SSUs $u\neq v$, write the joint inclusion probability as $\pi_{uv}:=\Pr(\xi_{u}=\xi_{v}=1)$.
+Finally, let $k(u):=(h,k)$ denote the PSU containing SSU $u$, so that $k(u)=k(v)$ iff $u$ and $v$ belong to the same PSU.
+S2 Motivating Example
+When learning a conditional distribution, using design-aware algorithms is not only about potential efficiency gains; it is essential in certain settings.
+Consider the following scenario.
+Denote by $Z\in\mathcal{Z}$ a design variable, $\xi\in\{0,1\}$ the inclusion indicator, and suppose that we are interested in $P(Y\mid X)$.
+Design variables may be available only on the sampling frame and used to construct strata, define measures of size (MOS) for PPS sampling, or model nonresponse.
+Examples include fine geographic indicators such as census blocks, facility size measures such as the number of providers, or other administrative flags.
+Although $X$ and $Z$ may overlap, $Z$ often contains frame-only or confidential information and therefore must be removed from the released sampled data $D=\{(X_{i},Y_{i},w_{i}):i=1,\dots,N;\xi_{i}=1\}$ used for analysis.
+Problem arise when $Z$ indexes subgroups with heterogeneous conditional outcome distributions and the design differentially samples these subgroups.
+Formally, when
+$\begin{split}Y\not\!\perp\!\!\!\perp Z\mid X\quad\text{and}\quad\xi\not\!\perp\!\!\!\perp Z\mid X.\end{split}$ (S1)
+Even if the design is conditionally noninformative given $(X,Z)$, i.e., $Y\perp\!\!\!\perp\xi\mid X,Z$, we have
+$P(Y\mid X,\xi=1)=\int P(Y\mid X,Z=z)\,P(dz\mid X,\xi=1)\neq P(Y\mid X),\text{ since }$
+$P(dz\mid X,\xi=1)=\frac{P(\xi=1\mid X,Z=z)}{\int_{\mathcal{Z}}P(\xi=1\mid X,Z=z)\,P(dz\mid X)}\,P(dz\mid X)\neq P(dz\mid X).$
+Thus, algorithms designed for i.i.d.
+data that implicitly learn $P(Y\mid X,\,\xi=1)$ lead to systematic bias relative to $P(Y\mid X)$, which does not vanish with increasing sample size or model complexity.
+This is the case even for non-parametric algorithms that free us from modeling error.
+Consistency for $P(Y\mid X)$ is impossible without incorporating the design.
+This phenomenon is showcased in Section 4 of the manuscript.
+S3 Survey Design
+Proof of Lemma 2.1.
+By (D3),
+$\sum_{i=1}^{N}\mathrm{Var}_{S^{N}\mid\omega}\left(\frac{\xi_{i}}{\pi_{i}}-1\right)\leq\sum_{i=1}^{N}\frac{1}{\pi_{i}}\leq N\cdot\frac{1}{\inf_{i}\pi_{i}}\leq N\cdot\frac{1}{\underline{\lambda}n/N}=\frac{N^{2}}{\underline{\lambda}n}.$
+By (D4) and (D3), for $i\neq j$,
+$\left|\mathrm{Cov}_{S^{N}\mid\omega}(\frac{\xi_{i}}{\pi_{i}}-1,\frac{\xi_{j}}{\pi_{j}}-1)\right|\;\leq\;\frac{1}{\pi_{i}\pi_{j}}\cdot\frac{C_{1}}{n_{s,N}}\;\leq\;\frac{C_{1}}{n_{s,N}}\cdot\frac{1}{(\inf_{i}\pi_{i})^{2}}\;\leq\;\frac{C_{1}}{n_{s,N}}\cdot\frac{1}{(\underline{\lambda}n/N)^{2}},$
+for some $C_{1}<\infty$.
+So summing over $\binom{N}{2}$ pairs gives $\sum_{1\leq i<j\leq N}\left|\mathrm{Cov}_{S^{N}\mid\omega}(\frac{\xi_{i}}{\pi_{i}}-1,\;\frac{\xi_{j}}{\pi_{j}}-1)\right|\;\leq\;\frac{N(N-1)}{2}\cdot\frac{C_{1}N^{2}}{\underline{\lambda}^{2}n^{2}}\cdot\frac{1}{n_{s,N}}.$ Thus
+$Var_{\mathcal{S}^{N}\mid\omega}\left(\sum_{i\in[N]}\frac{\xi_{i}}{\pi_{i}}-1\right)\lesssim\frac{N^{2}}{n}+N(N-1)\frac{N^{2}}{n^{2}n_{s,N}}.$
+By (D2), $n_{s,N}=\Theta_{p_{d}}(N)$, and $n=\Theta(N)$.
+By Chebyshev’s inequality and that $\mathbb{E}[\sum_{i\in[N]}(\frac{\xi_{i}}{\pi_{i}}-1)]=0$, for each $\epsilon>0$, $P_{\mathcal{S}^{N}\mid\omega}\!\left(|\frac{1}{N}\sum_{i\in[N]}(\frac{\xi_{i}}{\pi_{i}}-1)|>\epsilon\right)=O(1/N)\to 0,$ proving the LLN statement; and $P_{\mathcal{S}^{N}\mid\omega}\!\left(|\frac{1}{\sqrt{N}}\sum_{i\in[N]}(\frac{\xi_{i}}{\pi_{i}}-1)|>\epsilon\right)=O_{p_{d}}(1)$ uniformly in $N$, proving the CLT statement.
+∎
+Lemma S1.
+Each of the following first–stage designs satisfies assumptions (D2)(D4):
+  1. (a)
+    Poisson sampling. Conditionally independent indicators $\{\xi_{i}\}_{i=1}^{N}$ given $(X^{N},Z^{N})$ with $\xi_{i}\sim\mathrm{Bern}(\pi_{i})$, where the inclusion probabilities satisfy
+    $0<\pi_{\min}\leq\pi_{i}\leq\pi_{\max}<1\qquad\text{for all }i,N,$
+    and the expected sample size $n=\sum_{i=1}^{N}\pi_{i}$ satisfies $n/N\to f\in(0,1)$.
+  2. (b)
+    SRSWOR (fixed size). Simple random sampling without replacement of a fixed size $n_{N}$ from $[N]$, where $n_{N}/N\to f\in(0,1)$ as $N\to\infty$.
+  3. (c)
+    PPS with replacement (distinct–unit sample).
+Let MOS values $\{m_{i}\}_{i=1}^{N}$ satisfy $0<m_{\min}\leq m_{i}\leq m_{\max}<\infty$ for all $i,N$.
+Define per–draw probabilities $p_{i}:=m_{i}/\sum_{r=1}^{N}m_{r}$.
+Draw $m_{N}$ times with replacement, independently across draws, with $\Pr(I_{t}=i)=p_{i}$, and define distinct–unit membership indicators
+    $\xi_{i}:=\mathbbm{1}\{i\in\{I_{1},\dots,I_{m_{N}}\}\},\qquad n_{s,N}:=\sum_{i=1}^{N}\xi_{i}.$
+    Assume $m_{N}/N\to f_{0}\in(0,1)$ and that the expected number of distinct sampled units $n:=\mathbb{E}(n_{s,N})=\sum_{i=1}^{N}\pi_{i}$ satisfies $n/N\to f\in(0,1)$.
+Proof of Lemma S1.
+We verify (D2)–(D4) case by case.
+Poisson sampling.
+Let $n_{s,N}=\sum_{i=1}^{N}\xi_{i}$ with independent $\xi_{i}\sim\mathrm{Bern}(\pi_{i})$ and $n=\mathbb{E}(n_{s,N})=\sum_{i=1}^{N}\pi_{i}$.
+Since $n/N\to f\in(0,1)$, fix any $\delta\in(0,f)$ and set $\eta:=1-\delta/f\in(0,1)$.
+For all large $N$, $n\geq(f/2)N$ and $\delta N\leq(1-\eta/2)n$.
+A standard Chernoff bound for Poisson–binomial sums yields
+$\Pr\!\bigl(n_{s,N}\leq(1-\eta/2)n\bigr)\leq\exp\!\left(-\frac{\eta^{2}}{8}\,n\right).$
+Consequently, for all large $N$,
+$\Pr\!\left(\frac{n_{s,N}}{N}\leq\delta\right)\leq\Pr\!\bigl(n_{s,N}\leq(1-\eta/2)n\bigr)\leq\exp(-cN)$
+for some $c>0$.
+Hence, for any fixed $\kappa>4$, there exists $C<\infty$ such that $\Pr(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$ for all large $N$, proving (D2).
+Moreover, $n_{s,N}/N\to_{p_{d}}f$ by the weak law of large numbers for independent bounded summands.
+To verify (D3), note that $\pi_{\min}\leq\pi_{i}\leq\pi_{\max}$ for all $i,N$ implies
+$\pi_{\min}\leq\frac{n}{N}=\frac{1}{N}\sum_{i=1}^{N}\pi_{i}\leq\pi_{\max}.$
+Therefore, for all $i,N$,
+$\pi_{i}\leq\pi_{\max}\leq\frac{\pi_{\max}}{\pi_{\min}}\cdot\frac{n}{N},\qquad\pi_{i}\geq\pi_{\min}\geq\frac{\pi_{\min}}{\pi_{\max}}\cdot\frac{n}{N},$
+so (D3) holds with $\overline{\lambda}=\pi_{\max}/\pi_{\min}$ and $\underline{\lambda}=\pi_{\min}/\pi_{\max}$.
+Finally, independence gives $\pi_{ij}=\pi_{i}\pi_{j}$ for all $i\neq j$, hence (D4) holds with $C_{1}=0$.
+SRSWOR. Here $n_{s,N}\equiv n_{N}$ deterministically, so $n_{s,N}/N\to f$ and, for any $\delta\in(0,f)$, $\Pr(n_{s,N}/N\leq\delta)=0$ for all large $N$, proving (D2) with any $\kappa>0$.
+Moreover $\pi_{i}=n_{N}/N$ for all $i$, so (D3) holds with $\underline{\lambda}=\overline{\lambda}=1$ and $n=\mathbb{E}(n_{s,N})=n_{N}$.
+For $i\neq j$,
+$\pi_{ij}=\frac{n_{N}(n_{N}-1)}{N(N-1)},\qquad\pi_{i}\pi_{j}=\Big(\frac{n_{N}}{N}\Big)^{2},$
+hence
+$\pi_{ij}-\pi_{i}\pi_{j}=-\frac{n_{N}(N-n_{N})}{N^{2}(N-1)}.$
+Therefore
+$\max_{1\leq i<j\leq N}|\pi_{ij}-\pi_{i}\pi_{j}|=\frac{n_{N}(N-n_{N})}{N^{2}(N-1)}\leq\frac{n_{N}}{N(N-1)}.$
+Since $n_{N}/N\to f\in(0,1)$, there exists $N_{0}$ and $c>0$ such that $n_{N}\geq cN$ for all $N\geq N_{0}$. Thus, for $N\geq N_{0}$,
+$\frac{n_{N}}{N(N-1)}\leq\frac{1}{c(N-1)}\leq\frac{2}{c}\cdot\frac{1}{n_{N}}=\frac{2}{c}\cdot\frac{1}{n_{s,N}},$
+which is (D4) with $C_{1}=2/c$.
+PPS with replacement (distinct–unit sample). Define $p_{i}=m_{i}/\sum_{r=1}^{N}m_{r}$ and note that
+$\frac{m_{\min}}{Nm_{\max}}\leq p_{i}\leq\frac{m_{\max}}{Nm_{\min}}\qquad\text{for all }i,N.$
+Let $\pi_{i}=1-(1-p_{i})^{m_{N}}$ and $n_{s,N}=\sum_{i=1}^{N}\xi_{i}$ with $\xi_{i}=\mathbbm{1}\{i\in\{I_{1},\dots,I_{m_{N}}\}\}$.
+By assumption, $m_{N}/N\to f_{0}\in(0,1)$ and $n/N\to f\in(0,1)$, where $n=\mathbb{E}(n_{s,N})=\sum_{i=1}^{N}\pi_{i}$.
+(D2).
+Consider the function $F(I_{1},\dots,I_{m_{N}})=n_{s,N}$.
+Changing one draw $I_{t}$ can change $n_{s,N}$ by at most $1$, hence $F$ satisfies bounded differences with constants $c_{t}\equiv 1$.
+McDiarmid’s inequality yields, for all $t>0$,
+$\Pr\!\left(n_{s,N}\leq\mathbb{E}(n_{s,N})-t\right)\leq\exp\!\left(-\frac{2t^{2}}{m_{N}}\right).$
+Fix $\delta\in(0,f)$ and choose $\varepsilon:=(f-\delta)/2>0$.
+For all large $N$, $\mathbb{E}(n_{s,N})/N=n/N\geq f-\varepsilon=\delta+\varepsilon$, hence $\mathbb{E}(n_{s,N})\geq(\delta+\varepsilon)N$.
+Rearranging, we have $\delta N\leq\mathbb{E}(n_{s,N})-\varepsilon N$ for all large $N$.
+Therefore, for all large $N$,
+$\Pr\!\left(\frac{n_{s,N}}{N}\leq\delta\right)\leq\Pr\!\left(n_{s,N}\leq\delta N\right)\leq\Pr\!\left(n_{s,N}\leq\mathbb{E}(n_{s,N})-\varepsilon N\right)\leq\exp\!\left(-\frac{2\varepsilon^{2}N^{2}}{m_{N}}\right)\leq\exp(-cN)$
+for some $c>0$ (since $m_{N}\asymp N$).
+As before, exponential decay implies $\Pr(n_{s,N}/N\leq\delta)\leq CN^{-\kappa}$ for any $\kappa>4$ with suitable $C<\infty$, proving (D2).
+Moreover $n_{s,N}/N\to_{p_{d}}f$ by the same concentration.
+(D3). Define
+$a_{N}:=\min_{i\in[N]}\pi_{i},\qquad b_{N}:=\max_{i\in[N]}\pi_{i}.$
+Since $\pi_{i}$ is increasing in $p_{i}$ and $p_{i}\in[c_{1}/N,c_{2}/N]$ with $c_{1}:=m_{\min}/m_{\max}$ and $c_{2}:=m_{\max}/m_{\min}$, we have
+$1-\Big(1-\frac{c_{1}}{N}\Big)^{m_{N}}\leq a_{N}\leq\pi_{i}\leq b_{N}\leq 1-\Big(1-\frac{c_{2}}{N}\Big)^{m_{N}}.$
+Also $n/N$ is the average of the $\pi_{i}$, hence $a_{N}\leq n/N\leq b_{N}$. Therefore, for all $i$,
+$\pi_{i}\leq b_{N}=\frac{b_{N}}{a_{N}}\,a_{N}\leq\frac{b_{N}}{a_{N}}\,\frac{n}{N},\qquad\pi_{i}\geq a_{N}=\frac{a_{N}}{b_{N}}\,b_{N}\geq\frac{a_{N}}{b_{N}}\,\frac{n}{N}.$
+Since $m_{N}/N\to f_{0}\in(0,1)$, we have $a_{N}\to 1-e^{-f_{0}c_{1}}$ and $b_{N}\to 1-e^{-f_{0}c_{2}}$, so $b_{N}/a_{N}$ and $a_{N}/b_{N}$ are bounded for all large $N$.
+Hence there exist constants $0<\underline{\lambda}\leq\overline{\lambda}<\infty$ such that (D3) holds for all large $N$.
+(D4). For $i\neq j$,
+$\pi_{ij}=1-(1-p_{i})^{m_{N}}-(1-p_{j})^{m_{N}}+(1-p_{i}-p_{j})^{m_{N}},$
+and
+$\pi_{ij}-\pi_{i}\pi_{j}=(1-p_{i}-p_{j})^{m_{N}}-\bigl(1-p_{i}-p_{j}+p_{i}p_{j}\bigr)^{m_{N}}.$
+Let $a:=1-p_{i}-p_{j}$ and $b:=a+p_{i}p_{j}$, so $a,b\in[0,1]$. By the mean value theorem and the bound $\sup_{x\in[0,1]}|(x^{m_{N}})^{\prime}|=m_{N}$,
+$|\pi_{ij}-\pi_{i}\pi_{j}|=|a^{m_{N}}-b^{m_{N}}|\leq m_{N}|a-b|=m_{N}p_{i}p_{j}\leq m_{N}\left(\frac{c_{2}}{N}\right)^{2}\leq\frac{C}{N}$
+for a constant $C<\infty$ and all $N$ (since $m_{N}/N$ is bounded).
+Because $n_{s,N}\leq m_{N}$ almost surely (the number of distinct sampled units cannot exceed the number of draws), it follows that $1/m_{N}\leq 1/n_{s,N}$ a.s., hence for all large $N$,
+$\max_{1\leq i<j\leq N}|\pi_{ij}-\pi_{i}\pi_{j}|\leq\frac{C^{\prime}}{m_{N}}\leq\frac{C^{\prime}}{n_{s,N}}\qquad\text{a.s.}$
+for some $C^{\prime}<\infty$, which proves (D4). ∎
+We now verify through the Lemma S2 that a broad class of stratified two-stage cluster designs induces SSU-level inclusion probabilities satisfying (D2)–(D4).
+Let the SSU index be $u=(h,k,j)$, where $h\in[H]$ indexes strata, $k\in[K_{h}]$ indexes PSUs in stratum $h$, and $j\in[\mathcal{U}_{hk}]$ indexes SSUs within PSU $(h,k)$.
+Write $\xi^{(1)}_{hk}\in\{0,1\}$ for the stage–1 PSU inclusion indicator and $\xi^{(2)}_{hkj}\in\{0,1\}$ for the stage–2 SSU inclusion indicator, defined conditionally on $\xi^{(1)}_{hk}=1$.
+Let the overall SSU indicator be $\xi_{u}:=\xi^{(1)}_{hk}\xi^{(2)}_{hkj}$.
+Define the marginal inclusion probabilities $\pi^{(1)}_{hk}:=\Pr(\xi^{(1)}_{hk}=1)$, $\pi^{(2)}_{hkj\mid hk}:=\Pr(\xi^{(2)}_{hkj}=1\mid\xi^{(1)}_{hk}=1)$, and $\pi_{u}:=\Pr(\xi_{u}=1)=\pi^{(1)}_{hk}\pi^{(2)}_{hkj\mid hk}$.
+For $u\neq v$, let $\pi_{uv}:=\Pr(\xi_{u}=\xi_{v}=1)$.
+Lemma S2 (Stage-wise condition ensure SSU-level (D2)–(D4)).
+For this lemma, let $N:=\sum_{h=1}^{H}\sum_{k=1}^{K_{h}}M_{hk}$ be the SSU population size.
+Let $n_{s}:=\sum_{h=1}^{H}\sum_{k=1}^{K_{h}}\sum_{j=1}^{\mathcal{U}_{hk}}\xi_{u}$ be the possibly random SSU sample size, and let $n:=\mathbb{E}_{P_{\mathcal{S}^{N}\mid\omega}}(n_{s})$.
+Suppose the stratified two-stage clustering design satisfy
+  1. (D-TS1)
+    Across-stratum independence.
+The design is independent across strata: for any measurable functions of the stage–1 and stage–2 indicators, the collections $\{\xi^{(1)}_{hk},\xi^{(2)}_{hkj}:k\in[K_{h}],j\in[M_{hk}]\}$ are mutually independent over $h\in[H]$.
+  2. (D-TS2)
+    Stable overall sampling fraction.
+There exists $f\in(0,1]$ such that $n_{s}/N\xrightarrow{p_{d}}f$, and there exist constants $\delta\in(0,1)$, $\kappa>4$ and $C<\infty$ such that $P_{\mathcal{S}^{N}\mid\omega}(n_{s}/N\leq\delta)\leq CN^{-\kappa}$ for all $N\geq 1$.
+  3. (D-TS3)
+    Stagewise first-order inclusion bounds.
+There exist constants $0<\underline{\lambda}_{1}\leq\bar{\lambda}_{1}<\infty$ and $0<\underline{\lambda}_{2}\leq\bar{\lambda}_{2}<\infty$ such that for all $h\in[H]$, $k\in[K_{h}]$, $j\in[M_{hk}]$,
+    $\underline{\lambda}_{1}\,\frac{n}{\sum_{h^{\prime}=1}^{H}K_{h^{\prime}}}\ \leq\ \pi^{(1)}_{hk}\ \leq\ \bar{\lambda}_{1}\,\frac{n}{\sum_{h^{\prime}=1}^{H}K_{h^{\prime}}},\qquad\underline{\lambda}_{2}\,\frac{n}{N}\ \leq\ \pi^{(2)}_{hkj\mid hk}\ \leq\ \bar{\lambda}_{2}\,\frac{n}{N}.$
+    (Any equivalent bounds implying $0<\inf_{u}\pi_{u}\lesssim n/N\lesssim\sup_{u}\pi_{u}<\infty$ are admissible.)
+  4. (D-TS4)
+    Stagewise controlled second order. There exists $C<\infty$ such that for all $N\geq 1$:
+    1. (a)
+      Between-PSU dependence within strata: for each $h\in[H]$ and all $k\neq k^{\prime}\in[K_{h}]$,
+      $\bigl|\pi^{(1)}_{hkk^{\prime}}-\pi^{(1)}_{hk}\pi^{(1)}_{hk^{\prime}}\bigr|\ \leq\ \frac{C}{n_{s}}.$
+    2. (b)
+      Within-PSU dependence: for each $h\in[H]$, $k\in[K_{h}]$, and all $j\neq j^{\prime}\in[M_{hk}]$,
+      $\bigl|\pi^{(2)}_{hkj,hkj^{\prime}\mid hk}-\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hkj^{\prime}\mid hk}\bigr|\ \leq\ \frac{C}{n_{s}}.$
+Then the induced SSU-level inclusion probabilities $\{\pi_{u}\}_{u=1}^{N}$ and joint probabilities $\{\pi_{uv}\}_{1\leq u<v\leq N}$ satisfy (D2)–(D4) (with the same $n_{s}$, $N$ and $n$).
+In particular, there exist constants $0<\underline{\lambda}\leq\bar{\lambda}<\infty$ such that for all $N\geq 1$, $\inf_{u\in[N]}\pi_{u}\ \geq\ \underline{\lambda}\,\frac{n}{N},\,\sup_{u\in[N]}\pi_{u}\ \leq\ \bar{\lambda}\,\frac{n}{N},$ and there exists $C^{\prime}<\infty$ such that $\max_{1\leq u<v\leq N}\bigl|\pi_{uv}-\pi_{u}\pi_{v}\bigr|\ \leq\ \frac{C^{\prime}}{n_{s}}\quad\text{a.s.}$
+Proof of Lemma S2.
+We verify (D2)–(D4) in turn.
+(D2). directly follows from (D-TS2).
+(D3). Fix $u=(h,k,j)$. By definition, $\pi_{u}=\pi^{(1)}_{hk}\pi^{(2)}_{hkj\mid hk}$. Assumption (D-TS3) yields
+$\lambda_{1}\lambda_{2}\Bigl(\frac{n}{\sum_{h^{\prime}}K_{h^{\prime}}}\Bigr)\Bigl(\frac{n}{N}\Bigr)\ \leq\ \pi_{u}\ \leq\ \bar{\lambda}_{1}\bar{\lambda}_{2}\Bigl(\frac{n}{\sum_{h^{\prime}}K_{h^{\prime}}}\Bigr)\Bigl(\frac{n}{N}\Bigr).$
+Since $\sum_{h^{\prime}}K_{h^{\prime}}\leq N$ and $n\asymp N$ under (D2), we may absorb the factor $n/(\sum_{h^{\prime}}K_{h^{\prime}})$ into the constants to obtain $\underline{\lambda}(n/N)\leq\pi_{u}\leq\bar{\lambda}(n/N)$ uniformly over $u\in[N]$.
+(D4). Fix distinct SSUs $u=(h,k,j)$ and $v=(h^{\prime},k^{\prime},j^{\prime})$ and consider cases.
+Case 1: $h\neq h^{\prime}$. By (D-TS1), strata are sampled independently, hence $\pi_{uv}=\pi_{u}\pi_{v}$.
+Case 2: $h=h^{\prime}$ and $k\neq k^{\prime}$. Then $u$ and $v$ belong to different PSUs within the same stratum. By the two-stage construction,
+$\pi_{uv}=\pi^{(1)}_{hkk^{\prime}}\,\pi^{(2)}_{hkj\mid hk}\,\pi^{(2)}_{hk^{\prime}j^{\prime}\mid hk^{\prime}}$.
+Therefore,
+$\pi_{uv}-\pi_{u}\pi_{v}=\bigl(\pi^{(1)}_{hkk^{\prime}}-\pi^{(1)}_{hk}\pi^{(1)}_{hk^{\prime}}\bigr)\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hk^{\prime}j^{\prime}\mid hk^{\prime}}$.
+Using (D-TS4), part (a) and the uniform upper bounds in (D-TS3) gives $|\pi_{uv}-\pi_{u}\pi_{v}|\leq C/n_{s}$ for some $C<\infty$.
+Case 3: $h=h^{\prime}$ and $k=k^{\prime}$ and $j\neq j^{\prime}$. Then $u$ and $v$ lie in the same PSU. We have
+$\pi_{uv}=\pi^{(1)}_{hk}\,\pi^{(2)}_{hkj,hkj^{\prime}\mid hk},\qquad\pi_{u}\pi_{v}=(\pi^{(1)}_{hk})^{2}\,\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hkj^{\prime}\mid hk}$.
+Hence
+$\pi_{uv}-\pi_{u}\pi_{v}$ $=\pi^{(1)}_{hk}\Bigl(\pi^{(2)}_{hkj,hkj^{\prime}\mid hk}-\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hkj^{\prime}\mid hk}\Bigr)$
+$\quad+\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hkj^{\prime}\mid hk}\Bigl(\pi^{(1)}_{hk}-(\pi^{(1)}_{hk})^{2}\Bigr).$
+The first term is bounded by $C/n_{s}$ by (D-TS4), part (b) and the uniform bound $\pi^{(1)}_{hk}\leq 1$.
+For the second term, note that $|\pi^{(1)}_{hk}-(\pi^{(1)}_{hk})^{2}|=\pi^{(1)}_{hk}(1-\pi^{(1)}_{hk})\leq\pi^{(1)}_{hk}$, so (D-TS3) yields
+$\pi^{(2)}_{hkj\mid hk}\pi^{(2)}_{hkj^{\prime}\mid hk}\bigl|\pi^{(1)}_{hk}-(\pi^{(1)}_{hk})^{2}\bigr|\ \leq\ (\bar{\lambda}_{2}n/N)^{2}\,(\bar{\lambda}_{1}n/\sum_{h^{\prime}}K_{h^{\prime}})\ \lesssim\ \frac{1}{n_{s}},$
+where the final inequality uses $n_{s}\asymp n\asymp N$ from (D2) and absorbs constants.
+Combining the bounds gives $|\pi_{uv}-\pi_{u}\pi_{v}|\leq C^{\prime}/n_{s}$ uniformly over $u\neq v$.
+This establishes (D4) and completes the proof.
+∎
+S4 Embedding of Measures
+We first show the applicability of the kernel mean embedding (KME) framework on measures in survey sampling.
+For any finite signed measure $\nu\in\mathcal{M}_{b}(\mathcal{Y})$ with finite total variation, recall the kernel mean embedding definition
+$\mu_{\mathbf{k}}(\nu):=\int\mathbf{k}(\cdot,t)\,d\nu(t)\in\mathcal{H}.$ (S2)
+The integral is well-defined since $\int\|\mathbf{k}(\cdot,t)\|_{\mathcal{H}}\,d|\nu|(t)\leq K\,|\nu|(\mathcal{Y})<\infty$. Moreover, for any $f\in\mathcal{H}$,
+$\langle\mu_{\mathbf{k}}(\nu),f\rangle_{\mathcal{H}}=\int\langle\mathbf{k}(\cdot,t),f\rangle_{\mathcal{H}}\,d\nu(t)=\int f(t)\,d\nu(t),$ (S3)
+where the last equality uses the reproducing property.
+The takeaways are that for any given $\mathbf{x}$ and a region $A(\mathbf{x})\subseteq\mathcal{X}$ indexed by $\mathbf{x}$ with $\{|A(\mathbf{x})|>0\}$:
+  1. 1.
+    Under (K1), each of $P_{Y\mid X\in A(\mathbf{x})}$, the true conditional measure in the super population, $P^{N}_{Y\mid X\in A(\mathbf{x})}$, the true conditional measure in the finite population, and $\widehat{P}^{\mathrm{HT}}_{Y\mid X\in A(\mathbf{x})}$ and $\widehat{P}^{\mathrm{HJ}}_{Y\mid X\in A(\mathbf{x})}$, its estimators, is a finite–signed measure in $\mathcal{M}_{b}(\mathcal{Y})$.
+  2. 2.
+    For any finite collection $\{(a_{i},y_{i})\}$ with $\sum_{i}|a_{i}|<\infty$,
+    $\mu_{\mathbf{k}}\!\Big(\sum_{i}a_{i}\delta_{y_{i}}\Big)=\sum_{i}a_{i}\,\mathbf{k}(\cdot,y_{i}),\qquad\langle\mu_{\mathbf{k}}(\sum_{i}a_{i}\delta_{y_{i}}),f\rangle_{\mathcal{H}}=\sum_{i}a_{i}f(y_{i}).$
+    Thus, HT/Hájek properties transfer to $\mathcal{H}$-valued objects by linearity, as the additional randomness enters through from the weights.
+  • Dirac measure (at a fixed point). For $y\in\mathcal{Y}$, the Dirac measure $\delta_{y}$ satisfies
+    $\mu_{\mathbf{k}}(\delta_{y})=\int\mathbf{k}(\cdot,t)\,d\delta_{y}(t)=\mathbf{k}(\cdot,y)\in\mathcal{H}.$ (S4)
+    Consequently, for any $f\in\mathcal{H}$,
+    $\langle\mu_{\mathbf{k}}(\delta_{y}),f\rangle_{\mathcal{H}}=\langle\mathbf{k}(\cdot,y),f\rangle_{\mathcal{H}}=f(y).$
+  • Super population measure Let $P_{Y}$ denote the super-population law of $Y$. Define
+    $\mu_{\mathbf{k}}(P_{Y}):=\int\mathbf{k}(\cdot,t)\,dP_{Y}(t)\in\mathcal{H},$
+    where the Bochner integral is well-defined under (K1). Then for any $f\in\mathcal{H}$,
+    $\langle\mu_{\mathbf{k}}(P_{Y}),f\rangle_{\mathcal{H}}=\int\langle\mathbf{k}(\cdot,t),f\rangle_{\mathcal{H}}\,dP_{Y}(t)=\int f(t)\,dP_{Y}(t)=\mathbb{E}[f(Y)].$
+    Similarly, for a regular conditional law $P_{Y\mid X=\mathbf{x}}$,
+    $\langle\mu_{\mathbf{k}}(P_{Y\mid X=\mathbf{x}}),f\rangle_{\mathcal{H}}=\int f(t)\,dP_{Y\mid X=\mathbf{x}}(t)=\mathbb{E}[f(Y)\mid X=\mathbf{x}],\qquad f\in\mathcal{H}.$
+    Thus $\mu_{\mathbf{k}}(P_{Y\mid X=\mathbf{x}})$ is the Riesz representer in $\mathcal{H}$ of the functional $f\mapsto\mathbb{E}[f(Y)\mid X=\mathbf{x}]$ restricted to $f\in\mathcal{H}$ (and identifies $P_{Y\mid X=\mathbf{x}}$ if $\mathbf{k}$ is characteristic).
+  • Finite population measure
+    Denote $P_{N,Y}=\frac{1}{N}\sum_{i=1}^{N}\delta_{Y_{i}}$ the finite-population marginal measure of $Y$, its KME is
+    $\mu_{\mathbf{k}}(P_{N,Y})=\int\mathbf{k}(\cdot,t)\,dP_{N,Y}(t)=\frac{1}{N}\sum_{i=1}^{N}\mathbf{k}(\cdot,Y_{i})\in\mathcal{H}.$
+    And the inner product with any $f\in\mathcal{H}$ is
+    $\langle\mu_{\mathbf{k}}(P_{N,Y}),f\rangle_{\mathcal{H}}=\left\langle\frac{1}{N}\sum_{i=1}^{N}\mathbf{k}(\cdot,Y_{i}),f\right\rangle_{\mathcal{H}}=\frac{1}{N}\sum_{i=1}^{N}\langle\mathbf{k}(\cdot,Y_{i}),f\rangle_{\mathcal{H}}=\frac{1}{N}\sum_{i=1}^{N}f(Y_{i}).$
+    In the design space, $(X_{i},Y_{i})$ are considered fixed, so $\langle\mu_{\mathbf{k}}(P_{N,Y}),f\rangle_{\mathcal{H}}$ returns a fixed value.
+  • Finite population conditional measure (estimate)
+    Consider the HJ-type estimator $\hat{P}^{HJ}_{N}(Y|X\in A(\mathbf{x}))$ for the target $P_{N,Y|X\in A(\mathbf{x})}=\sum_{i\in[N]}\mathbbm{1}(X_{i}\in A(\mathbf{x}))\delta_{Y_{i}}/\sum_{i\in[N]}\mathbbm{1}(X_{i}\in A(\mathbf{x}))$, the conditional finite population measure on $A(\mathbf{x})$ s.t.
+$|A(\mathbf{x})|:=\sum_{i\in[N]}\mathbbm{1}(X_{i}\in A(\mathbf{x}))>0$.
+Specifically,
+[…]
+    $\widehat{\omega}_{N,i}(\mathbf{x}):={\dfrac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x}))}/{\hat{N}(\mathbf{x})},\quad\hat{N}(\mathbf{x})=\sum_{i}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in A(\mathbf{x})),$
+    The kernel mean embedding is
+    $\mu_{\mathbf{k}}(\hat{P}^{HJ}_{N}(Y|X\in A(\mathbf{x})))=\int\mathbf{k}(\cdot,y)d\hat{P}^{HJ}_{N}(y|X\in A(\mathbf{x}))=\sum_{i=1}^{N}\widehat{\omega}_{N,i}(\mathbf{x})\mathbf{k}(\cdot,Y_{i}),$
+    where the sum is over all units. While the inner product with any $f\in\mathcal{H}$ is
+    $\langle\mu_{\mathbf{k}}(\hat{P}^{HJ}_{N}(Y\mid X\in A(\mathbf{x}))),f\rangle_{\mathcal{H}}=\sum_{i=1}^{N}\widehat{\omega}_{N,i}(\mathbf{x})\langle\mathbf{k}(\cdot,Y_{i}),f\rangle_{\mathcal{H}}=\sum_{i=1}^{N}\widehat{\omega}_{N,i}(\mathbf{x})f(Y_{i}).$
+    Randomness from the design constitutes the extra weight $\hat{\omega}_{N,i}$.
+So, if one uses the HT estimator (with $N(\mathbf{x})=|A(\mathbf{x})|$ a fixed number rather than $\hat{N}(\mathbf{x})$),
+    $\langle\mathbb{E}_{\xi}[\mu_{\mathbf{k}}(\hat{P}^{HT}_{N}(Y\mid X\in A(\mathbf{x})))],f\rangle_{\mathcal{H}}=\frac{1}{N(\mathbf{x})}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in A(\mathbf{x}))f(Y_{i})=\int f\,dP_{N,Y\mid X\in A(\mathbf{x})},$
+    unbiased of $\int f\,dP_{N,Y\mid X\in A(\mathbf{x})}$, while the HJ-type estimator is generally biased (unless $\hat{N}(\mathbf{x})=N(\mathbf{x})$ but design-consistent.
+S5 Bagging and Honesty
+In the complex survey design forest setting, we need to introduce a more sophisticated resampling approach and a new criterion for the tree construction that mimic the design:
+  1. 1.
+    The resamples used to start a tree maintain the original design structure.
+  2. 2.
+    Honesty remains in that the data used for the estimation of $\widehat{\omega}_{i}$ are separate from the data used to determine the splits.
+S5.1 Resampling
+Proof of Lemma 3.1.
+We focus on proving (R2) and (R3) for some single–stage designs, as (R1) holds directly from the re–sampling procedure detailed below.
+Poisson Sampling: When the original sample is obtained by Poisson sampling, $\mathbf{p}(S,Z^{N})=\prod_{i\in S}\pi_{i}(Z^{N})\prod_{i\notin S}(1-\pi_{i}(Z^{N}))$, so $\xi_{i}\mid Z^{N}\sim Bernoulli(\pi_{i}(Z^{N}))$, where $\pi_{i}(Z^{N})$ is possibly different for all $i=1,\dots,N$.
+Thus, for subsamples indexed by $b=1,\dots,B$: Draw $(N^{*}_{1},\dots,N^{*}_{n_{s}})\mid Z^{N}\sim Multinomial(N,\bm{w}^{*})$, $\bm{w}^{*}=(\bm{w}_{j}^{*})_{j=1}^{n_{s}}$, $\bm{w}_{j}^{*}=(1/\pi_{j})/\sum_{i=1}^{n_{s}}1/\pi_{i}$.
+Obtain a rebuilt finite population by replicating $D_{i}=(X_{i},\mathbf{k}(Y_{i},\cdot))$ by $N^{*}_{i}$ times for $i=1,\dots,n_{s}$.
+Then draw inclusion indicators for the subsample.
+Equivalently, this is generating multipliers $n^{*}_{i}\mid N^{*}_{i}\sim Binomial(N^{*}_{i},\pi_{i})$, $i=1,\dots,n_{s}$.
+This results in the subsample ${D}^{*}_{b}=\{(D_{i},\hat{w}_{i}=(1/\pi_{i})n^{*}_{i})\}_{i=1}^{n_{s}}$.
+(R2) holds as $\mathbb{E}(n_{i}^{*}\mid\xi)=\mathbb{E}\!\left[\mathbb{E}(n_{i}^{*}\mid N_{i}^{*},\xi)\mid\xi\right]=\mathbb{E}(N_{i}^{*}\pi_{i}\mid\xi)=\pi_{i}\,\mathbb{E}(N_{i}^{*}\mid\xi)=\pi_{i}w_{i}\xi_{i}=\xi_{i}.$ Also, for any $r=2+\epsilon>2$, the two-stage procedure leads to $n^{*}_{i}\mid\xi\sim\mathrm{Binomial}(\sum_{i\in[n_{s}]}\frac{1}{\pi_{i}},\;\frac{1}{\sum_{i\in[n_{s}]}1/\pi_{i}})$, $\forall i\in[N].$ so it is sufficient to show
+$\sup_{m\geq 1,p\in(0,1]\;:\,pm=1}\mathbb{E}\Big[\mathrm{Binomial}\big(m,p\big)^{r}\Big]<\infty.$ (S5)
+This is true, as for random variable $X\sim Bin(m,\;p)$, for all $t>0$, $M_{X}(t)=\mathbb{E}(e^{tX})=(1-p+pe^{t})^{m}=\big(1+p(e^{t}-1)\big)^{m}\leq\exp(e^{t}-1)$ by $1+x\leq e^{x}$.
+As $x^{r}\leq\left(\frac{r}{t}\right)^{r}e^{tx}\quad\forall t>0,\,x\geq 0,$ we can bound moments of $n^{*}_{i}\mid\xi$ by its MGF as
+$\mathbb{E}[(n^{*}_{i})^{r}\mid\xi]\leq\left(\frac{r}{t}\right)^{r}\mathbb{E}[e^{tn^{*}_{i}}\mid\xi]\leq\left(\frac{r}{t}\right)^{r}\exp(e^{t}-1).$
+Pick $t=1$ gives us $\mathbb{E}[{(n^{*}_{i})}^{r}]\leq r^{r}\exp(e-1)=:C_{r},$ a finite constant depends only on finite $r$. Therefore (S5) holds, and
+$\sup_{i\in[N]}\mathbb{E}\!\Big[\;|n_{i}^{*}|^{2+\varepsilon}\mid\xi\Big]\leq C_{2+\varepsilon}<\infty.$
+(R3) can be deduced directly from the distribution result for $n^{*}\mid\xi$ obtained previously: Denote $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}$, then $\mathrm{Cov}(n_{i}^{*},\,n_{j}^{*}\mid\xi)=-\,\frac{1}{\widehat{N}}$, and $\mathrm{Var}(n_{i}^{*}\mid\xi)=\widehat{N}\left(\frac{1}{\widehat{N}}\right)\left(1-\frac{1}{\widehat{N}}\right)=1-\frac{1}{\widehat{N}}$.
+Therefore
+$\sum_{i\neq j}\bigl|\mathrm{Cov}(n_{i}^{*},n_{j}^{*}\mid\xi)\bigr|=\frac{n_{s}(n_{s}-1)}{\widehat{N}}\leq n_{s}(1-\frac{1}{\widehat{N}})=\sum_{i}\mathrm{Var}(n_{i}^{*}\mid\xi),$
+as $\widehat{N}=\sum_{i\in[n_{s}]}1/\pi_{i}\geq\sum_{i\in[n_{s}]}i=n_{s}$, so (R3) holds with $C=1$.
+SRSWOR (SRS without Replacement): When the original sampling scheme is SRSWOR, we specify a fixed sample size $n_{s}$, and $n=n_{s}$.
+Accordingly, $\mathbf{p}(S,z^{N})={1}/{\binom{N}{n_{s}}}\,\mathbbm{1}\{|S|=n_{s}\}$, and every subset $S$ of size $n_{s}$ is equally likely, with $\pi_{i}=n_{s}/N$, $\forall i=1,\dots,n_{s}$.
+So we first draw $(N^{*}_{1},\dots,N^{*}_{n_{s}})|Z^{N}\sim Multinomial(N,\bm{w}^{*})$, $\bm{w}^{*}=(\bm{w}_{j}^{*})_{j=1}^{n_{s}}$, $\bm{w}_{j}^{*}=1/n_{s}$.
+Replicate $D_{i}=(X_{i},\mathbf{k}(Y_{i},\cdot))$ by $N^{*}_{i}$ times for $i=1,\dots,n_{s}$.
+Then, realize inclusion indicators for the subsample $\bm{\xi}^{*}$ following the original sampling scheme.
+The resulting subsample is then ${D}^{*}_{b}=\{(D_{i},\hat{w}_{i}=(1/\pi_{i})\sum_{j=1}^{N^{*}_{i}}\xi^{*}_{j}\}_{i=1}^{n_{s}}$.
+This two–stage procedure implies $n_{i}^{*}\mid(N_{i}^{*},\xi)\sim\mathrm{Hypergeometric}\!\left({N},\,N_{i}^{*},\,n\right),\,\forall i.$ So
+$\mathbb{E}(n_{i}^{*}\mid\xi)=\mathbb{E}\!\left[\,n\,\frac{N_{i}^{*}}{{N}}\,\Big|\,\xi\right]=n\,\frac{\mathbb{E}(N_{i}^{*}\mid\xi)}{{N}}=n\,\frac{1/\pi_{i}}{{N}}=n\,\frac{{N}/n}{{N}}=1.$
+As $N^{*}_{i}/N<1$, $0\leq n^{*}_{i}\leq n$ always, and the MGF of hypergeometric is dominated by a binomial with the same mean:
+$\mathbb{E}[e^{tn^{*}_{i}}\mid N^{*},\xi]\leq\mathbb{E}[e^{tY_{i}}]=\left(1-\frac{N^{*}_{i}}{N}+\frac{N^{*}_{i}}{N}e^{t}\right)^{n},$
+with $Y_{i}\sim\mathrm{Bin}(n,N^{*}_{i}/N)$. By a similar argument as that for the Poisson sampling scenario, for $r=2+\epsilon$, for any $t>0$,
+$\mathbb{E}\!\left[(n_{i}^{*})^{r}\mid\xi\right]$ $\leq\left(\frac{r}{t}\right)^{r}\mathbb{E}\!\left[\exp\!\big(n\frac{N^{*}_{i}}{N}(e^{t}-1)\big)\mid\xi\right]$ (S6)
+Now from $N^{*}_{i}\xi\sim\mathrm{Bin}(N,\;1/n)$, let $c=e^{t}-1$,
+[…]
+where the last inequality uses $e^{cn/N}-1\leq(cn/N)e^{cn/N}$. By (D2), $n/N=f\in(0,1]$, bounded for all $N$, so setting $t=1$ gives
+$\sup_{i\in[N]}\,\mathbb{E}\!\left[(n_{i}^{*})^{r}\mid\xi\right]\leq r^{r}\cdot e^{(e-1)e^{(e-1)}}=C_{r}<\infty,\quad\text{proving $(R2)$}.$
+Also from the hyper–geometric distribution, $\mathrm{Cov}(n_{i}^{*},n_{j}^{*}\mid N^{*},\xi)=-\,\frac{n({N}-n)}{{N}-1}\,\frac{N_{i}^{*}}{{N}}\,\frac{N_{j}^{*}}{{N}},\,i\neq j,$ $\mathrm{Var}(n_{i}^{*}\mid N^{*},\xi)=\frac{n({N}-n)}{{N}-1}\,\frac{N_{i}^{*}}{{N}}\left(1-\frac{N_{i}^{*}}{{N}}\right).$ As $\sum_{i}N^{*}_{i}/N=1$, $\sum_{i\neq j}\frac{N^{*}_{i}}{N}\frac{N^{*}_{j}}{N}=\sum_{i}(\frac{N^{*}_{i}}{N}-(\frac{N^{*}_{i}}{N})^{2})$ so $\sum_{i\neq j}\mathbb{E}\left[\left|\mathrm{Cov}\!(n_{i}^{*},n_{j}^{*}\mid N^{*},\xi)\right|\mid\xi\right]\leq\sum_{i}\mathbb{E}\left[\mathrm{Var}\!\left(n_{i}^{*}\mid N^{*},\xi\right)\mid\xi\right],$ proving (R3) with $C=1$.
+PPS With Replacement: When the original sampling scheme is PPSWR, we specify a fixed sample size $n_{s}=n>0$ and for all units a measure of size (MOS), and draw indices $a(1),\dots,a(n)\in[N]$ i.i.d.
+with $P(a(j)=i)=p_{i}\propto\mathrm{MOS}_{i}$, $\sum_{i}p_{i}=1$.
+Re–sampling can be done by: (1) drawing $(\tilde{N}_{1}^{*},\ldots,\tilde{N}_{n_{s}}^{*})\mid\xi\sim\mathrm{Multinomial}(N-n,\bm{w}^{*})$ with $w_{j}^{*}\;=\;\frac{p_{a(j)}^{-1}}{\sum_{\ell=1}^{n_{s}}p_{a(\ell)}^{-1}},\,j=1,\ldots,n_{s}.$.
+Set $N^{*}_{i}:=1+\tilde{N}^{*}_{i}$.
+(2) Replicate the observed units $D_{a(j)}=(X_{a(j)},\mathbf{k}(Y_{a(j)},\cdot))$ exactly $N_{j}^{*}$ times to form the pseudo finite population.
+(3) Normalize the resulting bootstrap resampling probabilities for the realized draws to sum to one by $p^{*}_{i}=\frac{N^{*}_{i}p_{a(i)}}{\sum_{k\in[n_{s}]}N^{*}_{k}p_{a(k)}}$, $\forall i\in[n_{s}]$.
+(4) Re-apply the original PPSWR mechanism on the bootstrap population by drawing $b(1),\dots,b(n_{s})$ i.i.d.
+on $n_{s}]$ with $\Pr\{b(j)=i\}=p_{i}^{*}$.
+Finally, define count $m^{*}_{i}=\sum_{j\in[n_{s}]}\mathbbm{1}(b(j)=i)$, and let $n^{*}_{i}=\frac{m^{*}_{i}}{n_{s}p^{*}_{i}}$ for $i\in[n_{s}]$.
+By (PPS-MOS) $\exists c>0$, $C<\infty$ such that $c\leq MOS_{i}\leq C$, then
+$\frac{c_{-}}{N}\leq p_{i}=\frac{MOS_{i}}{\sum_{j}MOS_{j}}\leq\frac{c_{+}}{N}$ (S7)
+where $c_{-}:=c/C$, $c_{+}:=C/c$ both bounded constants away from 0.
+Conditional on $(p^{*},\xi)$, $(m_{1}^{*},\ldots,m_{n}^{*})\sim\mathrm{Multinomial}(n,p^{*})$, hence $\mathbb{E}\!\left(m_{i}^{*}\mid p^{*},\xi\right)=np_{i}^{*}$ and $\mathbb{E}\!\left(n_{i}^{*}\mid p^{*},\xi\right)=\frac{1}{np_{i}^{*}}\,\mathbb{E}\!\left(m_{i}^{*}\mid p^{*},\xi\right)=1.$ Taking expectation over $p^{*}\mid\xi$ yields $\mathbb{E}\!\left(n_{i}^{*}\mid\xi\right)=\xi_{i},$ which implies the required $\mathbb{E}(n_{i}^{*}\mid\xi)=1+o(1)$.
+Consider
+$\mathcal{E}_{N}:=\{\,f_{-}\leq n_{s}/N\leq f_{+}\,\}\text{ with $f_{-}$ and $f_{+}$ constants},$
+$\mathcal{E}_{N}$ satisfies $P_{\mathcal{S}^{N}\mid\omega}(\mathcal{E}_{N})\to 1$ by (D2).
+On $\mathcal{E}_{N}$, using (S7): since $\sum_{k=1}^{n}N_{k}^{*}=N$ and $p_{a(k)}\leq c_{+}/N$, $S^{*}=\sum_{k=1}^{n}N_{k}^{*}\,p_{a(k)}\leq\sum_{k=1}^{n}N_{k}^{*}\,\frac{c_{+}}{N}=c_{+};$ for each $i$, $N_{i}^{*}\geq 1$ and $p_{a(i)}\geq c_{-}/N$, so
+$p_{i}^{*}=\frac{N_{i}^{*}\,p_{a(i)}}{S^{*}}\geq\frac{1\cdot(c_{-}/N)}{c_{+}}=\frac{c_{-}}{c_{+}}\cdot\frac{1}{N}.$ (S8)
+Multiplying by $n$ and using $n/N\geq f_{-}$ on $\mathcal{E}_{N}$, we have that
+$np_{i}^{*}\geq f_{-}\frac{c_{-}}{c_{+}}=:\lambda_{-}>0,\qquad\forall i\in[n],\ \text{on }\mathcal{E}_{N}.$
+Now we prove for the $r=2+\epsilon$ moments.
+Fix $i$.
+$m_{i}^{*}\mid p^{*},\xi\sim\mathrm{Binomial}(n,p_{i}^{*}).$ Write $m_{i}^{*}=\sum_{t=1}^{n}B_{t,i}$ with $B_{t,i}:=\mathbbm{1}(b(t)=i)$, $B_{t,i}\mid p^{*},\xi\sim\mathrm{Bernoulli}$.
+Set $X_{t}:=B_{t,i}-p_{i}^{*}$.
+By Rosenthal’s inequality ([1970_Rosenthal]) on $X_{t}\mid p^{*},\xi$, there exists $K_{r}<\infty$ such that
+$\mathbb{E}\!\left(\bigl|m_{i}^{*}-np_{i}^{*}\bigr|^{r}\,\big|\,p^{*},\xi\right)$ $\leq K_{r}\!\left(np_{i}^{*}(1-p_{i}^{*})+\bigl(np_{i}^{*}(1-p_{i}^{*})\bigr)^{r/2}\right)\leq K_{r}\!\left(np_{i}^{*}+(np_{i}^{*})^{r/2}\right).$
+Hence, using $n_{i}^{*}=1+(m_{i}^{*}-np_{i}^{*})/(np_{i}^{*})$ and $|x+y|^{r}\leq 2^{r-1}(|x|^{r}+|y|^{r})$,
+$\mathbb{E}\!\left(|n_{i}^{*}|^{r}\,\big|\,p^{*},\xi\right)$ $\leq 2^{r-1}\left(1+\frac{\mathbb{E}\!\left(|m_{i}^{*}-np_{i}^{*}|^{r}\,\big|\,p^{*},\xi\right)}{(np_{i}^{*})^{r}}\right)$
+$\leq 2^{r-1}\left(1+K_{r}\left(\frac{1}{(np_{i}^{*})^{r-1}}+\frac{1}{(np_{i}^{*})^{r/2}}\right)\right).$
+On $\mathcal{E}_{N}$, apply the lower bound (S8) to obtain $\sup_{i\in[n]}\mathbb{E}\!\left(|n_{i}^{*}|^{r}\,\big|\,p^{*},\xi\right)\leq C_{r}<\infty\,\text{on }\mathcal{E}_{N},$ for a constant $C_{r}$ depending only on $r$ and $\lambda_{-}$.
+Taking expectation over $p^{*}\mid\xi$ preserves the bound, and since $P_{\mathcal{S}^{N}\mid\omega}(\mathcal{E}_{N})\to 1$, this completes (R2).
+For (R3), from the multinomial distribution, $\mathrm{Cov}(m_{i}^{*},m_{j}^{*}\mid p^{*},\xi)=-np_{i}^{*}p_{j}^{*},$, $\mathrm{Var}(m_{i}^{*}\mid p^{*},\xi)=np_{i}^{*}(1-p_{i}^{*}).$ So for the multipliers
+$\mathrm{Cov}(n_{i}^{*},n_{j}^{*}\mid p^{*},\xi)=\frac{\mathrm{Cov}(m_{i}^{*},m_{j}^{*}\mid p^{*},\xi)}{n^{2}p_{i}^{*}p_{j}^{*}}=-\frac{1}{n},\qquad i\neq j,$
+so $\sum_{i\neq j}\bigl|\mathrm{Cov}(n_{i}^{*},n_{j}^{*}\mid p^{*},\xi)\bigr|=\frac{n(n-1)}{n}=n-1.$ The conditional variance satisfy
+$\sum_{i=1}^{n}\mathrm{Var}(n_{i}^{*}\mid p^{*},\xi)=\sum_{i=1}^{n}\frac{\mathrm{Var}(m_{i}^{*}\mid p^{*},\xi)}{n^{2}(p_{i}^{*})^{2}}=\frac{1}{n}\sum_{i=1}^{n}\left(\frac{1}{p_{i}^{*}}-1\right)=\frac{1}{n}\left(\sum_{i=1}^{n}\frac{1}{p_{i}^{*}}-n\right).$
+Since $\sum_{i=1}^{n}p_{i}^{*}=1$ and $p_{i}^{*}>0$, the AM–HM inequality implies $\frac{1}{n}\sum_{i=1}^{n}\frac{1}{p_{i}^{*}}\geq\ n^{2}.$ Thus $\sum_{i=1}^{n}\mathrm{Var}(n_{i}^{*}\mid\mathbf{p}^{*},\xi)\geq\frac{1}{n}(n^{2}-n)=n-1.$ So (R3) holds with $C=1$.
+Now consider two–stage stratified designs with strata $h=1,\ldots,H$, $H\geq 2$.
+Lemma S3 (Two-stage product multipliers inherit (R1)–(R3)).
+Fix a stratum $h$.
+Let $\{G_{hk}^{*}\}_{k=1}^{K_{h}}$ be PSU-level bootstrap multipliers generated by a single-stage resampling scheme applied to PSUs in stratum $h$.
+For each PSU $(h,k)$ with $\xi_{hk}=1$, let $\{W_{hkj}^{*}\}_{j=1}^{m_{hk}}$, $m_{hk}=\sum_{j\in[M_{hk}]}\xi_{hkj}$ be SSU-level multipliers generated by a (possibly different) single-stage resampling scheme applied within PSU $(h,k)$.
+Let $\mathcal{F}:=\sigma\Big(\xi,\ \{Z_{i}\}_{i\leq N}\Big)$ be the $\sigma$–field generated by design and sampling available at time of re-sampling.
+Let $G^{*}:=\{\,G_{hk}^{*}:h=1,\ldots,H,\ k=1,\ldots,K_{h}\,\}$ the collection of PSU–level multipliers.
+If
+  1. (R1-TS)
+    Conditional on $(\mathcal{F},G^{*})$, the collections $\{W_{hkj}^{*}\}_{j=1}^{m_{hk}}$ are independent across distinct PSUs $(h,k)$.
+  2. (R2-TS)
+    There exists $\varepsilon>0$ such that,
+[…]
+    $\mathbb{E}(W_{hkj}^{*}\mid\mathcal{F},G^{*})=1,\,\forall h,k,j\qquad\sup_{h,k,j}\mathbb{E}\bigl(|W_{hkj}^{*}|^{2+\varepsilon}\mid\mathcal{F},G^{*}\bigr)<\infty.$
+  3. (R3-TS)
+    Denote $\mathcal{U}_{hk}$ the set of sampled SSU indices. (so that $m_{hk}=|\mathcal{U}_{hk}|$)
+    $\sum_{(h,k)\neq(h^{\prime},k^{\prime})}\left|Cov\!\left(G^{*}_{hk},\,G^{*}_{h^{\prime}k^{\prime}}\mid\mathcal{F}\right)\right|$ $\leq C_{G}\sum_{h,k}Var\!\left(G^{*}_{hk}\mid\mathcal{F}\right).$
+    $\sum_{\begin{subarray}{c}j\neq j^{\prime}\\ j,j^{\prime}\in\mathcal{U}_{hk}\end{subarray}}\left|Cov\!\left(W^{*}_{hkj},\,W^{*}_{hkj^{\prime}}\mid\mathcal{F},\,G^{*}\right)\right|$ $\leq C_{W}\sum_{j\in\mathcal{U}_{hk}}Var\!\left(W^{*}_{hkj}\mid\mathcal{F},\,G^{*}\right).$
+  4. (R4-TS)
+    For $m_{hk}=\sum_{j\in[M_{hk}]}\xi_{hkj}$ PSU sample sizes, $\sup_{h,k}m_{hk}<\infty$ .
+Then the unit multipliers $n_{hkj}^{*}:=G_{hk}^{*}\,W_{hkj}^{*}$ satisfy (R1)(R3).
+Proof of Lemma S3.
+Suppose we generate BS multipliers in this two–step approach:
+Stratified Two–Stage Clustering
+For each stratum $h$, separately,
+  • Generate PSU-level bootstrap multipliers $\{G_{hk}^{*}\}_{k=1}^{K_{h}}$ using one of the single-stage re–sampling schemes (e.g.
+Poisson, SRSWOR, and PPSWR verified above) applied at PSU level (using PSU-level weights or MOS), such that $G_{hk}^{*}\geq 0$;
+  • Conditional on the PSU bootstrap outcome for PSUs with $\xi_{hk}=1$, generate SSU-level multipliers $\{W_{hkj}^{*}\}_{j=1}^{m_{hk}}$ within each sampled PSU $(h,k)$, using possibly a different validated single-stage scheme applied at the SSU level (using SSU weights and MOS within the PSU).
+Then the first condition (i) is met.
+This is (R1) in Lemma 3.1.
+Now we show (R2).
+For any $u=(h,k,j)$, if a re-sampling scheme described for above single–stage designs are used, (R2) is met and $\mathbb{E}\!\left(n_{u}^{*}\mid\mathcal{F}\right)=\mathbb{E}\!\left(\mathbb{E}\!\left(n_{u}^{*}\mid\mathcal{F},G^{*}\right)\mid\mathcal{F}\right)=\mathbb{E}\!\left(G_{hk}^{*}\mid\mathcal{F}\right)=1.$ Let $r:=2+\varepsilon$.
+Conditional on $\mathcal{F}$, $\mathbb{E}\!\left(|n_{u}^{*}|^{r}\mid\mathcal{F}\right)=\mathbb{E}\!\left(|G_{hk}^{*}|^{r}\,|W_{hkj}^{*}|^{r}\mid\mathcal{F}\right).$ Apply conditional Hölder with conjugate exponents $p=q=2$,
+$\mathbb{E}\!\left(|G^{*}|^{r}|W^{*}|^{r}\mid\mathcal{F}\right)\leq\left(\mathbb{E}\!\left(|G^{*}|^{2r}\mid\mathcal{F}\right)\right)^{1/2}\left(\mathbb{E}\!\left(|W^{*}|^{2r}\mid\mathcal{F}\right)\right)^{1/2}\leq\infty,$
+as $\mathbb{E}[|W^{*}|^{2r}\mid\mathcal{F}]=\mathbb{E}\mathbb{E}[|W^{*}|^{2r}\mid\mathcal{F}]<C_{W}<\infty$ from (ii); and $\mathbb{E}[|G^{*}|^{2r}]<C_{G}<\infty$ directly from (ii).
+Hence $\sup_{u}\,\mathbb{E}\!\left(\,|n_{u}^{*}|^{r}\mid\xi\right)\leq\sqrt{C_{G}^{\prime}C_{W}^{\prime}}\;<\;\infty$.
+For (R3): for any two sampled SSU $u=(h,k,j)\neq v=(h^{\prime},k^{\prime},j^{\prime})$,
+$\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{F})$ $=\mathbb{E}\!\left[\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{G})\mid\mathcal{F}\right]+\mathrm{Cov}\!\left(\mathbb{E}(n_{u}^{*}\mid\mathcal{G}),\,\mathbb{E}(n_{v}^{*}\mid\mathcal{G})\mid\mathcal{F}\right)$
+$=\mathbb{E}\left[(G_{hk}^{*})^{2}\,\mathrm{Cov}(W_{u}^{*},W_{v}^{*}\mid\mathcal{G})\right]+\mathrm{Cov}\!\left(G^{*}_{\mathrm{psu}(u)},\,G^{*}_{\mathrm{psu}(v)}\mid\mathcal{F}\right),$
+[…]
+and equals $\mathrm{Cov}\!\big(G^{*}_{\mathrm{psu}(u)},\,G^{*}_{\mathrm{psu}(v)}\mid\mathcal{F}\big)$ if $u$ and $v$ belong to different PSUs, where we use $\mathbb{E}(n_{u}^{*}\mid\mathcal{G})=G_{\mathrm{psu}(u)}^{*}$ by (R2-TS).
+We prove (R3) under the two cases.
+Between PSU: The covariance term satisfy
+$\sum_{\begin{subarray}{c}u\neq v\\ \mathrm{psu}(u)\neq\mathrm{psu}(v)\end{subarray}}\bigl|\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{F})\bigr|=$ $\sum_{(h,k)\neq(h^{\prime},k^{\prime})}m_{hk}m_{h^{\prime}k^{\prime}}\,\bigl|\mathrm{Cov}(G_{hk}^{*},G_{h^{\prime}k^{\prime}}^{*}\mid\mathcal{F})\bigr|$
+$\leq$ $m_{\max}^{2}C_{G}\sum_{h,k}\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})$
+by (R4-TS) and (R3-TS). While by (R2-TS),
+$\mathrm{Var}(n_{u}^{*}\mid\mathcal{F})=\mathbb{E}\!\left[\mathrm{Var}(n_{u}^{*}\mid\mathcal{G})\mid\mathcal{F}\right]+\mathrm{Var}\!\left(\mathbb{E}(n_{u}^{*}\mid\mathcal{G})\mid\mathcal{F}\right)\geq\mathrm{Var}(G_{\mathrm{psu}(u)}^{*}\mid\mathcal{F}).$
+Since $m_{hk}\geq 1$ whenever $\mathcal{U}_{hk}\neq\varnothing$, it follows $\sum_{u\in\mathcal{U}_{hk}}\mathrm{Var}(n_{u}^{*}\mid\mathcal{F})\geq m_{hk}\,\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})\geq\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})$, thus
+$\sum_{\begin{subarray}{c}u\neq v\\ \mathrm{psu}(u)\neq\mathrm{psu}(v)\end{subarray}}\bigl|\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{F})\bigr|\leq m_{\max}^{2}C_{G}\sum_{u}\mathrm{Var}(n_{u}^{*}\mid\mathcal{F}).$ (S9)
+Within one PSU: For any fixed $(h,k)$, $u,v\in\mathcal{U}_{hk}$, by the triangle inequality and Jensen’s inequality,
+$\sum_{\begin{subarray}{c}u\neq v\\ u,v\in\mathcal{U}_{hk}\end{subarray}}\bigl|\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{F})\bigr|\leq\sum_{\begin{subarray}{c}u\neq v\\ u,v\in\mathcal{U}_{hk}\end{subarray}}\mathbb{E}\!\left[(G_{hk}^{*})^{2}\,\bigl|\mathrm{Cov}(W_{u}^{*},W_{v}^{*}\mid\mathcal{G})\bigr|\mid\mathcal{F}\right]+m_{hk}(m_{hk}-1)\,\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F}).$
+The first term, which we denote $A_{hk}$, satisfy by (R3-TS)
+$A_{hk}$ $\leq C_{W}\sum_{u\in\mathcal{U}_{hk}}\mathbb{E}\!\left[(G_{hk}^{*})^{2}\,\mathrm{Var}(W_{u}^{*}\mid\mathcal{G})\mid\mathcal{F}\right]\leq C_{W}\sum_{u\in\mathcal{U}_{hk}}\mathrm{Var}(n_{u}^{*}\mid\mathcal{F}),\text{ as }$
+$\mathrm{Var}(n_{u}^{*}\mid\mathcal{F})$ $=\mathbb{E}\!\left[(G_{hk}^{*})^{2}\,\mathrm{Var}(W_{u}^{*}\mid\mathcal{G})\mid\mathcal{F}\right]+\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})\geq\mathbb{E}\!\left[(G_{hk}^{*})^{2}\,\mathrm{Var}(W_{u}^{*}\mid\mathcal{G})\mid\mathcal{F}\right].$
+For the second term, as $m_{hk}(m_{hk}-1)\leq m_{\max}m_{hk}$,
+$m_{hk}(m_{hk}-1)\,\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})\leq m_{\max}m_{hk}\,\mathrm{Var}(G_{hk}^{*}\mid\mathcal{F})\leq m_{\max}\sum_{u\in\mathcal{U}_{hk}}\mathrm{Var}(n_{u}^{*}\mid\mathcal{F}).$ (S10)
+So together we have the inner part of below. Then summing all PSUs gives
+$\sum_{h,k}\sum_{\begin{subarray}{c}u\neq v\\ u,v\in\mathcal{U}_{hk}\end{subarray}}\bigl|\mathrm{Cov}(n_{u}^{*},n_{v}^{*}\mid\mathcal{F})\bigr|\leq(C_{W}+m_{\max})\sum_{h,k}\sum_{u\in\mathcal{U}_{hk}}\mathrm{Var}(n_{u}^{*}\mid\mathcal{F}),$ (S11)
+finishing the proof for (R3), as it holds for both cases. ∎
+Remark.
+The re-sampling can be only at the PSU level, even if the original design has two stages. Including SSU level resampling is beneficial when the sample contains a small number of PSUs of large sizes.
+S6 Splitting Criteria
+This section examines the theoretical properties at an arbitrary split node.
+S6.1 Equivalent expression
+Proposition S4.
+A weighted MMD is equivalent in concept to the usual CART split criteria. Specifically, that
+$\begin{split}&\hat{\theta}_{n_{s}}({D})=\arg\max_{\theta\in\Pi}\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\left\|\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{\{i\in S:X_{i}\in C_{L}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{R}}\sum_{\{i\in S:X_{i}\in C_{R}\}}\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}}\\ =&\arg\min_{\theta\in\Pi}\frac{1}{\hat{N}_{P_{a}}}\sum_{\{i\in S:X_{i}\in P_{a}\}}\left\|\mu_{\mathbf{k}}\left(\frac{n^{*}_{b,i}}{q\times\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\sum_{l\in\{L,R\}}\frac{1_{X_{i}\in C_{l}}}{\hat{N}_{l}}\sum_{\{j\in S:X_{j}\in C_{l}\}}\frac{n^{*}_{b,j}}{q\times\pi_{j}}\;\delta_{Y_{j}}\right)\right\|^{2}_{\mathcal{H}}.\end{split}$ (S12)
+denote $\hat{w}_{i}=\frac{n^{*}_{b,i}}{q\times\pi_{i}}$ and let $\hat{N}_{l}=\sum_{X_{i}\in D_{b}}\mathbbm{1}_{\{X_{i}\in C_{l}\}}\hat{w}^{*}_{i}.$
+Proof of Proposition S4.
+We use a total distance decomposition argument, using the linearity of $\mu_{\mathbf{k}}(\cdot)$. Denote
+$M_{i}=\mu_{\mathbf{k}}\left(\hat{w}_{i}\;\delta_{Y_{i}}\right),\quad\bar{M}=\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{P_{a}}}\sum_{X_{i}\in P_{a}}\hat{w}_{i}\;\delta_{Y_{i}}\right),\quad\bar{M}_{L}=\mu_{\mathbf{k}}\left(\frac{1}{\hat{N}_{L}}\sum_{X_{i}\in P_{a}}{1_{X_{i}\in C_{L}}}\hat{w}_{i}\;\delta_{Y_{i}}\right)$
+and similarly for $M_{R}$, so the RHS of (5) is equivalently
+$\frac{1}{\hat{N}_{P_{a}}}\left\{\sum_{i\in C_{L}}\|M_{i}-\bar{M}_{L}\|^{2}_{\mathcal{H}}+\sum_{i\in C_{R}}\|M_{i}-\bar{M}_{R}\|^{2}_{\mathcal{H}}\right\},$ (S13)
+the sum of distances within groups. As in $\hat{N}_{P_{a}}=\hat{N}_{L}+\hat{N}_{R}$ and $\bar{M}-\bar{M}_{L}=\frac{\hat{N}_{R}}{\hat{N}_{P_{a}}}(\bar{M}_{R}-\bar{M}_{L})$,
+$\sum_{i\in C_{L}}\|M_{i}-M_{L}\|^{2}_{\mathcal{H}}=\sum_{i\in C_{L}}\|M_{i}-\bar{M}\|^{2}_{\mathcal{H}}-\frac{\hat{N}^{2}_{R}\hat{N}_{L}}{\hat{N}^{2}_{P_{a}}}\|\bar{M}_{R}-\bar{M}_{L}\|^{2}_{\mathcal{H}}$ (S14)
+and similarly for $\sum_{i\in C_{R}}\|M_{i}-\bar{M}_{R}\|^{2}_{\mathcal{H}}$.
+Plugging in the expression into (6), and realizing that the total sum of distances $\frac{1}{\hat{N}_{P_{a}}}\sum_{i\in P_{a}}\|M_{i}-\bar{M}\|^{2}_{\mathcal{H}}$ is invariant w.r.t.
+the split results in $\operatorname*{arg\,max}(6)=\operatorname*{arg\,max}\frac{1}{\hat{N}_{P_{a}}}(-\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}_{P_{a}}})\|\bar{M}_{R}-\bar{M}_{L}\|^{2}_{\mathcal{H}}\equiv\operatorname*{arg\,min}\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}^{2}_{P_{a}}}\|\bar{M}_{R}-\bar{M}_{L}\|^{2}_{\mathcal{H}},$ completing the proof.
+∎
+S6.2 Consistency of the Split
+Lemma S5 (Lipschitz continuity of split scores).
+Under assumptions (K1), (A2), and (A3), for any parent region $P_{a}\in\mathcal{C}_{d-1}$ with $N_{P_{a}}:=|\{i\in[N]:X_{i}\in P_{a}\}|$, both the finite-population split score
+$M^{N}(\theta)=\frac{N_{L}(\theta)N_{R}(\theta)}{N_{P_{a}}^{2}}\left\|\mu_{k}\left(\frac{1}{N_{L}(\theta)}\sum_{X_{i}\in C_{L}(\theta)}\delta_{Y_{i}}\right)-\mu_{k}\left(\frac{1}{N_{R}(\theta)}\sum_{X_{i}\in C_{R}(\theta)}\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}}$
+and the design-weighted split score
+$M^{*}_{n_{s}}(\theta)=\frac{N_{L}^{n_{s}}(\theta)N_{R}^{n_{s}}(\theta)}{(N_{P_{a}}^{n_{s}})^{2}}\left\|\mu_{k}\left(\frac{1}{N_{L}^{n_{s}}(\theta)}\sum_{X_{i}\in C_{L}(\theta)}\frac{\xi_{i}}{\pi_{i}}\delta_{Y_{i}}\right)-\mu_{k}\left(\frac{1}{N_{R}^{n_{s}}(\theta)}\sum_{X_{i}\in C_{R}(\theta)}\frac{\xi_{i}}{\pi_{i}}\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}}$
+are Lipschitz continuous on $\Pi$ with respect to the metric $d(\theta,\theta^{\prime})=\mathbbm{1}(j\neq j^{\prime})+|t-t^{\prime}|$.
+Specifically, there exist constants $L_{1},L_{2}>0$ (depending on $K_{\max}$, the parameters from (A2)-(A3), and $\bar{\lambda},\underline{\lambda}$ from (D3)) such that for all $\theta,\theta^{\prime}\in\Pi$: $|M^{N}(\theta)-M^{N}(\theta^{\prime})|\leq L_{1}\cdot d(\theta,\theta^{\prime}),$ and $|M_{n_{s}}(\theta)-M_{n_{s}}(\theta^{\prime})|\leq L_{2}\cdot d(\theta,\theta^{\prime}).$
+Proof.
+We prove the result for $M^{N}(\theta)$; the proof for $M_{n_{s}}(\theta)$ follows the same structure with design-weighted counts.
+Write $M^{N}(\theta)=w(\theta)\cdot\Delta(\theta)^{2}$, where
+$w(\theta):=\frac{N_{L}(\theta)N_{R}(\theta)}{N_{P_{a}}^{2}},\quad\Delta(\theta):=\left\|\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{R}(\theta))\right\|_{\mathcal{H}},$
+$P_{\ell}(\theta):=\frac{1}{N_{\ell}(\theta)}\sum_{X_{i}\in C_{\ell}(\theta)}\delta_{Y_{i}},\quad\ell\in\{L,R\}.$
+We first bound $|w(\theta)-w(\theta^{\prime})|$. By assumption (A3), all candidate split children satisfy minimum size requirements, ensuring there exist constants $0<c_{\min}<c_{\max}<1$ such that
+$c_{\min}\leq\frac{N_{L}(\theta)}{N_{P_{a}}},\frac{N_{R}(\theta)}{N_{P_{a}}}\leq c_{\max}.$
+Case 1: $j=j^{\prime}$ (same axis).
+Let $\delta=|t-t^{\prime}|$.
+By assumption (A2), the empirical covariate distribution satisfies a local mass bound.
+Specifically, the slab $\{x\in P_{a}:t<x_{j}\leq t^{\prime}\}$ of width $\delta$ can be covered by approximately $(D_{P_{a}}/\delta)^{p-1}$ balls of radius $\delta$, where $D_{P_{a}}:=\text{diam}(P_{a})$.
+By (A2), each ball $B_{\delta}(x)$ contains at most $M\delta^{p}N$ observations, hence:
+$|N_{L}(\theta)-N_{L}(\theta^{\prime})|\leq\frac{D_{P_{a}}^{p-1}}{\delta^{p-1}}\cdot M\delta^{p}N=MD_{P_{a}}^{p-1}\delta N.$
+Since $N_{P_{a}}=\Theta(N)$ for the parent node, we have $|N_{L}(\theta)-N_{L}(\theta^{\prime})|\leq MD_{P_{a}}^{p-1}\delta N_{P_{a}}=:L_{X}\delta N_{P_{a}},$ where $L_{X}:=MD_{P_{a}}^{p-1}$ depends on the upper density bound $M$ from (A2) and the parent diameter.
+Since $w(\theta)=N_{L}N_{R}/N_{P_{a}}^{2}$ is smooth in $(N_{L},N_{R})$ with bounded derivatives:
+$|w(\theta)-w(\theta^{\prime})|$ $\leq\frac{1}{N_{P_{a}}^{2}}\left(|N_{L}(\theta)-N_{L}(\theta^{\prime})|\cdot N_{R}(\theta)+N_{L}(\theta^{\prime})\cdot|N_{R}(\theta)-N_{R}(\theta^{\prime})|\right)$
+$\leq\frac{2c_{\max}N_{P_{a}}\cdot L_{X}\delta N_{P_{a}}}{N_{P_{a}}^{2}}=2c_{\max}L_{X}\delta.$
+Therefore $|w(\theta)-w(\theta^{\prime})|\leq C_{w}\delta=C_{w}d(\theta,\theta^{\prime}),$ where $C_{w}=2c_{\max}L_{X}=2c_{\max}MD_{P_{a}}^{p-1}$.
+Case 2: $j\neq j^{\prime}$ (different axes). Then $d(\theta,\theta^{\prime})=1$. Since $w(\theta)\in[c_{\min}^{2},c_{\max}(1-c_{\min})]$ by (A3), we have
+$|w(\theta)-w(\theta^{\prime})|\leq 2c_{\max}(1-c_{\min})=:C_{w}\cdot d(\theta,\theta^{\prime}).$
+Now we bound $|\Delta(\theta)^{2}-\Delta(\theta^{\prime})^{2}|$.
+By (K1), $k$ is uniformly bounded with $K_{\max}:=\sup_{y,y^{\prime}\in\mathcal{Y}}|k(y,y^{\prime})|<\infty$.
+Using the identity $|a^{2}-b^{2}|=|a-b|\cdot|a+b|$ and $\Delta(\theta)\leq 2\sqrt{K_{\max}}$:
+$|\Delta(\theta)^{2}-\Delta(\theta^{\prime})^{2}|\leq 2\sqrt{K_{\max}}\cdot|\Delta(\theta)-\Delta(\theta^{\prime})|.$
+By the triangle inequality:
+$|\Delta(\theta)-\Delta(\theta^{\prime})|$ $\leq\|\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{L}(\theta^{\prime}))\|_{\mathcal{H}}+\|\mu_{k}(P_{R}(\theta))-\mu_{k}(P_{R}(\theta^{\prime}))\|_{\mathcal{H}}.$
+Consider Bounding $\|\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{L}(\theta^{\prime}))\|_{\mathcal{H}}$.
+When $j=j^{\prime}$: Define $\mathcal{S}_{\Delta}=\{i:X_{i}\in C_{L}(\theta)\triangle C_{L}(\theta^{\prime})\}$ (symmetric difference).
+By the derivation above, $|\mathcal{S}_{\Delta}|\leq L_{X}\delta N_{P_{a}}$.
+Decompose:
+$\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{L}(\theta^{\prime}))$
+$=\frac{1}{N_{L}(\theta)}\sum_{X_{i}\in C_{L}(\theta)}\mu_{k}(\delta_{Y_{i}})-\frac{1}{N_{L}(\theta^{\prime})}\sum_{X_{i}\in C_{L}(\theta^{\prime})}\mu_{k}(\delta_{Y_{i}})$
+$=\frac{1}{N_{L}(\theta)}\sum_{i\in\mathcal{S}_{\Delta}\cap C_{L}(\theta)}\mu_{k}(\delta_{Y_{i}})-\frac{1}{N_{L}(\theta^{\prime})}\sum_{i\in\mathcal{S}_{\Delta}\cap C_{L}(\theta^{\prime})}\mu_{k}(\delta_{Y_{i}})$
+$\quad+\left(\frac{1}{N_{L}(\theta)}-\frac{1}{N_{L}(\theta^{\prime})}\right)\sum_{X_{i}\in C_{L}(\theta)\cap C_{L}(\theta^{\prime})}\mu_{k}(\delta_{Y_{i}}).$
+By (K1), $\|\mu_{k}(\delta_{y})\|_{\mathcal{H}}\leq\sqrt{K_{\max}}$ for all $y\in\mathcal{Y}$. Thus
+$\left\|\frac{1}{N_{L}(\theta)}\sum_{i\in\mathcal{S}_{\Delta}\cap C_{L}(\theta)}\mu_{k}(\delta_{Y_{i}})\right\|_{\mathcal{H}}\leq\frac{|\mathcal{S}_{\Delta}|}{N_{L}(\theta)}\sqrt{K_{\max}}\leq\frac{L_{X}\delta N_{P_{a}}}{c_{\min}N_{P_{a}}}\sqrt{K_{\max}}=\frac{L_{X}\sqrt{K_{\max}}}{c_{\min}}\delta.$
+Similarly for the second term. For the third term,
+$\left|\frac{1}{N_{L}(\theta)}-\frac{1}{N_{L}(\theta^{\prime})}\right|\leq\frac{|N_{L}(\theta)-N_{L}(\theta^{\prime})|}{N_{L}(\theta)N_{L}(\theta^{\prime})}\leq\frac{L_{X}\delta N_{P_{a}}}{(c_{\min}N_{P_{a}})^{2}}=\frac{L_{X}\delta}{c_{\min}^{2}N_{P_{a}}}.$
+Since $|C_{L}(\theta)\cap C_{L}(\theta^{\prime})|\leq N_{P_{a}}$:
+$\left\|\left(\frac{1}{N_{L}(\theta)}-\frac{1}{N_{L}(\theta^{\prime})}\right)\sum_{X_{i}\in C_{L}(\theta)\cap C_{L}(\theta^{\prime})}\mu_{k}(\delta_{Y_{i}})\right\|_{\mathcal{H}}\leq\frac{L_{X}\delta}{c_{\min}^{2}N_{P_{a}}}\cdot N_{P_{a}}\sqrt{K_{\max}}=\frac{L_{X}\sqrt{K_{\max}}}{c_{\min}^{2}}\delta.$
+Combining together gives
+$\|\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{L}(\theta^{\prime}))\|_{\mathcal{H}}\leq C_{\mu}\delta=C_{\mu}d(\theta,\theta^{\prime}),\text{ where }$
+$C_{\mu}=\frac{L_{X}\sqrt{K_{\max}}}{c_{\min}}\left(2+\frac{1}{c_{\min}}\right)=\frac{MD_{P_{a}}^{p-1}\sqrt{K_{\max}}}{c_{\min}}\left(2+\frac{1}{c_{\min}}\right)$
+When $j\neq j^{\prime}$: $\|\mu_{k}(P_{L}(\theta))-\mu_{k}(P_{L}(\theta^{\prime}))\|_{\mathcal{H}}\leq\|\mu_{k}(P_{L}(\theta))\|_{\mathcal{H}}+\|\mu_{k}(P_{L}(\theta^{\prime}))\|_{\mathcal{H}}\leq 2\sqrt{K_{\max}}=2\sqrt{K_{\max}}\cdot d(\theta,\theta^{\prime}).$ So $|\Delta(\theta)^{2}-\Delta(\theta^{\prime})^{2}|\leq C_{\Delta}d(\theta,\theta^{\prime})$ for $C_{\Delta}=4\sqrt{K_{\max}}\max(C_{\mu},2\sqrt{K_{\max}})$.
+Together, using $|ab-a^{\prime}b^{\prime}|\leq|a-a^{\prime}|\cdot|b|+|a^{\prime}|\cdot|b-b^{\prime}|$, we have that
+$|M^{N}(\theta)-M^{N}(\theta^{\prime})|$ $=|w(\theta)\Delta(\theta)^{2}-w(\theta^{\prime})\Delta(\theta^{\prime})^{2}|$
+$\leq|w(\theta)-w(\theta^{\prime})|\cdot\Delta(\theta)^{2}+w(\theta^{\prime})\cdot|\Delta(\theta)^{2}-\Delta(\theta^{\prime})^{2}|$
+$\leq C_{w}d(\theta,\theta^{\prime})\cdot 4K_{\max}+c_{\max}(1-c_{\min})\cdot C_{\Delta}d(\theta,\theta^{\prime})$
+$=L_{1}d(\theta,\theta^{\prime}),$
+where $L_{1}=4K_{\max}\cdot 2c_{\max}MD_{P_{a}}^{p-1}+c_{\max}(1-c_{\min})\cdot 4\sqrt{K_{\max}}\max(C_{\mu},2\sqrt{K_{\max}}).$
+Finally, we extend to design-weighted score $M_{n_{s}}(\theta)$. For the design-weighted split score, define analogously:
+$w_{n_{s}}(\theta)$ $:=\frac{N_{L}^{n_{s}}(\theta)N_{R}^{n_{s}}(\theta)}{(N_{P_{a}}^{n_{s}})^{2}},\quad\text{where }N_{\ell}^{n_{s}}(\theta):=\sum_{X_{i}\in C_{\ell}(\theta)}\frac{\xi_{i}}{\pi_{i}},$
+$\Delta_{n_{s}}(\theta)$ $:=\left\|\mu_{k}(P_{L}^{n_{s}}(\theta))-\mu_{k}(P_{R}^{n_{s}}(\theta))\right\|_{\mathcal{H}},$
+$P_{\ell}^{n_{s}}(\theta)$ $:=\frac{1}{N_{L}^{n_{s}}(\theta)}\sum_{X_{i}\in C_{L}(\theta)}\frac{\xi_{i}}{\pi_{i}}\delta_{Y_{i}},\quad\ell\in\{L,R\}.$
+The proof structure is identical, but the bounds differ due to design weights.
+For $|w_{n_{s}}(\theta)-w_{n_{s}}(\theta^{\prime})|$, By (D3), $\pi_{i}\in[\lambda n/N,\bar{\lambda}n/N]$, so each unit contributes weight within the range $N/(\bar{\lambda}n),N/(\lambda n)]$ to $N_{\ell}^{n_{s}}$.
+For $j=j^{\prime}$: The number of units changing sides is bounded by $L_{X}\delta N_{P_{a}}$, and each contributes weight at most $N/(\lambda n)$, giving
+$|N_{L}^{n_{s}}(\theta)-N_{L}^{n_{s}}(\theta^{\prime})|\leq L_{X}\delta N_{P_{a}}\cdot\frac{N}{\lambda n}.$
+By (D2), $n/N\to f\in(0,1]$, so $N/n=O(1)$. Also, by (A3) and (D3), $N_{\ell}^{n_{s}}(\theta)\geq c_{\min}N_{P_{a}}^{n_{s}}\geq c_{\min}\lambda(n/N)N_{P_{a}}$. By smoothness of $w_{n_{s}}$,
+$|w_{n_{s}}(\theta)-w_{n_{s}}(\theta^{\prime})|\leq\frac{2c_{\max}}{(N_{P_{a}}^{n_{s}})^{2}}\cdot N_{P_{a}}^{n_{s}}\cdot L_{X}\delta N_{P_{a}}\cdot\frac{N}{\lambda n}=\frac{2c_{\max}L_{X}}{N_{P_{a}}^{n_{s}}}\cdot\frac{N_{P_{a}}N}{\lambda n}\delta.$
+By (D2), $N_{P_{a}}^{n_{s}}/N_{P_{a}}\xrightarrow{p_{d}}f$, and $N/n=O(1)$, giving:
+$|w_{n_{s}}(\theta)-w_{n_{s}}(\theta^{\prime})|\leq C_{w}^{\prime}\delta=C_{w}^{\prime}d(\theta,\theta^{\prime}),\text{where}$
+$C_{w}^{\prime}=O(L_{X}c_{\max}/(\lambda f))=O(Mc_{\max}D_{P_{a}}^{p-1}/(\lambda f)).$
+For $j\neq j^{\prime}$: As before, $|w_{n_{s}}(\theta)-w_{n_{s}}(\theta^{\prime})|\leq 2c_{\max}(1-c_{\min})$.
+To bound $|\Delta_{n_{s}}(\theta)^{2}-\Delta_{n_{s}}(\theta^{\prime})^{2}|$, following previous arguments gives
+$|\Delta_{n_{s}}(\theta)^{2}-\Delta_{n_{s}}(\theta^{\prime})^{2}|\leq C_{\Delta}^{\prime}d(\theta,\theta^{\prime}),\text{ where }$
+$C_{\Delta}^{\prime}=O(\sqrt{K_{\max}}MD_{P_{a}}^{p-1}/(\lambda fc_{\min}^{2}))$
+Together,
+$|M_{n_{s}}(\theta)-M_{n_{s}}(\theta^{\prime})|\leq C_{w}^{\prime}d(\theta,\theta^{\prime})\cdot 4K_{\max}+c_{\max}(1-c_{\min})\cdot C_{\Delta}^{\prime}d(\theta,\theta^{\prime})=L_{2}d(\theta,\theta^{\prime}),\text{ where }$
+$L_{2}=O\left(\frac{MD_{P_{a}}^{p-1}K_{\max}}{\lambda fc_{\min}^{2}}\left(c_{\max}+\frac{c_{\max}(1-c_{\min})}{c_{\min}}\right)\right)$
+depends on the kernel bound $K_{\max}$ from (K1), design parameters $\underline{\lambda},\bar{\lambda},f$ from (D2)-(D3), the upper density bound $M$ from (A2), and non-degeneracy constants $c_{\min},c_{\max}$ from (A3).
+∎
+Proof of Theorem 3.2.
+Let $P_{a}\subseteq\mathcal{X}$ denote the region corresponding to a given node at depth $d$, defined by the preceding splits.
+For a fixed candidate split $\theta=(j,t)$ partitioning $P_{a}$ into child regions $C_{L}(\theta)$ and $C_{R}(\theta)$ (hereafter suppressing the dependence on $\theta$ for notational clarity).
+Consider a fixed region $P_{a}\subseteq\mathcal{X}$ which may represent the region at a node at depth $d$ in the tree-growing process.
+Part A
+We first establish pointwise design and model consistency of the MMD splitting criterion $M^{*}_{n_{s}}(\theta)$ to the finite-population score $M^{N}(\theta)$ and the super-population score $M(\theta)$ for a fixed split $\theta$.
+Under the resampling scheme described in the SDRF Algorithm, the multipliers $\{n^{*}_{b,i}\}_{i=1}^{n_{s}}$ satisfy $\mathbb{E}[n^{*}_{b,i}\mid\xi,Z^{N}]=q$ by construction, along with properties (R1)-(R3).
+For each child region $\ell\in\{L,R\}$, define the bootstrap-weighted leaf size and the corresponding finite-population leaf size respectively as
+$\hat{N}_{\ell}=\sum_{i=1}^{n_{s}}\frac{n^{*}_{b,i}}{q}\frac{\xi_{i}}{\pi_{i}}\mathbb{I}(X_{i}\in C_{\ell}),\qquad N_{\ell}^{N}=\sum_{i=1}^{N}\mathbb{I}(X_{i}\in C_{\ell}).$
+By Lemma 2.1, under assumptions (D2)-(D4), $\frac{1}{N}\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}-1\xrightarrow{p_{d}}0$.
+Combined with the centering property $\mathbb{E}[n^{*}_{b,i}/q\mid\xi,Z^{N}]=1$ from Lemma 3.1 (R2) and the weak dependence condition (R3), the law of large numbers for bootstrap-weighted sums yields
+$\frac{\hat{N}_{\ell}}{N}=\frac{1}{N}\sum_{i=1}^{n_{s}}\frac{n^{*}_{b,i}}{q}\frac{\xi_{i}}{\pi_{i}}\mathbb{I}(X_{i}\in C_{\ell})\xrightarrow{{p_{d}}}\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(X_{i}\in C_{\ell})=\frac{N_{\ell}^{N}}{N}\eqqcolon p_{\ell}^{N}.$
+Define $p_{P_{a}}^{N}=p_{L}^{N}+p_{R}^{N}$. Since $\hat{N}_{P_{a}}/N=(\hat{N}_{L}+\hat{N}_{R})/N$, we have $\hat{N}_{P_{a}}/N\xrightarrow{{p_{d}}}p_{P_{a}}^{N}$. By Slutsky’s theorem,
+$\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}_{P_{a}}^{2}}\xrightarrow{{p_{d}}}\frac{p_{L}^{N}\cdot p_{R}^{N}}{(p_{P_{a}}^{N})^{2}}.$
+For the kernel mean embedding terms, define the KME and its finite-population counterpart, respectively as
+[…]
+By assumption (K1), the kernel $k$ is uniformly bounded: $\sup_{y,y^{\prime}}|k(y,y^{\prime})|\leq K_{\max}<\infty$, which implies $\sup_{i}\|\mu_{\mathbf{k}}(\delta_{Y_{i}})\|_{\mathcal{H}}\leq K_{\max}$.
+To show $\hat{\mu}_{\ell}=\frac{1}{\hat{N}_{\ell}}\sum_{\{i:X_{i}\in C_{\ell}\}}\frac{n^{*}_{b,i}}{q\pi_{i}}\mu_{\mathbf{k}}(\delta_{Y_{i}})\xrightarrow{{p_{d}}}\mu_{\ell}^{N}$, observe that
+$\|\hat{\mu}_{\ell}-\mu_{\ell}^{N}\|_{\mathcal{H}}$ $\leq\left\|\frac{1}{\hat{N}_{\ell}}\sum_{i:X_{i}\in C_{\ell}}\frac{n^{*}_{b,i}}{q\pi_{i}}\mu_{\mathbf{k}}(\delta_{Y_{i}})-\frac{1}{N_{\ell}^{N}}\sum_{i:X_{i}\in C_{\ell}}\mu_{\mathbf{k}}(\delta_{Y_{i}})\right\|_{\mathcal{H}}$
+$\leq\left\|\frac{1}{\hat{N}_{\ell}}\sum_{i:X_{i}\in C_{\ell}}\left(\frac{n^{*}_{b,i}}{q\pi_{i}}-1\right)\mu_{\mathbf{k}}(\delta_{Y_{i}})\right\|_{\mathcal{H}}+\left\|\left(\frac{1}{\hat{N}_{\ell}}-\frac{1}{N_{\ell}^{N}}\right)\sum_{i:X_{i}\in C_{\ell}}\mu_{\mathbf{k}}(\delta_{Y_{i}})\right\|_{\mathcal{H}}$
+$\leq K_{\max}\left|\frac{1}{\hat{N}_{\ell}}\sum_{i:X_{i}\in C_{\ell}}\left(\frac{n^{*}_{b,i}}{q\pi_{i}}-1\right)\right|+K_{\max}N_{\ell}^{N}\left|\frac{1}{\hat{N}_{\ell}}-\frac{1}{N_{\ell}^{N}}\right|.$
+The first term converges to zero in $P_{\mathcal{S}^{N}|\omega}$-probability by Lemma 3.1 (properties (R1)-(R3)) combined with Lemma 2.1.
+The second term also vanishes since $\hat{N}_{\ell}/N\xrightarrow{p_{d}}p_{\ell}^{N}>0$ (on the event $\{N_{\ell}^{N}>0\}$).
+Therefore,
+$\hat{\mu}_{\ell}\xrightarrow{p_{d}}\mu_{\ell}^{N},\quad\ell\in\{L,R\}.$
+Since the squared Hilbert norm $\|\cdot\|_{\mathcal{H}}^{2}$ is continuous, the continuous mapping theorem implies
+$\|\hat{\mu}_{L}-\hat{\mu}_{R}\|_{\mathcal{H}}^{2}\xrightarrow{p_{d}}\|\mu_{L}^{N}-\mu_{R}^{N}\|_{\mathcal{H}}^{2}.$
+Combining this with the convergence of the balancing weights via Slutsky’s theorem, we obtain
+$M^{*}_{n_{s}}(\theta)=\frac{\hat{N}_{L}\hat{N}_{R}}{\hat{N}_{P_{a}}^{2}}\|\hat{\mu}_{L}-\hat{\mu}_{R}\|_{\mathcal{H}}^{2}\xrightarrow{p_{d}}M^{N}(\theta)=\frac{p_{L}^{N}\cdot p_{R}^{N}}{(p_{P_{a}}^{N})^{2}}\|\mu_{L}^{N}-\mu_{R}^{N}\|_{\mathcal{H}}^{2},$
+establishing design consistency for any fixed split $\theta$.
+To establish pointwise model consistency, we require $M^{N}(\theta)\to M(\theta)$ as $N\to\infty$, where $M(\theta)$ is the super-population splitting criterion.
+Under the super-population model with $(X_{i},Y_{i})\stackrel{{\scriptstyle\text{iid}}}{{\sim}}P$ for $i=1,\ldots,N$, the strong law of large numbers yields
+$p_{\ell}^{N}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}(X_{i}\in C_{\ell})\to p_{\ell}:=P(X\in C_{\ell})\quad\text{a.s. for $\ell\in\{L,R\}$}.$
+For the kernel mean embeddings, by (K1), $\mathbb{E}_{P}[k(Y,Y)]\leq K_{\max}^{2}<\infty$, which implies $\mathbb{E}_{P}[\|\mu_{\mathbf{k}}(\delta_{Y})\|_{\mathcal{H}}]\leq K_{\max}<\infty$.
+By the strong law of large numbers in separable Hilbert spaces, for each fixed $\ell\in\{L,R\}$:
+$\mu_{\ell}^{N}=\frac{1}{N_{\ell}^{N}}\sum_{\{i:X_{i}\in C_{\ell}\}}\mu_{\mathbf{k}}(\delta_{Y_{i}})\to\mathbb{E}_{P}[\mu_{\mathbf{k}}(\delta_{Y})\mid X\in C_{\ell}]=:\mu_{\ell}\quad\text{in }\mathcal{H}\text{ a.s.}$
+By the continuous mapping theorem applied to the function $(p_{L},p_{R},\mu_{L},\mu_{R})\mapsto\frac{p_{L}p_{R}}{(p_{L}+p_{R})^{2}}\|\mu_{L}-\mu_{R}\|_{\mathcal{H}}^{2}$, we obtain
+$M^{N}(\theta)\to M(\theta)=\frac{p_{L}\cdot p_{R}}{(p_{L}+p_{R})^{2}}\|\mu_{L}-\mu_{R}\|_{\mathcal{H}}^{2}\quad\text{a.s.}$
+Combining the two stages, we conclude that for each fixed $\theta$,
+$M^{*}_{n_{s}}(\theta)\xrightarrow{p_{\otimes}}M(\theta)\quad\text{as }N\to\infty.$
+Part B
+We now establish $\sup_{\theta\in\Pi}|M^{*}_{n_{s}}(\theta)-M(\theta)|=o_{p_{\otimes}}(1)$.
+Denote $\Pi:=\{\theta=(j,t):j\in[p],t\in[a_{j},b_{j}]\}$ the class of axis-aligned splits in the bounded set $\mathcal{X}\subset\mathbb{R}^{p}$.
+Then $\Pi$ is a compact metric space under $d((j,t),(j^{\prime},t^{\prime}))=\mathbbm{1}(j\neq j^{\prime})+|t-t^{\prime}|$.
+For any $\theta=(j,t)\in\Pi$, define the child regions: $C_{L}(\theta)=\{x\in\mathbb{R}^{p}:x_{j}\leq t\},\quad C_{R}(\theta)=\{x\in\mathbb{R}^{p}:x_{j}>t\}.$ Define the primitive class of axis-aligned halfspaces: $\mathcal{A}:=\{C_{L}(\theta),C_{R}(\theta):\theta\in\Pi\}.$ Since each element of $\mathcal{A}$ is a halfspace in $\mathbb{R}^{p}$, $\mathcal{A}$ is a VC class with $\mathrm{VC}(\mathcal{A})\leq p+1$.
+Since parent regions at depth $d$ are formed by intersecting up to $d$ halfspaces from ancestor splits, we define:
+$\mathcal{C}_{d}$ $:=\left\{\bigcap_{\ell=1}^{m}C^{\ell}(\theta^{\ell}):1\leq m\leq d,\ \theta^{\ell}\in\Pi,\ C^{\ell}(\theta^{\ell})\in\{C_{L}(\theta^{\ell}),C_{R}(\theta^{\ell})\}\right\}$
+$\subseteq\mathcal{A}^{\cap d}:=\left\{\bigcap_{\ell=1}^{m}A_{\ell}:1\leq m\leq d,\ A_{\ell}\in\mathcal{A}\right\}.$
+By the closure bound for $d$-fold intersections of VC classes, $\mathrm{VC}(\mathcal{C}_{d})\leq\mathrm{VC}(\mathcal{A}^{\cap d})=O(\mathrm{VC}(\mathcal{A})\cdot d\log d)=O(pd\log d)<\infty.$ Define an augmented function class:
+$\mathcal{H}_{d}:=\mathbbm{1}(X\in L):L\in\mathcal{C}_{d}\}\cup\{k(Y,y)\cdot\mathbbm{1}(X\in L):L\in\mathcal{C}_{d},\ y\in\mathcal{Y}\}.$
+By (K1), $\sup_{y,y^{\prime}}|k(y,y^{\prime})|=:K_{\max}<\infty$, so $\mathcal{H}_{d}$ has envelope $H\equiv K_{\max}$ and inherits the VC structure from $\mathcal{C}_{d}$.
+Specifically, $\mathcal{H}_{d}$ is a VC-subgraph class with $\mathrm{VC}(\mathcal{H}_{d})=O(pd\log d)$ and square-integrable envelope.
+Therefore, for any fixed $d$, $\mathcal{H}_{d}$ satisfies the uniform entropy condition
+$\int_{0}^{\infty}\sup_{Q}\sqrt{\log N(\epsilon\|H\|_{Q,2},\mathcal{H}_{d},L_{2}(Q))}\,d\epsilon<\infty,$
+where the supremum is over all finitely discrete probability measures $Q$ (Theorem 2.6.7, [van-der-Vaart2023-ur]).
+For any parent region $P_{a}\in\mathcal{C}_{d-1}$, define the global empirical measures, for $g\in\mathcal{H}_{d}$:
+$P(g)$ $:=\mathbb{E}_{\omega}[g(X,Y)],$
+$P_{N}(g)$ $:=\frac{1}{N}\sum_{i=1}^{N}g(X_{i},Y_{i}),$
+$P^{n_{s}}_{N}(g)$ $:=\frac{1}{n_{s}}\sum_{i=1}^{N}\xi_{i}g(X_{i},Y_{i}),\quad n_{s}=\sum_{i=1}^{N}\xi_{i},$
+$P^{*}_{N}(g)$ $:=\frac{1}{n^{*}_{s}}\sum_{i=1}^{N}n^{*}_{i}\xi_{i}g(X_{i},Y_{i}),\quad n^{*}_{s}=\sum_{i=1}^{N}n^{*}_{i}\xi_{i}.$
+For a fixed parent $P_{a}$, the conditional empirical measures are:
+$P_{N|P_{a}}(g)$ $:=\frac{P_{N}(\mathbbm{1}_{P_{a}}\cdot g)}{P_{N}(\mathbbm{1}_{P_{a}})}=\frac{1}{N_{P_{a}}}\sum_{i:X_{i}\in P_{a}}g(X_{i},Y_{i}),$
+$P^{n_{s}}_{N|P_{a}}(g)$ $:=\frac{P^{n_{s}}_{N}(\mathbbm{1}_{P_{a}}\cdot g)}{P^{n_{s}}_{N}(\mathbbm{1}_{P_{a}})}=\frac{1}{N^{n_{s}}_{P_{a}}}\sum_{i:X_{i}\in P_{a}}\frac{\xi_{i}}{\pi_{i}}g(X_{i},Y_{i}),$
+$P^{*}_{N|P_{a}}(g)$ $:=\frac{P^{*}_{N}(\mathbbm{1}_{P_{a}}\cdot g)}{P^{*}_{N}(\mathbbm{1}_{P_{a}})}=\frac{1}{N^{*}_{P_{a}}}\sum_{i:X_{i}\in P_{a}}\frac{n^{*}_{i}\xi_{i}}{\pi_{i}}g(X_{i},Y_{i}).$
+For any split $\theta\in\Pi$ and parent $P_{a}\in\mathcal{C}_{d-1}$, the split scores can be expressed as:
+$M(\theta)$ $=\Phi_{\theta}\left(P(g_{1}^{\theta}),\ldots,P(g_{m}^{\theta})\right),$ (S15)
+$M^{N}_{P_{a}}(\theta)$ $=\Phi_{\theta}\left(P_{N|P_{a}}(g_{1}^{\theta}),\ldots,P_{N|P_{a}}(g_{m}^{\theta})\right),$ (S16)
+$M^{n_{s}}_{P_{a}}(\theta)$ $=\Phi_{\theta}\left(P^{n_{s}}_{N|P_{a}}(g_{1}^{\theta}),\ldots,P^{n_{s}}_{N|P_{a}}(g_{m}^{\theta})\right),$ (S17)
+$M^{*}_{n_{s},P_{a}}(\theta)$ $=\Phi_{\theta}\left(P^{*}_{N|P_{a}}(g_{1}^{\theta}),\ldots,P^{*}_{N|P_{a}}(g_{m}^{\theta})\right),$ (S18)
+where $g_{1}^{\theta},\ldots,g_{m}^{\theta}\in\mathcal{H}_{d}$ are the building-block functions:
+$\mathcal{G}_{\theta}:=\mathbbm{1}(X\in C_{L}(\theta)),\mathbbm{1}(X\in C_{R}(\theta))\}\cup\{k(Y,Y_{i})\cdot\mathbbm{1}(X\in C_{\ell}(\theta)):i\in[N],\ell\in\{L,R\}\},$
+and $\Phi_{\theta}:\mathbb{R}^{m}\to\mathbb{R}$ is a bounded Lipschitz function. Specifically, since
+$M^{N}_{P_{a}}(\theta)=\frac{N_{L}(\theta)N_{R}(\theta)}{N_{P_{a}}^{2}}\left\|\mu_{k}(P_{L|P_{a}})-\mu_{k}(P_{R|P_{a}})\right\|^{2}_{\mathcal{H}},$
+where $N_{\ell}(\theta)/N_{P_{a}}=P_{N|P_{a}}(\mathbbm{1}(X\in C_{\ell}(\theta)))$ for $\ell\in\{L,R\}$, and
+$\|\mu_{k}(P_{L})-\mu_{k}(P_{R})\|^{2}_{\mathcal{H}}$ $=\langle\mu_{k}(P_{L}),\mu_{k}(P_{L})\rangle_{\mathcal{H}}+\langle\mu_{k}(P_{R}),\mu_{k}(P_{R})\rangle_{\mathcal{H}}-2\langle\mu_{k}(P_{L}),\mu_{k}(P_{R})\rangle_{\mathcal{H}}$
+$=P_{L}\times P_{L}(k)+P_{R}\times P_{R}(k)-2P_{L}\times P_{R}(k),$
+each component is a polynomial in empirical averages of functions in $\mathcal{H}_{d}$.
+By assumption (A3), parent regions satisfy non-degeneracy: there exist $0<c_{\min}<c_{\max}<1$ such that $c_{\min}\leq N_{P_{a}}/N\leq c_{\max}$ for all $P_{a}\in\mathcal{C}_{d-1}$ considered by the algorithm, ensuring the denominators remain bounded away from zero.
+Therefore, $\Phi_{\theta}$ is Lipschitz continuous with constant $C$ depending on $K_{\max}$ (from (K1)), $c_{\min},c_{\max}$ (from (A3)), and thus for any pair of measures $Q_{1},Q_{2}$:
+$\sup_{\theta\in\Pi}|M_{Q_{1}}(\theta)-M_{Q_{2}}(\theta)|\leq C\sup_{g\in\mathcal{H}_{d}}|Q_{1}(g)-Q_{2}(g)|,$ (S19)
+where $M_{Q}(\theta):=\Phi_{\theta}(\{Q(g)\}_{g\in\mathcal{G}_{\theta}})$.
+As $\sup_{g\in\mathcal{H}_{d}}|(P^{*}_{N}-P)(g)|\leq\sup_{g\in\mathcal{H}_{d}}|(P^{*}_{N}-P^{n_{s}}_{N})(g)|+\sup_{g\in\mathcal{H}_{d}}|(P^{n_{s}}_{N}-P_{N})(g)|+\sup_{g\in\mathcal{H}_{d}}|(P_{N}-P)(g)|,$ we now show for the three parts separately to complete proof for Part B.
+First, since $\mathcal{H}_{d}$ is $P$-Glivenko-Cantelli (by Step 1 and Theorem 2.4.1, [Kosorok2008-xs]): $\sup_{g\in\mathcal{H}_{d}}|(P_{N}-P)(g)|\xrightarrow{p_{\omega}}0.$ For the conditional measures, by the ratio representation and (A3).
+$\sup_{g\in\mathcal{H}_{d}}|P_{N|P_{a}}(g)-P_{|P_{a}}(g)|\leq\frac{1}{c_{\min}}\sup_{g\in\mathcal{H}_{d}}|(P_{N}-P)(g)|\xrightarrow{p_{\omega}}0,$ where $P_{|P_{a}}(g):=\mathbb{E}_{\omega}[g(X,Y)|X\in P_{a}]$.
+So by (S19)
+$\sup_{P_{a}\in\mathcal{C}_{d-1}}\sup_{\theta\in\Pi}|M^{N}_{P_{a}}(\theta)-M(\theta)|\leq C\sup_{g\in\mathcal{H}_{d}}|P_{N|P_{a}}(g)-P_{|P_{a}}(g)|\xrightarrow{p_{\omega}}0.$ (S20)
+Then, for any $g\in\mathcal{H}_{d}$, decompose $P^{n_{s}}_{N}(g)-P_{N}(g)=R_{N}(g)+S_{N}(g)$ where
+$R_{N}(g):=\left(\frac{1}{n_{s}}-\frac{1}{N}\right)\sum_{i=1}^{N}\xi_{i}g(X_{i},Y_{i}),\quad S_{N}(g):=\frac{1}{N}\sum_{i=1}^{N}(\xi_{i}-\pi_{i})g(X_{i},Y_{i}).$
+By (D2), $n_{s}/N\xrightarrow{p_{d}}f\in(0,1]$, hence $|R_{N}(g)|\leq K_{\max}|n_{s}/N-1|\xrightarrow{p_{d}}0$, so $\sup_{g\in\mathcal{H}_{d}}|R_{N}(g)|=o_{p_{d}}(1)$.
+For $S_{N}(g)$, by (D2)-(D4) and the design-based maximal inequality (Lemma 2.1 with Theorem 2.6.7, [van-der-Vaart2023-ur]),
+$\mathbb{E}_{\xi|\omega}\left[\sup_{g\in\mathcal{H}_{d}}|S_{N}(g)|\right]=O\left(\sqrt{\frac{\mathrm{VC}(\mathcal{H}_{d})\log N}{N}}\right)=O\left(\sqrt{\frac{pd\log d\cdot\log N}{N}}\right).$
+By Markov’s inequality, $\sup_{g\in\mathcal{H}_{d}}|S_{N}(g)|=o_{p_{d}}(1)$.
+Therefore $\sup_{g\in\mathcal{H}_{d}}|P^{n_{s}}_{N}(g)-P_{N}(g)|=o_{p_{d}}(1)$.
+By the ratio representation and (A3), $\sup_{g\in\mathcal{H}_{d}}|P^{n_{s}}_{N|P_{a}}(g)-P_{N|P_{a}}(g)|\leq\frac{1}{c_{\min}}\sup_{g\in\mathcal{H}_{d}}|P^{n_{s}}_{N}(g)-P_{N}(g)|+O_{p_{d}}(|n_{s}/N-1|)=o_{p_{d}}(1).$ Thus by (S19) again,
+$\sup_{P_{a}\in\mathcal{C}_{d-1}}\sup_{\theta\in\Pi}|M^{n_{s}}_{P_{a}}(\theta)-M^{N}_{P_{a}}(\theta)|\leq C\sup_{g\in\mathcal{H}_{d}}|P^{n_{s}}_{N|P_{a}}(g)-P_{N|P_{a}}(g)|=o_{p_{d}}(1).$ (S21)
+Now consider the part $(P^{*}_{N}-P^{n_{s}}_{N})$ regarding resampling.
+By (R1)-(R3), the multipliers $\{n^{*}_{i}\}_{i=1}^{N}$ satisfy $\mathbb{E}(n^{*}_{i}|\xi,Z^{N})=1$ with appropriate moment bounds.
+Following the same previous argument, $\sup_{g\in\mathcal{H}_{d}}|P^{*}_{N}(g)-P^{n_{s}}_{N}(g)|=o_{p_{d}}(1).$ By the ratio representation, (A3), and (S19),
+$\sup_{P_{a}\in\mathcal{C}_{d-1}}\sup_{\theta\in\Pi}|M^{*}_{n_{s},P_{a}}(\theta)-M^{n_{s}}_{P_{a}}(\theta)|\leq C\sup_{g\in\mathcal{H}_{d}}|P^{*}_{N|P_{a}}(g)-P^{n_{s}}_{N|P_{a}}(g)|=o_{p_{d}}(1).$ (S22)
+Finally, by the triangle inequality and (S22), (S21), and (S20),
+$\sup_{\theta\in\Pi}|M^{*}_{n_{s}}(\theta)-M(\theta)|$ $\leq\sup_{P_{a}\in\mathcal{C}_{d-1}}\sup_{\theta\in\Pi}\Big[|M^{*}_{n_{s},P_{a}}(\theta)-M^{n_{s}}_{P_{a}}(\theta)|$
+$\qquad+|M^{n_{s}}_{P_{a}}(\theta)-M^{N}_{P_{a}}(\theta)|+|M^{N}_{P_{a}}(\theta)-M(\theta)|\Big]$
+$=o_{p_{d}}(1)+o_{p_{d}}(1)+o_{p_{\omega}}(1)=o_{p_{\otimes}}(1),$
+completing the proof for part B.
+Part C
+We now combine previous result with some additional regularization assumptions to show design and model consistency in the split $\hat{\theta}_{n_{s}}$.
+Suppose $M^{N}(\theta)$ the finite population score has a unique maximizer $\theta_{0}$.
+Denote $U$ the selection random variable for mtry possible candidate axis.
+Suppose also that for $dist(\theta_{0},\Pi(U)):=\inf_{\theta\in\Pi(U)}d(\theta_{0},\theta),$ $\mathbb{E}_{U}[dist(\theta_{0},\Pi(U))]\rightarrow 0$.
+By Lemma S5, $M^{N}$ satisfies the modulus of continuity condition.
+That is, there exists a deterministic non-decreasing function $g(\cdot)$ with $g(t)\to 0$ as $t\downarrow 0$, such that $|M^{N}(\theta)-M^{N}(\theta^{\prime})|\leq g(d(\theta,\theta^{\prime})),\forall\theta,\theta^{\prime}\in\Pi.$ We need to show that $\hat{\theta}_{n_{s}}=\operatorname*{arg\,max}_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$, $M^{*}_{n_{s}}(\hat{\theta}_{n_{s}})\geq\sup_{\theta\in\Pi}M^{*}_{n_{s}}(\theta)-\;o_{p_{d}}(1)$.
+Denote $\delta_{n}:=\sup_{\theta\in\Pi}|M^{*}_{n_{s}}(\theta)-M^{N}(\theta)|$, for which we have shown $\delta_{n}=o_{p_{d}}(1)$ in Part B.
+For any realization of $U$, let $\hat{\theta}_{n_{s}}\in\arg\max_{\theta\in\Pi(U)}M^{*}_{n_{s}}(\theta)$.
+Since $\Pi(U)\subseteq\Pi$, $\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)\leq\sup_{\theta\in\Pi}M^{N}(\theta)+\delta_{n}=M^{N}(\theta_{0})+\delta_{n},$ and $M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})\geq\sup_{\theta\in\Pi(U)}M^{N}(\theta)-\delta_{n}.$ Since $\Pi(U)$ is a non-empty closed subset of the compact set $\Pi$, there exists $\tilde{\theta}_{U}\in\Pi(U)$ such that $d(\theta_{0},\tilde{\theta}_{U})=\text{dist}(\theta_{0},\Pi(U))$.
+Hence $\sup_{\theta\in\Pi(U)}M^{N}(\theta)\geq M^{N}(\tilde{\theta}_{U})$, and by modulus of continuity,
+$M^{N}(\theta_{0})-\sup_{\theta\in\Pi(U)}M^{N}(\theta)\leq M^{N}(\theta_{0})-M^{N}(\tilde{\theta}_{U})\leq g(\text{dist}(\theta_{0},\Pi(U))).$
+Combining these inequalities yields
+$\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)-M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})\leq\big[M^{N}(\theta_{0})-\sup_{\theta\in\Pi(U)}M^{N}(\theta)\big]+2\delta_{n}\leq g(\text{dist}(\theta_{0},\Pi(U)))+2\delta_{n}.$
+Conditional on the finite population, taking expectations over $U$ while holding the design resample fixed gives, $\mathbb{E}_{U}\left[\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)-M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})\right]\leq\mathbb{E}_{U}\big[g(\text{dist}(\theta_{0},\Pi(U)))\big]+2\delta_{n}.$ By assumption, $\mathbb{E}_{U}[\text{dist}(\theta_{0},\Pi(U))]\to 0$.
+Since $g$ is continuous at 0 with $g(0)=0$ and non-decreasing, and $\text{dist}(\theta_{0},\Pi(U))\in[0,\text{diam}(\Pi)]$ is bounded, the dominated convergence theorem gives $\mathbb{E}_{U}[g(\text{dist}(\theta_{0},\Pi(U)))]\to 0$.
+Since $\delta_{n}=o_{p_{d}}(1)$, for any $\epsilon>0$, $P_{\mathcal{S}^{N}|\omega}(\mathbb{E}_{U}[\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)$- $M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})]>\epsilon)\to 0.$ For fixed design resample, $\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)-M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})\geq 0$, so by Markov’s inequality,
+$P_{\mathcal{S}^{N}|\omega}\left(P_{U}\left(\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)-M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})>\epsilon\right)>\epsilon\right)\to 0.$
+This establishes that, jointly over $(P_{\mathcal{S}^{N}|\omega},U)$, $M_{n_{s}}^{*}(\hat{\theta}_{n_{s}})\geq\sup_{\theta\in\Pi}M_{n_{s}}^{*}(\theta)-o_{p_{d}\otimes p_{U}}(1).$ By Theorem 3.2.2 of [van-der-Vaart2023-ur], since $\sup_{\theta\in\Pi}|M_{n_{s}}^{*}(\theta)-M^{N}(\theta)|\to_{p_{d}}0$ and $M^{N}$ has a unique maximizer $\theta_{0}$, we conclude $d(\hat{\theta}_{n_{s}},\theta_{0})\to_{p_{d}\otimes p_{U}}0.$ Integrating out $U$ gives $\hat{\theta}_{n_{s}}\to_{p_{d}}\theta_{0}$.
+Under the super-population model, by Part B we have $\sup_{\theta\in\Pi}|M_{n_{s}}^{*}(\theta)-M(\theta)|\to_{p_{\otimes}}0$, where $M(\theta)$ is the population score with unique maximizer $\theta^{\dagger}$.
+The same approximate optimality argument applied with $(p_{\otimes},M,\theta^{\dagger})$ in place of $(p_{d},M^{N},\theta_{0})$ yields $\hat{\theta}_{n_{s}}\to_{p_{\otimes}}\theta^{\dagger}$.
+∎
+Remark.
+Note that for the consistency in the selected split to hold, we need that $\mathbb{E}[dist(\theta_{0},\Pi(U))]\to 0$, meaning that on average, over random draws of $U$, the class of possible splits gets arbitrarily close to $\theta_{0}$.
+It is okay to have $U$ that draws from features $\{1,\dots,p\}$ that fails to include the true optimal $j_{0}$.
+For finite feature set and random sampling of $U$, $\mathbb{E}[dist(\theta_{0},\Pi(U))]\leq cP(j_{0}\not\in U)=c(1-\mathrm{mtry}/p)$, $0<c<\infty$, so a sufficient condition is to grow mtry so that $\mathrm{mtry}/p\to 1$.
+We also note that for consistency of the forest, which we prove latter, $\mathrm{mtry}$ need not grow with $N$.
+This condition is only needed for the local leaf to match the truth, where by truth we mean the split if the entire finite population data is available.
+S6.3 Effect of multiplier averaging
+Although design-aware re-sampling allows multiple trees, which enables bagging, the resampled weights $n^{*}$ might result in large variation of $\hat{w}_{i}=1/\pi_{i}\times n^{*}$.
+Averaging multiple conditional independent realizations of $n^{*}$: $\hat{w}_{i}=1/\pi_{i}\times\bar{n}^{*}_{b,i}$, where $\bar{n}^{*}_{b,i}$ the average of $M$ realizations, is a simple and direct technique for stabilizing the resulting weights.
+The price to pay is the reduction in diversity between trees.
+We analyze the effect of $M$ in the following.
+Lemma S6.
+Conditional on $\xi$, multipliers $\delta_{i}=n^{*}_{i}-1\in\Gamma(\nu,c)$, $\nu>0,c\geq 0$, for all $i$, are sub-gamma-centered random variables for all the resampling schemes described in Lemma 3.1, with $Cov(\delta_{i},\delta_{j}\mid\xi)<0$.
+Consequently, when $M>1$, $\delta_{i}=1/M\sum_{b=1}^{M}n^{*}_{b,i}-1$, $\delta_{i}\in\Gamma(\nu/M,c/M)$.
+Proof of Lemma S6.
+For the Poisson design,
+$\mathbb{E}\!\left[e^{\lambda\delta_{i}}\mid\xi\right]=e^{-\lambda}\,\mathbb{E}\!\left[\mathbb{E}\!\left[e^{\lambda n_{i}^{*}}\mid N_{i}^{*}\right]\Big|\,\xi\right]=e^{-\lambda}\,\mathbb{E}\!\left[(1-\pi_{i}+\pi_{i}e^{\lambda})^{N_{i}^{*}}\Big|\,\xi\right]$
+$=$ $e^{-\lambda}\left(1-w_{i}^{*}+w_{i}^{*}(1-\pi_{i}+\pi_{i}e^{\lambda})\right)^{\widehat{N}}=e^{-\lambda}\left(1+w_{i}^{*}\pi_{i}(e^{\lambda}-1)\right)^{\widehat{N}}=e^{-\lambda}\left(1+\frac{e^{\lambda}-1}{\widehat{N}}\right)^{\widehat{N}}$,
+where $w^{*}_{i}=(1/\pi)/\sum(1/\pi_{i})$, $\widehat{N}=\sum_{i}1/\pi_{i}$ fixed.
+Using $(1+(e^{\lambda}-1)/\widehat{N})^{\widehat{N}}\leq e^{e^{\lambda}-1}$ and $e^{\lambda}-1-\lambda\leq\frac{\lambda^{2}}{2(1-\lambda)}\quad\text{for}\lambda\in(0,1)$, we have $\log\mathbb{E}\!\left[e^{\lambda\delta_{i}}\mid\xi\right]\leq\frac{\lambda^{2}}{2(1-|\lambda|)}\quad\text{for }|\lambda|<1$, implying $\delta_{i}$ sub-gamma with variance factor $\nu=1$ and scale $c=1$, i.e., sub-exponential with parameters $(\nu^{2}=2,\,c=1)$.
+Also, $\mathbb{E}[n^{*}_{i}\mid\xi]=\mathbb{E}[N^{*}_{i}\pi_{i}\mid\xi]=1$.
+For SRSWOR, $n_{s}$ is non-random.
+For any $i\in[n_{s}]$, $N^{*}_{i}\sim Binomial(N,1/n_{s})$.
+Denote $Z_{j}$ indicator for if the jthj-th copy, $j\in[N]$, is from subject $i\in[n_{s}]$.
+Equivalently, $N^{*}_{i}=\sum_{j\in[N]}Z_{j}$.
+When a uniformly random subset from $[N]$ of size $n_{s}$ is picked, $n^{*}_{i}=\sum_{j\in[N]}Z_{j}\stackrel{{\scriptstyle d}}{{=}}\sum_{j=1}^{n_{s}}Z_{j}\sim Binomial(n_{s},1/n_{s})$.
+Thus $n^{*}_{i}\mid\xi\sim Binomial(n_{s},1/n_{s})$, with $\mathbb{E}[n^{*}_{i}\mid\xi]=1$, $Cov[n^{*}_{i},n^{*}_{j}\mid\xi]=Cov(\mathbb{E}[n^{*}_{i}|N^{*},\xi],\mathbb{E}[n^{*}_{j}|N^{*},\xi])+\mathbb{E}[Cov(n^{*}_{i},n^{*}_{j}|N^{*},\xi)]=0+\frac{n_{s}(N-n_{s})}{N^{2}(N-1)}<0$, and from Bernstein lemma for bounded random variable $Z_{j}\in[0,1]$,
+$\mathbb{E}[e^{\lambda\delta_{i}}\mid\xi]=\big(\mathbb{E}e^{\lambda(Z_{1}-\frac{1}{n_{s}})}\big)^{n_{s}}=\left(\frac{1}{n_{s}}e^{\lambda(1-\frac{1}{n_{s}})}+\left(1-\frac{1}{n_{s}}\right)e^{-\lambda/n_{s}}\right)^{n_{s}}\leq exp\Big(\frac{\lambda^{2}(1-\tfrac{1}{n_{s}})}{2(1-|\lambda|/3)}\Big),\quad\text{$|\lambda|<3$},$
+implying $\delta_{i}\in\Gamma(\mu=1-1/n_{s},c=1/3)\subset\Gamma(\mu=1,c=1)$.
+For a stratified cluster design, note that sampling and resampling across strata are independent and follow the same structure; so it suffices to focus on an arbitrary stratum.
+As above, multinomial draws within strata produce $Cov(n^{*}_{hk},\,n^{*}_{hk^{\prime}}\mid\xi)<0$.
+∎
+Proof of Proposition 3.3.
+How large $M$ is transfers directly to the concentration of $M^{*}_{n_{s}}$ to
+$M_{n_{s}}=\frac{{N}^{n_{s}}_{L}{N}^{n_{s}}_{R}}{{N}^{{}^{n_{s}}\,2}_{P_{a}}}\left\|\mu_{\mathbf{k}}\left(\frac{1}{{N}^{n_{s}}_{L}}\sum_{X_{i}\in C_{L},i\in S}\frac{1}{\pi_{i}}\;\delta_{Y_{i}}\right)-\mu_{\mathbf{k}}\left(\frac{1}{{N}^{n_{s}}_{R}}\sum_{X_{i}\in C_{R},i\in S}\frac{1}{\pi_{i}}\;\delta_{Y_{i}}\right)\right\|^{2}_{\mathcal{H}}.$
+The mapping $\bar{n}^{*}_{b}\mapsto M^{*}_{n_{s}}$ is Lipschitz as it is a composition of Lipschitz maps.
+Denote $\delta_{i}=\bar{n}^{*}_{i}-1$, $c_{L,i}=\frac{1/\pi_{i}}{\sum_{\{j\in S:X_{j}\in C_{L}\}}1/\pi_{j}}$, then
+$\Big\|\mu_{\mathbf{k}}({P}^{*}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{L})\Big\|=\Big\|\sum_{i\in C_{L}}\left(\frac{(1+\delta_{i})/\pi_{i}}{\sum_{j\in C_{L}}(1+\delta_{j})/\pi_{j}}-\frac{1/\pi_{i}}{\sum_{j\in C_{L}}1/\pi_{j}}\right)\mathbf{k}(Y_{i},\cdot)\Big\|$
+$\leq$ $\sqrt{K}\sum_{i\in C_{L}}\left(\frac{1/\pi_{i}}{\sum 1/\pi_{j}}\left|\delta_{i}-\bar{\delta}_{L}\right|+|R_{i}|\right)\leq\sqrt{K}\left(\frac{|\bar{\delta}_{L}|}{1-|\bar{\delta}_{L}|}+1\right)\frac{1}{\sqrt{n_{\text{eff},L}}}\left\|{\delta}_{L}-\bar{\delta}_{L}\mathbbm{1}\right\|_{2},$
+where $\bar{\delta}_{L}=\sum_{i}c_{L,i}\delta_{i}$, $R_{i}=-c_{L,i}\,(\delta_{i}-\bar{\delta}_{L})\,\frac{\bar{\delta}_{L}}{1+\bar{\delta}_{L}}$ so $|R_{i}|\leq\frac{|\bar{\delta}_{L}|}{1-|\bar{\delta}_{L}|}(c_{L,i}|\delta_{i}-\bar{\delta}_{L}|)$ and $\sum_{i}1/c^{2}_{L,i}=n_{\text{eff},L}=\frac{(\sum_{j}1/\pi_{j})^{2}}{\sum_{i}(1/\pi_{i})^{2}}$.
+A similar inequality holds for $\|\mu({P}^{*}_{R})-\mu(P^{n_{s}}_{R})\|$.
+Consider
+$\begin{split}&|M^{*}-M_{n_{s}}|\leq\Big|\frac{\hat{N}^{*}_{L}\hat{N}^{*}_{R}}{\hat{N}^{*\;2}_{P_{a}}}-\frac{{N}^{n_{s}}_{L}{N}^{n_{s}}_{R}}{{N}^{n_{s}\,2}_{P_{a}}}\Big|\|\mu_{\mathbf{k}}(P^{n_{s}}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{R})\|^{2}\\ +&\frac{{N}^{n_{s}}_{L}{N}^{n_{s}}_{R}}{{N}^{n_{s}\,2}_{P_{a}}}\left(\|\mu_{\mathbf{k}}(P^{n_{s}}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{R})\|^{2}-\|\mu_{\mathbf{k}}(P^{*}_{L})-\mu_{\mathbf{k}}(P^{*}_{R})\|^{2}\right).\end{split}$ (S23)
+By $S=N_{L}N_{R}/N_{P_{a}}$ a smooth function of $(N_{L},N_{R})$ with $|\partial S/\partial N_{L}|\leq N_{R}/N^{2}_{P_{a}}$, $|\partial S/\partial N_{R}|\leq N_{L}/N^{2}_{P_{a}}$, and $S\leq 1/4$, $|\frac{\hat{N}^{*}_{L}\hat{N}^{*}_{R}}{\hat{N}^{*\,2}_{P_{a}}}-\frac{{N}^{n_{s}}_{L}{N}^{n_{s}}_{R}}{{N}^{n_{s}\,2}_{P_{a}}}|\leq 1/4(|\bar{\delta}_{L}|+|\bar{\delta}_{R}|)+C(\bar{\delta}^{2}_{L}+\bar{\delta}^{2}_{R}),C<\infty,$ with the second order term of higher order.
+Also, apply that for $a,b>0$, which is true by regularity of child node size, $\|a\|^{2}-\|b\|^{2}\leq\|a-b\|(\|a-b\|+2\|b\|)\leq(\|b\|+1)(2\|a-b\|+\|a-b\|^{2})$, where $\|a-b\|\leq\|\mu({P}^{*}_{R})-\mu(P^{n_{s}}_{R})\|+\|\mu({P}^{*}_{R})-\mu(P^{n_{s}}_{R})\|$, so
+$|M_{n_{s}}^{*}-M_{n_{s}}|\lesssim\left(\sum_{l\in\{L,R\}\}}|\bar{\delta}_{l}|\right)\|\mu_{\mathbf{k}}(P^{n_{s}}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{R})\|^{2}$
+$+$ $\frac{\sqrt{K}}{2}(\|\mu_{\mathbf{k}}(P^{n_{s}}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{R})\|+1)\sum_{l\in\{L,R\}}\left(\frac{|\bar{\delta}_{l}|}{|1+\bar{\delta}_{l}|}+1\right)\frac{1}{\sqrt{n_{\text{eff},l}}}\left\|{\delta}_{l}-\bar{\delta}_{l}\mathbbm{1}\right\|_{2}.$
+$\delta_{i}$ are mean-centered sub-gamma $\Gamma(\nu,c)$ with $Cov(n^{*}_{i},n^{*}_{j}\mid\xi)\leq 0$, implying $\mathbb{E}[e^{\lambda\sum_{i}c_{i}n^{*}_{i}}\mid\xi]\leq\prod_{i}\mathbb{E}[e^{\lambda c_{i}n^{*}_{i}}\mid\xi]$ for $\lambda\geq 0$, $c_{min}<c_{l,i}=\frac{1/\pi_{i}}{\sum_{j}1/\pi_{j}}<1$.
+Thus $\bar{\delta}_{l}$ is also sub-gamma with $\nu_{l}=(\nu/M)\sum_{i}c_{l,i}^{2}\leq(\frac{\bar{\lambda}}{\underline{\lambda}})^{4}\nu n^{-1}_{\mathrm{eff},l}M^{-1}$, $c_{l}=\frac{c}{n_{l}M}\max_{i}|c_{l,i}|\leq(\frac{\bar{\lambda}}{\underline{\lambda}})^{2}c(n_{\mathrm{eff},l}M)^{-1}$, using $n_{l}=\Theta(n_{\mathrm{eff},l})$, and $P(|\bar{\delta}_{l}|\geq\sqrt{2\nu t}+ct)\leq 2e^{-t},\,t>0,$ implying $\sum_{\ell}\lvert\bar{\delta}_{\ell}\rvert=O_{p}\!\left(\sum_{\ell}\sqrt{\nu_{l}}+c_{l}\right)=O_{p}\!\left(\sum_{\ell}n_{\mathrm{eff},\ell}^{-1/2}M^{-1/2}\right).$ Bounded kernel implies $\|\mu_{\mathbf{k}}(P^{n_{s}}_{L})-\mu_{\mathbf{k}}(P^{n_{s}}_{R})\|\leq 2\sqrt{K}$.
+Also, $\frac{|\bar{\delta}_{L}|}{|1+\bar{\delta}_{L}|}+1\leq 1+2|\bar{\delta}_{L}|+2\mathbbm{1}\{|\bar{\delta}_{L}|>\frac{1}{2}\},$ so $\frac{|\bar{\delta}_{L}|}{|1+\bar{\delta}_{L}|}+1=1+o_{p}(1)$.
+By $\delta_{i}$ sub-gamma, $\sup_{i}\mathbb{E}[\delta^{2}_{l,i}|\xi]<\infty$, implying $\mathbb{E}[\sum_{i}(\delta_{l,i}-\bar{\delta}_{l})^{2}\mid\xi]\leq 1/\min_{i}c_{l,i}\mathbb{E}[\sum_{i}c_{l,i}\delta^{2}_{l,i}\mid\xi]\leq(1/\min_{i}c_{l,i})\mathbb{E}[\delta^{2}_{l,i}|\xi]<\infty,$ as $\sum_{i}c_{l,i}=1$.
+By (D3), $n_{l}=|\{i\in S\cap C_{l}\}|=O(n_{\mathrm{eff},l})$, so $\mathbb{E}[\|\delta_{l}-\bar{\delta}_{l}\mathbf{1}\|_{2}^{2}\mid\xi]\leq n_{l}\cdot\mathbb{E}[\delta^{2}_{l,i}\mid\xi]\leq Cn_{\mathrm{eff},l}\cdot\frac{\nu}{M}.$ Therefore, $\|\delta_{l}-\bar{\delta}_{l}\mathbf{1}\|_{2}=O_{p}\left(\sqrt{\frac{n_{\mathrm{eff},l}}{M}}\right),$ and hence $\frac{1}{\sqrt{n_{\mathrm{eff},l}}}\|\delta_{l}-\bar{\delta}_{l}\mathbf{1}\|_{2}=O_{p}\left(M^{-1/2}\right).$ Put together, combining the bound from the first term involving $|\bar{\delta}_{l}|$ and the second term involving $\frac{1}{\sqrt{n_{\mathrm{eff},l}}}\|\delta_{l}-\bar{\delta}_{l}\mathbf{1}\|_{2}$:
+$|M^{*}_{n_{s}}-M_{n_{s}}|$ $\lesssim K\sum_{l\in\{L,R\}}O_{p}\left(M^{-1/2}n_{\text{eff},l}^{-1/2}\right)$
+$\quad+\sqrt{K}\sum_{l\in\{L,R\}}\left(1+O_{p}(M^{-1/2}n_{\text{eff},l}^{-1/2})\right)O_{p}(M^{-1/2})$
+$=O_{p}\left(M^{-1/2}\sum_{l\in\{L,R\}}n_{\text{eff},l}^{-1/2}\right).$
+From the sample score to the finite population score, by the usual Hajek linearization, for $l\in\{L,R\}$,
+$\|\mu_{\mathbf{k}}(P^{n_{s}}_{l})-\mu_{\mathbf{k}}(P_{l}^{N})\|=\Big\|\frac{1}{N_{l}}\sum_{i\in[N]:X_{i}\in C_{l}}\left(\frac{\xi_{i}}{\pi_{i}}-1\right)(\mathbf{k}(Y_{i},\cdot)-\mu_{\mathbf{k}}(P_{l}^{N}))+R_{l}\Big\|$
+$\leq$ $\frac{2\sqrt{K}}{N_{l}}\sum_{i\in[N]:X_{i}\in C_{l}}\left|\,\frac{\xi_{i}}{\pi_{i}}-1\,\right|+\|R_{l}\|=O_{p}(N^{-1/2}_{l})+o_{p}(1).$
+The last equation hold due to the assumption $\frac{1}{N}\sum_{i\in[N]}\mathbbm{1}(X_{i}\in C_{l})\rightarrow p\in(0,1]$, then the rate follows from the assumed CLT for $\sum_{i\in[N]}(\frac{\xi_{i}}{\pi_{i}}-1)$.
+Also,
+$\left|\frac{N^{n_{s}}_{L}N^{n_{s}}_{R}}{N^{n_{s}2}_{P_{a}}}-\frac{N_{L}N_{R}}{N^{2}_{P_{a}}}\right|=O_{p}(N_{L}^{-1/2}+N_{R}^{-1/2})=O_{p}(N_{P_{a}}^{-1/2}),\text{ when }N_{L}=\Theta(N_{R})=\Theta(N_{P_{a}}/2).$
+This is true if there exists $\alpha\in(0,1/2]$ with $N_{L}/N_{P_{a}},N_{R}/N_{P_{a}}\in[\alpha,1-\alpha]$.
+Follow the decomposition similar to S23 and a similar argument, $\left|M_{n_{s}}-M^{N}\right|=O_{p}(N^{-1/2}_{P_{a}}).$ Comparing the two orders, we can see when $M>\tilde{M}$, where $\tilde{M}$ satisfies $\frac{1}{\sqrt{\tilde{M}}}\sum_{l}n_{\text{eff},l}^{-1/2}\asymp N^{-1/2}_{P_{a}}$, equivalently $\tilde{M}\asymp N_{P_{a}}(\sum_{l}n_{\text{eff},l}^{-1/2})^{2}$, the rate from the design part $|M^{*}_{n_{s}}-M^{N}|$ dominates and further increasing $M$ cannot stabilize the split any further.
+∎
+S7 Consistency of MMD
+We now establish auxiliary results ensuring that all estimators in Theorem 3.4 are well-defined.
+Recall that $L(\mathbf{x})$ denotes the oracle leaf containing $\mathbf{x}$ when the splitting algorithm is applied to the full finite population $\{(X_{i},Y_{i})\}_{i=1}^{N}$.
+In this oracle setting, the design reduces to sampling with or without replacement from the finite population, and the splitting criterion simplifies to that of standard i.i.d.
+distributional random forests as all weights become $1/N$.
+In contrast, $L^{*}_{\xi}(\mathbf{x})$ denotes the estimated leaf when the algorithm is applied to the design-weighted sample with resampling.
+Both $L(\cdot)$ and $L^{*}_{\xi}(\cdot)$ involve algorithmic randomness (e.g., random selection of candidate split variables at each node); however, this randomness is independent of the sampling design and averages out as $B\to\infty$, following standard practice in random forests.
+We therefore suppress this randomness in our notation.
+The following lemmas, combining algorithmic assumptions (A1)–(A3) and (B1)–(B2) with design assumptions (D2)–(D4), establish that leaf sizes remain non-degenerate with high probability.
+Lemma S7.
+By (A2)-(A3), (B2)-(B3), for any given $\mathbf{x}$, $N_{L(\mathbf{x})}=\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))=\Theta_{p}(Nk^{p}_{N})$, and so $N_{L(\mathbf{x})}\to_{p}\infty$ as $N\to\infty$.
+Proof of Lemma S7.
+For any $M>0$,
+$P(N_{L(\mathbf{x})}\leq M)\leq P(N_{L(\mathbf{x})}\leq M,\;diam(L)<k_{N})+MP(diam(L(\mathbf{x}))\geq k_{N}).$
+The second part goes to zero by (B1).
+On $\{diam(L(\mathbf{x}))<k_{N}\}$ for which the event happens with probability 1o(1)1-o(1), by (A2), $\bar{B}_{r_{N}}(\mathbf{x})\subseteq L(\mathbf{x})\subseteq B_{k_{N}}(\mathbf{x}),$ so by (A3),
+$mc_{1}^{p}Nk_{N}^{p}\leq\#\{\,i:X_{i}\in B_{c_{1}k_{N}}(\mathbf{x})\}\;\leq\;\#\{\,i:X_{i}\in L(\mathbf{x})\}\;\leq\;\#\{\,i:X_{i}\in B_{k_{N}}(\mathbf{x})\}\leq MNk_{N}^{p},$
+so $N_{L(\mathbf{x})}=\Theta_{p}(Nk^{p}_{N})$. Finally, as, $\lim_{N\to\infty}Nk^{p}_{N}=\infty$, completing the proof. ∎
+Lemma S8.
+For any given $\mathbf{x}$, with $n_{s,N}=\sum_{i}\xi_{i}$, $n_{L(\mathbf{x})}=\sum_{i}\xi_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))$, it holds that $n_{s,N}=\Theta_{p_{d}}(N)$ and $n_{L(\mathbf{x})}/N=\Theta_{p}(k_{N}^{p})$.
+Proof of Lemma S8.
+(D2) implies for any $\epsilon>0$, $\exists\;N_{0}(\epsilon)\geq 1$ such that for all $N\geq N_{0}$, $P(\Big\{\frac{r}{2}\leq\frac{n_{s,N}}{N}\leq\frac{3r}{2}\Big\})\geq 1-\epsilon,$ so $n_{s,N}/N=O_{p_{d}}(1)$ and $N/n_{s,N}=O_{p_{d}}(1)$ by $r$ some finite constant bounded away from 0.
+For $n_{L(\mathbf{x})}$,
+$P\left(\left|\frac{n_{L(\mathbf{x})}}{\mathbb{E}[n_{L(\mathbf{x})}]}-1\right|\geq\epsilon\right)=P\left(\big|{n_{L(\mathbf{x})}}-\mathbb{E}[n_{L(\mathbf{x})}]\big|\geq\epsilon\mathbb{E}[n_{L(\mathbf{x})}]\right)\leq\frac{Var(n_{L(\mathbf{x})})}{\epsilon^{2}(\mathbb{E}[n_{L(\mathbf{x})}])^{2}}=\frac{1}{\epsilon^{2}\Theta_{p_{d}}(Nk_{N}^{p})}\to 0,$
+as $\pi_{i}\in(0,1)$ and result from Lemma (S7), it holds that
+$Var(n_{L(\mathbf{x})})$ $=\sum_{i}\pi_{i}(1-\pi_{i})I_{i}+\sum_{i\neq j}(\pi_{ij}-\pi_{i}\pi_{j})I_{i}I_{j}$
+$\leq N_{L(\mathbf{x})}+\frac{C_{1}}{2}\frac{N_{L(\mathbf{x})}}{n_{s,N}}(N_{L(\mathbf{x})}-1)\leq N_{L(\mathbf{x})}+\frac{C_{1}}{2}\frac{N}{n_{s,N}}(N_{L(\mathbf{x})}-1)\leq\Theta_{p}(Nk_{N}^{p})\Theta_{p}(1),\text{ and }$
+$\mathbb{E}[n_{L(\mathbf{x})}]$ $=\sum_{i}\pi_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\geq N_{L(\mathbf{x})}\underline{\lambda}\frac{n_{s,N}}{N}=\Theta_{p}(Nk_{N}^{p})\Theta_{p}(1).$
+So $n_{L(\mathbf{x})}=E[n_{L(\mathbf{x})}](1+o_{p}(1))=\Theta_{p}(Nk_{N}^{p})$, as desired. ∎
+Lemma S9.
+Assuming (D2), it holds that for $1\leq p\leq 4$
+$\sup_{N\geq 1}\mathbb{E}\left[\left(\frac{N}{n_{s,N}}\right)^{p}\mid n_{s,N}\geq 1\right]<\infty.$ (S24)
+Proof of Lemma S9.
+For any $\delta\in(0,r]$, $r>0$ Since $n_{s,N}/N\to_{p_{d}}f>0$, $P_{\mathcal{S}^{N}|\omega}(n_{s,N}\geq 1)\to 1$, so for $N$ large enough,
+$P_{\mathcal{S}^{N}|\omega}\!\left(\frac{n_{s,N}}{N}\leq\delta\,\Big|\,n_{s,N}\geq 1\right)\;\leq\;\frac{P_{\mathcal{S}^{N}|\omega}\!\left({n_{s,N}}/{N}\leq\delta\right)-P_{\mathcal{S}^{N}|\omega}(n_{s,N}=0)}{P_{\mathcal{S}^{N}|\omega}(n_{s,N}\geq 1)}\;\leq\;2C\,N^{-\kappa}.$
+So my splitting up into regions $\{n_{s,N}/N>\delta\}$ and $\{n_{s,N}/N\leq\delta\}$, it holds
+$\mathbb{E}\left[\left(\frac{N}{n_{s,N}}\right)^{p}\mid n_{s,N}\geq 1\right]\leq\delta^{-p}+2CN^{p-\kappa},\text{ with }p-\kappa<0,$
+completing the proof. ∎
+Lemma S10.
+Fix $\mathbf{x}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, $p\geq 1$.
+For $L(\mathbf{x})$ the oracle leaf, and $L^{*}_{\xi}(\mathbf{x})$ the leaf produced by the SDRF algorithm.
+Let $\{n^{*}_{b,i}\}_{i=1}^{N}$ be the unit-level resample multipliers used in the estimation stage for tree $b\in[B]$.
+Define indicators
+$I_{i}:=\mathbbm{1}\{X_{i}\in L(\mathbf{x})\},\quad I_{i}^{*}:=\mathbbm{1}\{X_{i}\in L^{*}_{\xi}(\mathbf{x})\},$
+and the (design-sample) leaf size $n_{L}(\mathbf{x}):=\sum_{i=1}^{N}\xi_{i}I_{i}.$ Assume (R1)–(R3), and (A1)–(A3), (B1)–(B2), Then the index set of “good trees”
+$\mathcal{B}_{L}(\mathbf{x}):=\left\{\,b\in[B]:\sum_{i=1}^{N}n^{*}_{b,i}\,I_{i}\in\big[\,n_{L}(\mathbf{x})-n_{d,N},\;n_{L}(\mathbf{x})+n_{d,N}\,\big]\right\},$
+$\mathcal{B}_{L^{*}}(\mathbf{x}):=\left\{\,b\in[B]:\sum_{i=1}^{N}n^{*}_{b,i}\,I_{i}^{*}\in\big[\,n_{L}(\mathbf{x})-n_{d,N},\;n_{L}(\mathbf{x})+n_{d,N}\,\big]\right\}$
+satisfy
+  1. 1.
+    $\lvert\mathcal{B}_{L}(\mathbf{x})\rvert/B\to_{p_{d}}1$ and $\lvert\mathcal{B}_{L^{*}}(\mathbf{x})\rvert/B\to_{p_{d}}1$.
+  2. 2.
+    For any tree index $b$, $P_{\mathcal{S}^{N}\mid\omega}\!\left(b\in\mathcal{B}_{L}(\mathbf{x})\,\Delta\,\mathcal{B}_{L^{*}}(\mathbf{x})\right)=o\!\left(\frac{1}{Nk_{N}^{p}}\right)+o\!\left(\frac{d_{N}}{k_{N}^{p}}\right),$ so in particular when $d_{N}=o(k_{N}^{p})$ and $Nk^{p}_{N}\to\infty$ then
+    $P_{\mathcal{S}^{N}\mid\omega}\!\left(b\in\mathcal{B}_{L}(\mathbf{x})\,\Delta\,\mathcal{B}_{L^{*}}(\mathbf{x})\right)=o(1).$
+Proof of Lemma S10.
+Part 1: We first control the mean and variance for the oracle leaf–based sizes. Define $S_{b}:=\sum_{i=1}^{N}n^{*}_{b,i}I_{i}\,\forall b.$ By (R2), the conditional mean
+$\mu$ $:=\mathbb{E}(S_{b}\mid\xi,L)=\sum_{i=1}^{N}\mathbb{E}(n^{*}_{b,i}\mid\xi)\,I_{i}=\sum_{i=1}^{N}\xi_{i}(1+o(1))I_{i}=n_{L}\{1+o(1)\}.$
+For its conditional variance,
+$Var(S_{b}\mid\xi,L)$ $=\sum_{i}I_{i}^{2}\,Var\!\bigl(n_{i}^{*}\mid\xi\bigr)\;+\;\sum_{i\neq j}I_{i}I_{j}\,Cov\!\bigl(n_{i}^{*},n_{j}^{*}\mid\xi\bigr)$
+$\;\leq\;\sum_{i=1}^{N}Var(n^{*}_{b,i}\mid\xi)\,I_{i}+\sum_{i\neq j}\bigl|Cov(n^{*}_{b,i},n^{*}_{b,j}\mid\xi)\bigr|\,\;\leq\;(1+C_{\mathrm{cov}})\sum_{i=1}^{N}Var(n^{*}_{b,i}\mid\xi)\,I_{i}$
+$\leq(1+C_{\mathrm{cov}})C_{2}\sum_{i=1}^{N}I_{i}=(1+C_{\mathrm{cov}})C_{2}\,N_{L},\quad N_{L}:=\sum_{i=1}^{N}I_{i}.$
+The first inequality by $\lvert I_{i}I_{j}\rvert\leq 1$, the second by (R3), and the third holds as by (R2), $\sup_{i}\mathbb{E}((n^{*}_{b,i})^{2}\mid\xi)\leq C_{2}$ a.s., and $\mathbb{E}(n^{*}_{b,i}\mid\xi)=\xi_{i}(1+o(1))$, so $Var(n^{*}_{b,i}\mid\xi)\leq C_{2}$ a.s.
+Together, $N_{L}=\Theta_{p}(Nk_{N}^{p})$ by Lemma S7, hence
+$Var(S_{b}\mid\xi,L)=O_{p_{d}}(Nk_{N}^{p}).$ (S25)
+By Lemma S7, $n_{L}(\mathbf{x})=\Theta_{p_{d}}(Nk^{p}_{N})$, thus one can always choose a sequence $n_{min,N}\to\infty$ with $n_{min,N}=o(Nk^{p}_{N})$ such that $P_{\mathcal{S}^{N}\mid\omega}(n_{L}(\mathbf{x})\geq 2n_{min,N})\to 1.$ Now choose $n_{d,N}\leq n_{\min,N}$ where $d$ denotes it is characterizing a tolerance region, so that on the high-probability event $\{n_{L}\geq 2n_{\min,N}\}$ we also have $n_{d,N}\leq n_{L}/2$.
+So $\{S_{b}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\}\subseteq\{|S_{b}-\mu|>\mu/2\},$ and Chebyshev gives
+$P\!\big(S_{b}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\mid\xi,L\big)\leq P\!\big(|S_{b}-\mu|>\mu/2\mid\xi,L\big)\leq\frac{4\,Var(S_{b}\mid\xi,L)}{\mu^{2}}.$
+Using $\mu=n_{L}=\Theta_{p_{d}}(Nk_{N}^{p})$ (Lemma S7) together with (S25), we obtain
+$P\!\left(S_{b}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\mid\xi,L\right)=O_{p_{d}}\!\left(\frac{1}{Nk_{N}^{p}}\right).$ (S26)
+Part2: We then control the mean and variance for the estimated leaf–based sizes.
+The argument is similar: For $S_{b}^{*}:=\sum_{i=1}^{N}n^{*}_{b,i}\,I_{i}^{*}$, by (R2) $\mu^{*}:=\mathbb{E}(S_{b}^{*}\mid\xi,L^{*})=\sum_{i=1}^{N}\mathbb{E}(n^{*}_{b,i}\mid\xi)\,I_{i}^{*}=\sum_{i=1}^{N}\\ \xi_{i}(1+o(1))I_{i}^{*}=n^{*}_{L}(\mathbf{x})\{1+o(1)\}.$ Note that $\mu^{*}-n_{L}=\sum_{i=1}^{N}\xi_{i}\,(I_{i}^{*}-I_{i}).$ By the same argument used in Lemma S10, $\mu^{*}-n_{L}=O_{p_{d}}(Nd_{N})$.
+Combined with $d_{N}=o(k_{N}^{p})$ and $n_{L}=\Theta_{p_{d}}(Nk_{N}^{p})$, it follows that $\mu^{*}=n_{L}\bigl(1+o_{p_{d}}(1)\bigr)=\Theta_{p_{d}}(Nk_{N}^{p}).$ For the conditional variance,
+$Var(S_{b}^{*}\mid\xi,L^{*})$ $\leq(1+C_{\mathrm{cov}})\sum_{i=1}^{N}Var(n^{*}_{b,i}\mid\xi)\,I_{i}^{*}\leq(1+C_{\mathrm{Cov}})C_{2}\sum_{i=1}^{N}I_{i}^{*}$
+$=(1+C_{\mathrm{cov}})C_{2}\,\sum_{i}I_{i}^{*}=O_{p_{d}}(Nk_{N}^{p}),$
+as $N_{L^{*}}=\Theta_{p}(Nk_{N}^{p})$ (Lemma S7) and $Nd_{N}=o(Nk_{N}^{p})$ (B2).
+Re–applying the previous argument on the same high-probability event $\{n_{L}\geq 2n_{\min,N}\}$, $\mu^{*}\geq n_{L}/2$ w.h.p., Therefore,
+$\{S_{b}^{*}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\}\subseteq\{|S_{b}^{*}-\mu^{*}|>\mu^{*}/2\}\ \cup\ \{|\mu^{*}-n_{L}|>n_{L}/2\}.$
+The second event is $o_{p_{d}}(1)$ by $|\mu^{*}-n_{L}|=O_{p_{d}}(Nd_{N})=o_{p_{d}}(Nk_{N}^{p})$ and $n_{L}\asymp Nk_{N}^{p}$.
+The first event is $O_{p_{d}}\!\left(\frac{1}{Nk_{N}^{p}}\right)$ from Chebyshev and $\mu^{*}=\Theta_{p_{d}}(Nk_{N}^{p})$ shown previously.
+Hence
+$P\!\left(S_{b}^{*}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\mid\xi,L^{*}\right)=O_{p_{d}}\!\left(\frac{1}{Nk_{N}^{p}}\right)+O_{p_{d}}\!\left(\frac{d_{N}}{k_{N}^{p}}\right).$ (S27)
+Part3: We now combine previous findings to finish the proof. Define the “bad” indicators
+$Z_{b}:=\mathbbm{1}\Big\{S_{b}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\Big\},\quad Z_{b}^{*}:=\mathbbm{1}\Big\{S_{b}^{*}\notin[n_{L}-n_{d,N},\,n_{L}+n_{d,N}]\Big\}.$
+Then $1-|\mathcal{B}_{L}|/B=B^{-1}\sum_{b=1}^{B}Z_{b}$.
+Fix any $\delta\in(0,1)$.
+By Markov and that (R1) implies the vectors $\{n_{b}^{*}\}_{b=1}^{B}$ are i.i.d.
+across $b$ under the resampling distribution $P(\,\cdot\mid\xi)$, so $\{\mathbbm{1}(b\notin\mathcal{B}_{L}(\mathbf{x}))\}_{b=1}^{B}$ conditional on $\xi$ is i.i.d.
+Bernoulli.
+[…]
+Therefore $|\mathcal{B}_{L}|/B\to_{p_{d}}1$ and a same argument with (S27) gives $|\mathcal{B}_{L^{*}}|/B\to_{p_{d}}1$, proving (a).
+Finally, for any fixed $b$, $\mathbbm{1}\{b\in\mathcal{B}_{L}\Delta\mathcal{B}_{L^{*}}\}\leq\mathbbm{1}\{b\notin\mathcal{B}_{L}\}+\mathbbm{1}\{b\notin\mathcal{B}_{L^{*}}\}.$ So taking expectation for each part and apply (S26) and (S27) again gives
+$P_{\mathcal{S}^{N}|\omega}\!\left(b\in\mathcal{B}_{L}\;\Delta\;\mathcal{B}_{L^{*}}\right)=O\!\left(\frac{1}{Nk_{N}^{p}}\right)+O\!\left(\frac{d_{N}}{k_{N}^{p}}\right),$
+proving (b). ∎
+Implications for non-degeneracy. The preceding lemmas ensure that all denominators in the SDRF estimator are bounded away from zero with probability approaching one. Specifically:
+  • By Lemma S7, the oracle leaf size satisfies $N_{L}(\mathbf{x})=\Theta_{p}(Nk_{N}^{p})\to_{p}\infty$, ensuring $|L(\mathbf{x})|\geq 1$ with high probability.
+  • By Lemma S8, the design-sampled leaf size satisfies $n_{L}(\mathbf{x})=\Theta_{p_{d}}(Nk_{N}^{p})$. Combined with the inclusion probability bounds in (D3), the design-weighted leaf size satisfies
+    $|L_{\xi}(\mathbf{x})|=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\geq\frac{n_{L}(\mathbf{x})}{\bar{\lambda}}\cdot\frac{N}{n_{s,N}}=\Theta_{p_{d}}(Nk_{N}^{p})\gg 1.$ (S28)
+  • By Lemma S10, for any tree $b\in[B]$, the bootstrap-weighted size $\sum_{i=1}^{N}n^{*}_{b,i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))$ concentrates around $n_{L}(\mathbf{x})$ and is thus bounded away from zero on a high-probability event, ensuring $|L^{*}_{\xi}(\mathbf{x})|\gg 1$.
+Thus all ratio estimators appearing in the proof of Theorem 3.4 are well-defined and bounded.
+The remainder of this proof is implicitly conditional on the intersection of these high-probability events, which occurs with probability approaching one.
+Proof of Theorem 3.4.
+We now prove the main theorem.
+We show design consistency of the MMD between the KME of the conditional distribution estimate and that of the truth.
+Namely, $\|\mu_{\mathbf{k}}(\hat{P}_{forest}(\mathbf{x}))-\mu_{\mathbf{k}}(P_{N}(\mathbf{x}))\|_{\mathcal{H}}=o_{p_{d}}(1)$, through the decomposition
+$\left\lVert\mu_{\mathbf{k}}\!\big(\hat{P}_{\mathrm{forest}}(\mathbf{x})\big)-\mu_{\mathbf{k}}\!\big(P_{N}(\mathbf{x})\big)\right\rVert_{\mathcal{H}}\leq A_{N}(\mathbf{x})+B_{N}(\mathbf{x})+C_{N}(\mathbf{x})+D_{N}(\mathbf{x})+E_{N}(\mathbf{x})$, where
+$A_{N}(\mathbf{x})$ $=\left\lVert\frac{1}{|\mathcal{B}_{L^{*}}(\mathbf{x})|}\!\sum_{b\in\mathcal{B}_{L^{*}}(\mathbf{x})}T^{*}_{N,b}(\mathbf{x})-\!\frac{1}{|\mathcal{B}_{L}(\mathbf{x})|}\!\sum_{b\in\mathcal{B}_{L}(\mathbf{x})}\sum_{i=1}^{N}\frac{n^{*}_{i}/\pi_{i}}{|L(\mathbf{x})|}\,\mathbbm{1}\{X_{i}\in L(\mathbf{x})\}\,\mathbf{k}(Y_{i},\cdot)\right\rVert_{\mathcal{H}},$
+$B_{N}(\mathbf{x})$ $=\left\lVert\frac{1}{|\mathcal{B}_{L}(\mathbf{x})|}\sum_{b\in\mathcal{B}_{L}(\mathbf{x})}\sum_{i=1}^{N}\frac{n^{*}_{i}/\pi_{i}}{|L(\mathbf{x})|}\,\mathbbm{1}\{X_{i}\in L(\mathbf{x})\}\,\mathbf{k}(Y_{i},\cdot)-\mathbb{E}_{n^{*}}\!\left[T^{*}_{N,b}(\mathbf{x})\,\mathbbm{1}\{b\in\mathcal{B}_{L}(\mathbf{x})\}\,\middle|\,\xi,L(\mathbf{x})\right]\right\rVert_{\mathcal{H}},$
+$C_{N}(\mathbf{x})$ $=\Bigg\|\mathbb{E}_{n^{*}}\!\left[T^{*}_{N,b}(\mathbf{x})\,\mathbbm{1}\{b\in\mathcal{B}_{L}(\mathbf{x})\}\,\middle|\,\xi,L(\mathbf{x})\right]-T_{\xi,L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}},$
+$D_{N}(\mathbf{x})$ $=\Bigg\|T_{\xi,L,b}(\mathbf{x})-T_{L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}},$
+$E_{N}(\mathbf{x})$ $=\Bigg\|T_{L,b}(\mathbf{x})-\mu_{N}(\mathbf{x})\Bigg\|_{\mathcal{H}}.$
+Where $\mathcal{B}(\mathbf{x})=\{b\in[B]:\;|\sum_{i=1}^{N}n^{*}_{b,i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))-n_{L(\mathbf{x})}|\leq n_{d}\},$ $\mathcal{B}(\mathbf{x})=\{b\in[B]:\;|\sum_{i=1}^{N}n^{*}_{b,i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))|\leq n_{d}\}$.
+$\hat{\mu}_{N}(\mathbf{x})=\frac{1}{B}\sum_{b=1}^{B}T^{*}_{N,b}(\mathbf{x}),$ with
+$T^{*}_{N,b}(\mathbf{x})$ $=\sum_{i\in D_{\text{est}}}\frac{n^{*}_{b,i}/\pi_{i}}{|L^{*}_{\xi,b}(\mathbf{x})|}\;\mathbbm{1}(X_{i}\in L^{*}_{\xi,b}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot),$
+$|L^{*(1)}_{\xi,b}(\mathbf{x})|$ $=\sum_{i\in D_{\text{est}}}{n^{*}_{b,i}/\pi_{i}}\;\mathbbm{1}(X_{i}\in L^{*}_{\xi,b}(\mathbf{x})),$
+$T_{\xi,L,b}(\mathbf{x})=\sum_{i=1}^{N}\frac{1}{|L_{\xi}(\mathbf{x})|}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L_{\xi}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot),\;\;|L_{\xi}(\mathbf{x})|=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x})),$
+$T_{L,b}(\mathbf{x})=\sum_{i=1}^{N}\frac{1}{|L(\mathbf{x})|}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot),\;\;|L(\mathbf{x})|=\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x})),$
+and the target quantity
+$\mu_{\mathbf{k}}(P_{N}(\mathbf{x}))=\frac{\sum_{i=1}^{N}\mathbbm{1}(X_{i}=\mathbf{x})\mathbf{k}(Y_{i},\cdot)}{\sum_{i=1}^{N}\mathbbm{1}(X_{i}=\mathbf{x})}.$
+After decomposition, $A_{N}$ concerns the effect of using resampled data for leaf construction rather than the full finite population.
+$B_{N}$ concerns the variance from resampling once the partition at the oracle and design $\xi$ are considered fixed.
+$C_{N}$ measures the mismatch introduced by using the resamples per tree instead of the original weights.
+$D_{N}$ isolates the impact of the sampling design.
+And $E_{N}$ concerns the approximation (bias) from using a leaf to estimate a point.
+Part A
+By (B2), for each unit $i\in[n]$, $\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))=\mathbbm{1}(X_{i}\in L(\mathbf{x}))+o_{p}(d_{N})$.
+Denote $\mathcal{B}^{\cap}(\mathbf{x})=\mathcal{B}_{L}(\mathbf{x})\cap\mathcal{B}_{L^{*}}(\mathbf{x})$.
+By Lemma S10, $P(b\in\mathcal{B}^{\cap}(\mathbf{x}))\to_{p_{d}}1$.
+Denote the set limit $\mathcal{B}(\mathbf{x})$.
+Lemma S10 gives $|\mathcal{B}(\mathbf{x})|=B(1+o_{p}(1))$.
+Consider the quantity
+$A_{N}=$ $\Bigg\|\frac{1}{|\mathcal{B}_{L^{*}}(\mathbf{x})|}\sum_{b\in\mathcal{B}_{L^{*}}(\mathbf{x})}\sum_{i=1}^{N}\frac{n_{b,i}^{*}/\pi_{i}}{|L^{*}_{\xi}(\mathbf{x})|}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)$
+- $\quad\frac{1}{|\mathcal{B}_{L}(\mathbf{x})|}\sum_{b\in\mathcal{B}_{L}(\mathbf{x})}\sum_{i=1}^{N}\frac{n_{b,i}^{*}/\pi_{i}}{|L(\mathbf{x})|}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)\Bigg\|_{\mathcal{H}}.$
+Denote $T^{*}_{L^{*},b}(\mathbf{x})=\sum_{i=1}^{N}\frac{n_{b,i}^{*}/\pi_{i}}{|L^{*}_{\xi}(\mathbf{x})|}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot),$ $T_{L,b}(\mathbf{x})=\sum_{i=1}^{N}\frac{n_{b,i}^{*}/\pi_{i}}{|L(\mathbf{x})|}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot),$ then
+$A_{N}=$ $\Bigg\|\frac{1}{|\mathcal{B}_{L^{*}}(\mathbf{x})|}\sum_{b\in\mathcal{B}_{L^{*}}(\mathbf{x})}T^{*}_{L^{*},b}(\mathbf{x})-\frac{1}{|\mathcal{B}_{L}(\mathbf{x})|}\sum_{b\in\mathcal{B}_{L}(\mathbf{x})}T_{L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}}$
+$\leq$ $\frac{1}{|\mathcal{B}^{\cap}|}\sum_{b\in\mathcal{B}^{\cap}}\Bigg\|\frac{|\mathcal{B}^{\cap}|}{|\mathcal{B}_{L^{*}}|}T^{*}_{L^{*},b}(\mathbf{x})-\frac{|\mathcal{B}^{\cap}|}{|\mathcal{B}_{L}|}T_{L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}}+\frac{1}{|\mathcal{B}_{L^{*}}|}\sum_{b\in\mathcal{B}_{L^{*}}\setminus\mathcal{B}^{\cap}}\Bigg\|T^{*}_{L^{*},b}(\mathbf{x})-\frac{|\mathcal{B}_{L^{*}}|}{|\mathcal{B}_{L}|}T_{L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}}$
+$+$ $\frac{1}{|\mathcal{B}_{L}|}\sum_{b\in\mathcal{B}_{L}\setminus\mathcal{B}^{\cap}}\Bigg\|\frac{|\mathcal{B}_{L}|}{|\mathcal{B}_{L^{*}}|}T^{*}_{L^{*},b}(\mathbf{x})-T_{L,b}(\mathbf{x})\Bigg\|_{\mathcal{H}}.$
+For the second part, its expectation conditioning on $\xi$ is upper bounded by
+$\frac{1}{|\mathcal{B}_{L^{*}}|}\sum_{b\in\mathcal{B}_{L}\cup\mathcal{B}_{L^{*}}}\left\|T^{*}_{L^{*},b}(\mathbf{x})-\frac{|\mathcal{B}_{L^{*}}|}{|\mathcal{B}_{L}|}T_{L,b}(\mathbf{x})\right\|_{L^{2}}P(b\in\mathcal{B}_{L^{*}}\setminus\mathcal{B}^{\cap})$
+$\leq$ $\Big(\frac{1}{B}+o_{p}(1)\Big)\sum_{b\in\mathcal{B}}\Bigg\{\|T^{*}_{L^{*},b}(\mathbf{x})\|_{L^{2}}+(1+o_{p}(1))\|T_{L,b}(\mathbf{x})\|_{L^{2}}\Bigg\}O_{p}(\frac{1}{Nk^{p}_{N}})=O_{p}(\frac{1}{Nk^{p}_{N}}),$
+where the last equality holds as by (D2)-(D3), and intermediate results proved that $\sum_{i}n^{*}_{i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))=O_{p}(Nk^{p}_{N})$ during Lemma S10, $\|T^{*}_{L^{*},b}(\mathbf{x})\|_{\mathcal{H}}=\|\frac{\sum_{i}n^{*}_{i}/\pi_{i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)}{\sum_{i}n^{*}_{i}/\pi_{i}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))}\|_{\mathcal{H}}=O_{p}(1),$ and similarly $\|T_{L}(\mathbf{x})\|_{\mathcal{H}}=O_{p}(1)$.
+By a similar argument, the third term in the upper bound of $A_{N}$ is also $o_{p}(1)$.
+Thus
+$A_{N}\;\leq\;$ $\frac{1}{\mathcal{B}}\sum_{b\in\mathcal{B}}\Big\|\sum_{i=1}^{N}\frac{n^{*}_{i}}{\pi_{i}}\Big\{\frac{\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))}{|L^{*}_{\xi}(\mathbf{x})|}-\frac{\mathbbm{1}(X_{i}\in L(\mathbf{x}))}{|L(\mathbf{x})|}\Big\}\mathbf{k}(Y_{i},\cdot)\Big\|_{\mathcal{H}}+o_{p}(1)$
+$\;\leq\;$ $\frac{K}{\min\{|L^{*}_{\xi}(\mathbf{x})|,|L(\mathbf{x})|\}}\frac{N}{\underline{\lambda}n_{s,N}}\sum_{i=1}^{N}n^{*}_{i}\Big|\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))-\mathbbm{1}(X_{i}\in L(\mathbf{x}))\Big|+o_{p}(1).$
+We now bound the sum $\sum_{i=1}^{N}n^{*}_{i}|\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))-\mathbbm{1}(X_{i}\in L(\mathbf{x}))|$.
+Define the symmetric difference set $S_{\triangle}:=\{i\in[N]:X_{i}\in L^{*}_{\xi}(\mathbf{x})\triangle L(\mathbf{x})\}.$ So $\sum_{i=1}^{N}n^{*}_{i}|\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x}))-\mathbbm{1}(X_{i}\in L(\mathbf{x}))|=\sum_{i\in S_{\triangle}}n^{*}_{i}.$ The cardinality of $S_{\triangle}$ satisfies $|S_{\triangle}|=\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x})\triangle L(\mathbf{x})).$ By the strong law of large numbers applied to the finite population,
+$\frac{1}{N}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L^{*}_{\xi}(\mathbf{x})\triangle L(\mathbf{x}))\xrightarrow{p}P(X\in L^{*}_{\xi}(\mathbf{x})\triangle L(\mathbf{x})).$
+By assumption (B2), $\sup_{\mathbf{x}}P(X\in L^{*}(\mathbf{x})\triangle L(\mathbf{x}))=o_{p}(d_{N})$. Therefore,
+$|S_{\triangle}|=N\cdot o_{p}(d_{N})=o_{p}(Nd_{N}).$
+For the weighted sum, by (R2), $\mathbb{E}[n^{*}_{i}\mid\xi]=1+o(1)$ uniformly in $i$, and $\sup_{i}\mathbb{E}[(n^{*}_{i})^{2}\mid\xi]\leq C$ by bounded second moments.
+Conditional on $|S_{\triangle}|\geq 1$ which occurs with high probability for large $N$,
+$\sum_{i\in S_{\triangle}}n^{*}_{i}$ $=|S_{\triangle}|\cdot\frac{1}{|S_{\triangle}|}\sum_{i\in S_{\triangle}}n^{*}_{i}=|S_{\triangle}|\cdot(1+o_{p}(1))=o_{p}(Nd_{N})\cdot(1+o_{p}(1))=o_{p}(Nd_{N}),$ (S29)
+where the second equality holds by LLN for multipliers.
+Substituting (S29) back into the bound for $A_{N}$, together with By $\min\{|L^{*}_{\xi}(\mathbf{x})|,|L(\mathbf{x})|\}=\Theta_{p}(Nk_{N}^{p})$ from Lemma S7 and (B2) and $n_{s,N}=\Theta_{p_{d}}(N)$ from (D2):
+$A_{N}$ $\leq\frac{K}{\min\{|L^{*}_{\xi}(\mathbf{x})|,|L(\mathbf{x})|\}}\frac{N}{\underline{\lambda}n_{s,N}}\cdot o_{p}(Nd_{N})+o_{p}(1)$
+$\leq\frac{K\cdot N\cdot o_{p}(Nd_{N})}{\Theta_{p}(Nk_{N}^{p})\cdot\Theta_{p_{d}}(N)}+o_{p}(1)=\frac{o_{p}(Nd_{N})}{\Theta_{p}(k_{N}^{p})}+o_{p}(1)=o_{p}\left(\frac{d_{N}}{k_{N}^{p}}\right)+o_{p}(1).$
+Since (B2) requires $d_{N}=o(k_{N}^{p})$, we have $o_{p}(d_{N}/k_{N}^{p})=o_{p}(1)$, completing Part A.
+Part B
+Conditioning on the oracle leaf $L$ in the finite population (which is irrespective of design) as well as the inclusion indicators, across $b$ the resamples are independent and identically distributed.
+For any $\epsilon>0$,
+$\begin{split}&P_{n^{*},\xi,L(\mathbf{x})}\Big(\Bigg\|\frac{1}{\mathcal{B}_{L}(\mathbf{x})}\sum_{b\in\mathcal{B}_{L}(\mathbf{x})}T^{*}_{N,b}(\mathbf{x})-\mathbb{E}_{n^{*}}[T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\;|\;\xi,L(\mathbf{x})]\Bigg\|_{\mathcal{H}}\geq\epsilon\Big)\\ =&\mathbb{E}_{\xi,L(\mathbf{x})}\Big[P_{n^{*}|\xi,L(\mathbf{x})}\Big(\Big\|\frac{1}{B}\sum_{b=1}^{B}\frac{B}{|\mathcal{B}_{L}(\mathbf{x})|}\Bigg\{T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))-\mu\Bigg\}\Big\|_{\mathcal{H}}\geq\epsilon\Big)\Big]\text{ where }\end{split}$ (S30)
+$\mu$ $=\mathbb{E}_{n^{*}}\Big[T^{*}_{N,b}(\mathbf{x})\;|\;\xi,L(\mathbf{x})\Big]P(b\in\mathcal{B}_{L}(\mathbf{x})\mid\xi,L(\mathbf{x})).$
+By that for any fixed $\delta\in(0,1/2)$,
+$I\left(\frac{B}{|\mathcal{B}_{L}(\mathbf{x})|}\|\frac{1}{B}\sum_{b}\Big\{T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))-\mu\Big\}\|_{\mathcal{H}}\geq\epsilon\right)$
+$\leq\;$ $I\left(\frac{|\mathcal{B}_{L}(\mathbf{x})|}{B}\leq 1-\delta\right)+I\left(\frac{1}{B}\sum_{b}\Big\{T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))-\mu\Big\}\|_{\mathcal{H}}\geq\epsilon(1-\delta)\right),$
+[…]
+$\begin{split}&\mathbb{E}\Big[Var(T_{b}^{*})\mathbbm{1}(b\in\mathcal{B}(\mathbf{x}))|\xi,\;L(\mathbf{x}))\Big]\\ =&\mathbb{E}\Big[Var_{n^{*}}\Big(\frac{N_{\xi,L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}/\pi_{i}}{N_{\xi,L,x}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\;\Big|\;\xi,L(\mathbf{x})\Big)\Big].\end{split}$ (S31)
+Note that $b\in\mathcal{B}_{L}(\mathbf{x})$ if $|\sum_{i=1}^{N}n^{*}_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))-n_{L(\mathbf{x})}|\leq n_{d}\overset{\text{denote}}{=}\mathcal{E}_{b}$.
+As conditioning on $\xi$, the indicators $\mathbbm{1}(\mathcal{E})$ are i.i.d.
+Bernoulli, $\{|\mathcal{B}_{L}(\mathbf{x})|/B\leq 1-\delta\}=\frac{1}{B}\sum_{b=1}^{B}\mathbbm{1}(\mathcal{E}_{b}^{c})\geq\delta\}$ so by Markov and Lemma S10,
+$\mathbb{E}_{\xi,L(\mathbf{x})}\left[P_{n^{*}}\big(\frac{|\mathcal{B}_{L}(\mathbf{x})|}{B}\leq 1-\delta\mid\xi,L(\mathbf{x})\big)\right]\leq\frac{1}{\delta}\mathbb{E}[\mathbbm{1}(\mathcal{E}_{b}^{c})]=\frac{1}{\delta}O_{p}\left(\frac{1}{Nk^{p}_{N}}\right)=o_{p}(1).$
+[…]
+$\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}=1+o_{P(n^{*},\xi|L(\mathbf{x}))}(1)\text{,\; as $\forall\epsilon>0$, }$ (S32)
+$P(|\frac{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}{{N}_{L(\mathbf{x})}}-1|\geq\epsilon\mid L(\mathbf{x}))=P(|{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-{N}_{L(\mathbf{x})}|\geq\epsilon{N}_{L(\mathbf{x})}\mid L(\mathbf{x}))$
+$\leq$ $\frac{Var(\widehat{N}^{*}_{\xi,L(\mathbf{x})}\mid L(\mathbf{x}))}{\epsilon^{2}N_{L(\mathbf{x})}^{2}}=O_{p}(N^{-1}_{L(\mathbf{x})}),\text{ as }$
+$\begin{split}&Var(\sum_{i=1}^{N}\frac{n^{*}_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mid L(\mathbf{x}))\\ =&\mathbb{E}_{\xi}\Big[Var\big(\sum_{i=1}^{N}\frac{n^{*}_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mid\xi,\;L(\mathbf{x})\big)\mid L(\mathbf{x})\Big]+Var_{\xi}\Big(\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mid L(\mathbf{x})\Big)\\ =&\mathbb{E}[\sum_{i=1}^{N}Var(n^{*}_{i}\mid\xi,\;L(\mathbf{x}))\frac{I_{i}}{\pi^{2}_{i}}]+\mathbb{E}[\sum_{i\neq j}Cov(n^{*}_{i},n^{*}_{j}\mid\xi,\;L(\mathbf{x}))\frac{I_{i}I_{j}}{\pi_{i}\pi_{j}}]+\sum_{i}\frac{1-\pi_{i}}{\pi_{i}}I_{i}+\sum_{i}\frac{\pi_{ij}-\pi_{i}\pi_{j}}{\pi_{i}\pi_{j}}I_{i}I_{j}\\ \leq&\mathbb{E}[\sum_{i=1}^{N}\Big\{\xi_{i}(1-\pi_{i})+\frac{\xi_{i}}{\pi_{i}}\big(1-\frac{\xi_{i}/\pi_{i}}{\hat{N}}\big)\Big\}\frac{I_{i}}{\pi^{2}_{i}}]+\sum_{i}\underline{\lambda}\mathbb{E}_{\xi}[\frac{N}{n_{s,N}}]I_{i}+\sum_{i\neq j}\mathbb{E}_{\xi}[\frac{cN^{2}}{\underline{\lambda}^{2}n^{3}_{s,N}}]I_{i}I_{j}\\ \leq&2\underline{\lambda}^{-2}\mathbb{E}\big[\big(\frac{N}{n_{s,N}}\big)^{2}\big]N_{L(\mathbf{x})}+\underline{\lambda}\mathbb{E}\big[\frac{N}{n_{s,N}}\big]N_{L(\mathbf{x})}+\mathbb{E}\big[\frac{cN^{2}}{\underline{\lambda}^{2}n^{3}_{s,N}}\big]N_{L(\mathbf{x})}(N_{L(\mathbf{x})}-1).\end{split}$ (S33)
+The first inequality by oracle partition independent of $\xi$ and $Cov(n^{*}_{i},n^{*}_{j}|\xi)\leq 0$; the second by $1-(\xi_{i}/\pi_{i})/\hat{N}\in[0,1]$, $\pi_{i}\in[\;\underline{\lambda}n_{s,N}/N,\;1)$, and then (D2)-(F5), implying
+$Var(\widehat{N}^{*}_{\xi,L(\mathbf{x})}\mid L(\mathbf{x}))/N^{2}_{L(\mathbf{x})}=O_{p}(1/N_{L(\mathbf{x})})+O_{p}(1/N(1+1/N_{L(\mathbf{x})}))=O_{p}(1/N_{L(\mathbf{x})})$
+[…]
+Additionally,
+$\sup_{N\geq 1}\mathbb{E}_{n^{*},\xi}\left[\Big|\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1\Big|^{p}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid L(\mathbf{x})\right]<\infty,\text{ for any $p\geq 1$},$ (S34)
+as $\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1=\frac{-W_{N}}{1+W_{N}}$ with $W_{N}=\frac{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}{N_{L(\mathbf{x})}}-1$, $|\frac{-W_{N}}{1+W_{N}}|^{p}$ achieves the largest value when $W_{N}$ reaches its smallest negative value.
+on $b\in\mathcal{B}_{L}(\mathbf{x})$, $W_{N}\geq\frac{\pi_{max}}{\pi_{min}}(1-n_{d}/n_{L(\mathbf{x})})-1\geq\frac{\bar{\lambda}}{\underline{\lambda}}(1-n_{d}/n_{min})-1>-1$ for all $N$.
+Write variance of product of two terms in terms of variance and moments of each gives
+$\mathbb{E}_{\xi,L(\mathbf{x})}\Big[Var\Big(T_{b}^{*}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid\xi,L(\mathbf{x})\Big)\Big]$
+$=$ $\mathbb{E}\Big[\mathbb{E}\Big[Var(\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x}))\mid L(\mathbf{x})\Big]\Big]$
+$+$ $\mathbb{E}\Big[\mathbb{E}\big[\Big\{(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}})^{2}-1\Big\}\Big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big\|^{2}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid\xi,L(\mathbf{x})\Big]\Big]$
+- $2\mathbb{E}\Big[\mathbb{E}\big[\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}|\;\xi,L(\mathbf{x})\Big]\mathbb{E}\Big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})\Big]\Big]$
+- $\mathbb{E}\Big[\Big(\mathbb{E}\big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})\Big]\Big)^{2}\Big].$
+We show two intermediate results that we will use latter: Firstly,
+$\sup_{N\geq 1}\mathbb{E}\left[\Big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big\|^{4}_{\mathcal{H}}\;|\;L(\mathbf{x})\right]<\infty,\text{ as }$ (S35)
+$\begin{split}\Big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big\|_{L^{4}(P_{n^{*},\xi\mid L(\mathbf{x})})}\leq&\frac{1}{N_{L(\mathbf{x})}}\sum_{i}I_{i}\Big\{\mathbb{E}_{\xi}[\mathbb{E}_{n^{*}}\big[\|\frac{{n^{*}_{i}}k_{i}}{\pi_{i}}\|^{4}_{\mathcal{H}}\mid\xi,L(\mathbf{x})\big]\mid L(\mathbf{x})]\Big\}^{1/4}\\ \leq&K\frac{1}{N_{L(\mathbf{x})}}\sum_{i}I_{i}\Big\{\mathbb{E}_{\xi}[\mathbb{E}\big[\|\frac{n^{*}_{i}}{\pi_{i}}\|^{4}_{\mathcal{H}}\mid\xi\big]]\Big\}^{1/4}\\ \leq&K15^{1/4}\underline{\lambda}^{3/4}\mathbb{E}_{\xi}[(\frac{N}{n_{s,N}})^{3/4}]\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))/N_{L(\mathbf{x})}\end{split}$ (S36)
+w.r.t. $P(n^{*},\xi\mid L(\mathbf{x}))$, where the second inequality uses independence of $L(\mathbf{x})$ and $\xi$; the third inequality uses
+$\begin{split}&\mathbb{E}[(n_{i}^{*})^{4}\mid\xi]=\mathbb{E}_{N^{*}}[\mathbb{E}[(n_{i}^{*})^{4}\mid\xi,\;N^{*}]\mid\xi]\\ =&\pi_{i}^{4}\mathbb{E}_{N^{*}}[N^{*}_{i}(N^{*}_{i}-1)(N^{*}_{i}-2)(N^{*}_{i}-3)\mid\xi]+\,6\pi_{i}^{3}(1-\pi_{i})\,\mathbb{E}\big[N_{i}^{*}(N_{i}^{*}-1)(N_{i}^{*}-2)\,\big|\,\xi\big]\\ &+\,7\pi_{i}^{2}(1-\pi_{i})^{2}\,\mathbb{E}\big[N_{i}^{*}(N_{i}^{*}-1)\mid\xi\big]+\,\pi_{i}(1-\pi_{i})^{3}\,\mathbb{E}\big[N_{i}^{*}\,\big|\,\xi\big]\\ =&\xi_{i}\Big\{(1-\pi_{i})^{3}+7\frac{\widehat{N}-1}{\widehat{N}}\,(1-\pi_{i})^{2}\\ &+6\frac{(\widehat{N}-1)(\widehat{N}-2)}{\widehat{N}^{2}}\,(1-\pi_{i})+\frac{(\widehat{N}-1)(\widehat{N}-2)(\widehat{N}-3)}{\widehat{N}^{3}}\,\Big\}\leq 15\xi_{i},\end{split}$ (S37)
+[…]
+Secondly,
+$\sup_{N\geq 1}\mathbb{E}\left[\Big(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}+1\Big)^{4}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid L(\mathbf{x})\right]<\infty,\text{ as }$ (S38)
+[…]
+$\begin{split}&\sup_{N\geq 1}\mathbb{E}\left[\big(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}+1\big)^{4}\mid L(\mathbf{x})\right]\lesssim\sup_{N\geq 1}\mathbb{E}\Big[(\frac{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}{N_{L(\mathbf{x})}}-1)^{4}\mid L(\mathbf{x})\Big]\\ \leq&\frac{1}{N^{4}_{L(\mathbf{x})}}\mathbb{E}\Big[\Big(\sum_{i}(\frac{n^{*}_{i}}{\pi_{i}}-1)I_{i}\mid\Big)^{4}\mid L(\mathbf{x})\Big]\leq\frac{1}{N^{4}_{L(\mathbf{x})}}\mathbb{E}\Big[\Big(N_{L(\mathbf{x})}^{3/4}(\sum_{i}|\frac{n^{*}_{i}}{\pi_{i}}-1|^{4}I_{i})^{1/4}\Big)^{4}\mid L(\mathbf{x})\Big]\\ \leq&\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbb{E}\Big[|\frac{n^{*}_{i}}{\pi_{i}}-1|^{4}\mid L(\mathbf{x})\Big]I_{i}\\ \leq&\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}2^{3}\Big\{\mathbb{E}_{\xi}[\mathbb{E}\Big[(\frac{n^{*}_{i}}{\pi_{i}})^{4}\mid\xi\Big]]+1\Big\}I_{i}\\ \leq&\mathbb{E}_{\xi}\left\{(\frac{2N}{\underline{\lambda}n_{s,N}})^{3}(15+\frac{N}{\underline{\lambda}n_{s,N}})\right\}\sum_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))/N_{L(\mathbf{x})}\end{split}$ (S39)
+w.r.t.
+$P(n^{*},\xi|L(\mathbf{x}))$.
+The third inequality by Holder; the fifth inequality by $(a+b)^{4}\leq 2^{3}(a^{4}+b^{4})$; and the last inequality by the previously shown result $\mathbb{E}[(n^{*}_{i})^{4}\mid\xi]\leq 15\;\xi_{i}$; and finally by Lemma S9.
+[…]
+$\mathbb{E}\Big[\mathbb{E}[\Big\{(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}})^{2}-1\Big\}\Big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big\|^{2}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid\xi,L(\mathbf{x})]\Big]$
+$=$ $\mathbb{E}_{L(\mathbf{x})}\Big[\mathbb{E}_{\xi}\Big[\mathbb{E}_{n^{*}}[\Big\{(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}})^{2}-1\Big\}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\Big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big\|^{2}\mid\xi,L(\mathbf{x})]\mid L(\mathbf{x})\Big]\Big]$
+$\leq$ $\mathbb{E}_{L(\mathbf{x})}\Bigg[\Big(\mathbb{E}_{n^{*},\xi}\Big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1)^{4}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid L(\mathbf{x})\Big]\Big)^{1/4}\Big(\mathbb{E}_{n^{*},\xi}\Big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}+1)^{4}\mid L(\mathbf{x})\Big]\Big)^{1/4}$
+$\quad\quad\times\sqrt{\mathbb{E}\Big[\big\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\big\|^{4}\;|\;L(\mathbf{x})\Big]}\;\;\Bigg]=\mathbb{E}_{L(\mathbf{x})}[o_{P(n^{*},\xi|L(\mathbf{x}))}(1)]=o_{p}(1),$
+Also, dominated convergence again for $N_{L(\mathbf{x})}/\widehat{N}^{*}_{L(\mathbf{x})}-1$ plus Cauchy Schwarz and $\|\cdot\|_{L^{p}}$ norm monotonicity implies
+$\mathbb{E}_{\xi,L(\mathbf{x})}\Big[\mathbb{E}_{n^{*}}\Big[\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})\Big]\mathbb{E}_{n^{*}}\Big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,\mathbf{x}}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})\Big]\Big]$
+$=$ $\mathbb{E}_{L(\mathbf{x})}\Big[\mathbb{E}\Big[\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid L(\mathbf{x})\Big]\mathbb{E}\Big[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid L(\mathbf{x})\Big]\Big]$
+$\leq$ $\mathbb{E}_{L(\mathbf{x})}\Big[K\frac{\sum_{i}I_{i}}{N_{L(\mathbf{x})}}\sqrt{\mathbb{E}[\big(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{L(\mathbf{x})}}-1\big)^{2}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid L(\mathbf{x})]}\sqrt{\mathbb{E}[\big(\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\big)^{2}\mid L(\mathbf{x})]}\;\;\Big]=o_{p}(1).$
+Further, by Jensen inequality and a similar argument as above,
+$\mathbb{E}\Big[\Big(\mathbb{E}[(\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{\xi,L(\mathbf{x})}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})]\Big)^{2}\Big]$
+$\leq$ $\mathbb{E}\Big[\mathbb{E}[\Big((\frac{N_{L(\mathbf{x})}}{\widehat{N}^{*}_{L(\mathbf{x})}}-1)\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\Big)^{2}\mid L(\mathbf{x})]\Big]\to 0.$
+Finally, using $\sup_{y_{i}\in\mathcal{Y}}\|k(y_{i},\cdot)\|_{\mathcal{H}}\leq K$, by similar reasoning used when showing $\widehat{N}^{*}_{L(\mathbf{x})}/N_{L(\mathbf{x})}\to_{p}1$,
+$\mathbb{E}\Big[\mathbb{E}[Var\big(\frac{1}{N_{L(\mathbf{x})}}\sum_{i}\frac{n^{*}_{i}}{\pi_{i}}I_{i}k_{i}\mid\xi,L(\mathbf{x})\big)\mid L(\mathbf{x})]\Big]$
+$\leq$ $\sup_{y}\|\mathbf{k}(Y_{i},\cdot)\|^{2}_{\mathcal{H}}\frac{1}{N^{2}_{L(\mathbf{x})}}\Big\{\mathbb{E}\big[\mathbb{E}[\sum_{i=1}^{N}Var(n^{*}_{i}\mid\xi,L(\mathbf{x}))\frac{I_{i}}{\pi^{2}_{i}}]+\mathbb{E}[\sum_{i\neq j}Cov(n^{*}_{i},n^{*}_{j}\mid\xi,L(\mathbf{x}))\frac{I_{i}I_{j}}{\pi_{i}\pi_{j}}]\big]\Big\}$
+$\leq$ $K^{2}\frac{1}{N^{2}_{L(\mathbf{x})}}\underline{\lambda}^{-2}\mathbb{E}[\big(\frac{N}{n_{s,N}}\big)^{2}]N_{L(\mathbf{x})}=o_{p}(1).$
+Together, we have shown
+$\frac{1}{\epsilon^{2}(1-\delta)^{2}B}\mathbb{E}_{\xi,L(\mathbf{x})}\Big[Var\Big(T_{b}^{*}\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid\xi,L(\mathbf{x})\Big)\Big]=\frac{O_{p}(1/(Nk_{N}^{p}))}{\epsilon^{2}(1-\delta)^{2}B}=o_{p}(1),$
+completing this section of the proof.
+Remark.
+Note that convergence happens without needing B to grow with $N$.
+From another perspective though, it also means that for $k^{p}_{N}$ not shrinking fast enough, this part of the MMD can still be controlled by letting $B$ compensate for the rate.
+Part C
+We now aim to find the rate for
+$\begin{split}&\mathbb{E}_{\xi,L(\mathbf{x})}\Bigg[\Big\|\mathbb{E}_{n^{*}}\left[T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\mid\xi,L(\mathbf{x})]-T_{\xi,L,b}(\mathbf{x})\right]\Big\|_{\mathcal{H}}\Bigg]\\ =&\mathbb{E}\Bigg[\Big\|\mathbb{E}_{n^{*}}[T^{*}_{N,b}(\mathbf{x})\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))-T_{\xi,L,b}(\mathbf{x})\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\Bigg]\end{split}$ (S40)
+We omit the subscript $N$ and $b$, and the evaluation point $\mathbf{x}$ for simplicity and without introducing confusion. Consider linearization of
+$T_{b}^{*}(\mathbf{x};D)=\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))\sum_{i=1}^{N}{n}_{i}^{*}\frac{\xi_{i}/\pi_{i}}{|L^{*}(\mathbf{x})|}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)\overset{\text{denote}}{=}\frac{T^{*}_{nu}}{T^{*}_{de}},\;\;|L^{*}(\mathbf{x})|=\sum_{i=1}^{N}\frac{{n}_{i}^{*}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))$
+centered at
+$T_{\xi,L}(\mathbf{x})=\sum_{i=1}^{N}\frac{\xi_{i}/\pi_{i}}{|L_{\xi}(\mathbf{x})|}\mathbbm{1}(X_{i}\in L_{\xi}(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)\overset{\text{denote}}{=}\frac{T_{L,nu}}{T_{L,de}},\;\;|L_{\xi}(\mathbf{x})|=\sum_{i=1}^{N}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x})).$
+For notation simplicity, we use subscript $i$ or $b$ to differentiate between $I_{b}=\mathbbm{1}(b\in\mathcal{B}_{L}(\mathbf{x}))$ and $I_{i}=\mathbbm{1}(X_{i}\in L(\mathbf{x}))$.
+For $g(t_{nu},t_{de})=t_{nu}/t_{de}$, with $\nabla g(t_{nu},t_{de})=\begin{pmatrix}\dfrac{1}{t_{de}},-\dfrac{t_{nu}}{t_{de}^{2}}\end{pmatrix}^{\!\top}$, and Hessian of $g$
+$H_{g}(t_{nu},t_{de})=\begin{pmatrix}0&-{1}/{t_{\mathrm{de}}^{2}}\\[6.0pt] -{1}/{t_{\mathrm{de}}^{2}}&{2\,t_{\mathrm{nu}}}/{t_{\mathrm{de}}^{3}}\end{pmatrix}$.
+Denote $\delta_{i}=\frac{{n}^{*}_{i}}{\pi_{i}}-\frac{\xi_{i}}{\pi_{i}}$, satisfying $\mathbb{E}[\delta_{i}\mid\xi,L(\mathbf{x})]=0$.
+(Note that ${n}^{*}_{i}={n}^{*}_{i}\xi_{i}$ a.s.
+$\mathbb{E}[{n}^{*}_{i}]=\mathbb{E}[\mathbb{E}[{n}^{*}_{i}|\xi]]=\pi_{i}$.) Accordingly,
+$\Delta_{nu}=T^{*}_{nu}-T_{L,nu}=I_{b}\sum_{i}\delta_{i}I_{i}k_{i}+(1-I_{b})T_{L,nu}\in\mathcal{H}\text{, and}$
+$\Delta_{de}=T^{*}_{de}-T_{L,de}=\sum_{i}\delta_{i}I_{i}\in\mathbb{R}$. and we have
+$T^{*}-T_{L}=$ $\frac{1}{T^{*}_{de}}\bigl(T^{*}_{nu}-T_{L,nu}\bigr)\;-\;\frac{T_{L,nu}}{T_{L,de}^{2}}\bigl(T^{*}_{de}-T_{L,de}\bigr)\;+\;R_{2}$
+$=$ $I_{b}\sum_{i=1}^{N}(\frac{{n}^{*}_{i}}{\pi_{i}}-\frac{\xi_{i}}{\pi_{i}})\,\underbrace{\left\{\frac{\mathbbm{1}(X_{i}\in L(\mathbf{x}))}{\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))}(k_{i}-T_{L}(\mathbf{x}))\right\}}_{u_{i}}\;+\;(I_{b}-1)T_{L}(\mathbf{x})\;+\;R_{2}(\mathbf{x}).$
+Denote $T^{(1)}=\sum_{i=1}^{N}(\frac{{n}^{*}_{i}}{\pi_{i}}-\frac{\xi_{i}}{\pi_{i}})u_{i}$.
+Now $u_{i}$ are fixed functions in the finite population; while the multiplier ${n}^{*}_{i}/\pi_{i}-\xi_{i}/\pi_{i}$ contains all randomness from the design and the bootstrapping of $D_{\text{est}}$.
+By unbiasedness of the multiplier $n^{*}_{i}$ for $\xi_{i}$, $\mathbb{E}_{n^{*}}[I_{b}T^{(1)}(\mathbf{x})\mid\xi,L(\mathbf{x})]=0$, and
+[…]
+$\leq\;$ $\mathbb{E}_{\xi,L(\mathbf{x})}\Big[\Big\|T_{\xi,L}(\mathbf{x})\Big\|_{\mathcal{H}}\Big\|\;\mathbb{E}[b\not\in\mathcal{B}_{L}(\mathbf{x})\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\Big]\;+\;\mathbb{E}_{\xi,L(\mathbf{x})}\Big[\Big\|\mathbb{E}[R_{2}(\mathbf{x})\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\Big].$
+By regularity conditions (D3) and boundedness of kernel, $\|T_{\xi,L}(\mathbf{x})\|_{\mathcal{H}}\leq\frac{\bar{\lambda}}{\underline{\lambda}}K<\infty,$ so Lemma S10 gives
+$\mathbb{E}_{\xi,L(\mathbf{x})}\Big[\Big\|T_{\xi,L}(\mathbf{x})\Big\|_{\mathcal{H}}\Big\|\;\mathbb{E}[b\not\in\mathcal{B}_{L}(\mathbf{x})\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\Big]$
+$\leq$ $\frac{\bar{\lambda}}{\underline{\lambda}}K\mathbb{E}\left[\|P(|\sum_{i}n^{*}_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))-n_{L(\mathbf{x})}|\geq n_{d})\|_{\mathcal{H}}\right]=O_{p}\left(\frac{1}{Nk_{N}^{p}}\right)$
+and it remains to show for the remainder part. By exact Taylor expansion,
+$R_{2}$ $=2\int_{0}^{1}(1-s)\left\{\frac{T_{L,nu}+s\Delta_{nu}}{(T_{L,de}+s\Delta_{de})^{3}}\Delta^{2}_{de}-\frac{1}{(T_{L,de}+s\Delta_{de})^{2}}\Delta_{nu}\Delta_{de}\right\}ds,\text{ with }$
+$\Delta_{nu}\Delta_{de}=I_{b}\left\{\sum_{i=1}^{N}\delta^{2}_{i}I_{i}\mathbf{k}(Y_{i},\cdot)+2\sum_{i\neq j,i<j}\delta_{i}\delta_{j}I_{i}I_{j}\mathbf{k}(Y_{i},\cdot)\right\}-(1-I_{b})T_{L,nu}\sum_{i=1}^{N}\delta_{i}I_{i},$
+$\Delta_{de}^{2}=\left\{\sum_{i=1}^{N}\delta^{2}_{i}I_{i}+2\sum_{i\neq j,i<j}\delta_{i}\delta_{j}I_{i}I_{j}\right\}.$
+Denote $R_{2}=\int(1-s)M_{s}ds$.
+Clearly $M_{s}$ is measurable as all $\Delta_{nu}$, $\Delta_{de}$, $T_{L,nu}$, and $T_{L,de}$ are measurable.
+Also, $\|T_{L,nu}\|_{\mathcal{H}}=\|\sum_{i}\frac{\xi_{i}}{\pi_{i}}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbf{k}(Y_{i},\cdot)\|_{\mathcal{H}}\leq K\frac{N}{\underline{\lambda}n_{s,N}}n_{L(\mathbf{x})}.$ Further, $\mathbb{E}[\|\Delta_{nu}\|^{2}_{\mathcal{H}}|\xi]\leq K\sum_{i}Var(\delta_{i}|\xi)I_{i}\leq 2KN/(\underline{\lambda}n_{s,N})n_{L(\mathbf{x})}$, and similarly $\mathbb{E}[\|\Delta_{de}\|^{2}|\xi]\leq O_{p}(1)n_{L(\mathbf{x})},$ so Chebyshev implies $\|\Delta_{mu}\|=O_{p}(1)\sqrt{n_{L(\mathbf{x})}}$ and smilarly for $|\Delta_{de}|$.
+Additionally, the denominator is away from zero, as
+$|T_{L,\text{de}}+s\Delta_{\text{de}}|\geq\big|n_{L(\mathbf{x})}-|\Delta_{\text{de}}|\big|\geq\big|\tfrac{1}{2}n_{L(\mathbf{x})}-O_{p}(\sqrt{n_{L(\mathbf{x})}})\big|=O_{p}(n_{L(\mathbf{x})}).\text{ So}$
+[…]
+for a.s.
+$\xi$ and $L(\mathbf{x})$, so it follows that for any $N$, $\sup_{0\leq s\leq 1}\mathbb{E}\left[\|M_{s}\|_{\mathcal{H}}\mid\xi,L(\mathbf{x})\right]=O_{p}(Nk_{N}^{p}).$ By Fubini, $\mathbb{E}[R_{2}\mid\xi,L(\mathbf{x})]=\int(1-s)\mathbb{E}[M_{s}\mid\xi,L(\mathbf{x})]ds$, and Minkowski together with Jensen’s inequality gives
+$\Big\|\mathbb{E}[R_{2}\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\leq\int_{0}^{1}(1-s)\Big\|\mathbb{E}[M_{s}\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}ds\leq\int_{0}^{1}(1-s)\mathbb{E}\Big[\|M_{s}\|_{\mathcal{H}}\;\Big|\;\xi,L(\mathbf{x})\Big]ds.$
+Denote vector $(\Delta_{nu},\Delta_{de})=\mathbf{\Delta}$, so $M_{s}=\langle\mathbf{\Delta},H(\tilde{T}_{nu}(s),\tilde{T}_{de}(s))\mathbf{\Delta}\rangle_{\mathcal{H}}$, $\|M_{s}\|_{\mathcal{H}}\leq\|H(\tilde{T}_{s})\|_{op}(\|\Delta_{nu}\|^{2}_{\mathcal{H}}+\Delta_{de}^{2})$.
+Also, $H(\cdot)$ is Lipschitz as on a.e.
+set $\mathcal{E}=\{b\in\mathcal{B}_{L}(\mathbf{x})\}$, $T^{*}_{de}$ is bounded from below by at least 1, and
+$\|H_{g}(t_{nu},t_{de})-H_{g}(\tilde{t}_{nu},\tilde{t}_{de})\|_{op}\leq\|H_{g}(t_{nu},t_{de})-H_{g}(\tilde{t}_{nu},\tilde{t}_{de})\|_{F}$
+$\leq$ $\sqrt{2|1/\tilde{t}^{2}_{de}-1/{t}^{2}_{de}|^{2}+2\|t_{nu}/t^{3}_{de}-\tilde{t}_{nu}/\tilde{t}^{3}_{de}\|^{2}_{\mathcal{H}}}$
+$\leq$ $\sqrt{2\,(2c^{-3})^{2}\,(t_{de}-\tilde{t}_{de})^{2}+\bigl(2c^{-3}\|t_{nu}-\tilde{t}_{nu}\|+6Kc^{-4}\lvert t_{de}-\tilde{t}_{de}\rvert\bigr)^{2}}$
+$\leq$ $L_{lip}\|(t_{nu},t_{de})-(\tilde{t}_{nu},\tilde{t}_{de})\|_{\mathcal{H}},$
+by taking $L_{lip}=\sqrt{16\,c^{-6}\;+\;2\,(6\,K\,c^{-4})^{2}}<\infty$. Take $(\tilde{t}_{nu},\tilde{t}_{de})$ to be $({T}^{*}_{nu},{T}^{*}_{de})$, thus
+$\int_{0}^{1}(1-s)\mathbb{E}\Big[\|M_{s}\|_{\mathcal{H}}\;\Big|\;\xi,L(\mathbf{x})\Big]ds$
+$\leq$ $\int_{0}^{1}(1-s)\mathbb{E}\Big[(\|H(T_{L})\|_{op}+\|H(\tilde{T}_{s})-H(T_{L})\|_{op})\|\mathbf{\Delta}\|^{2}_{\mathcal{H}\times\mathcal{H}}\;\Big|\;\xi,L(\mathbf{x})\Big]ds$
+$\leq$ $L_{lip}\int_{0}^{1}(1-s)\mathbb{E}\Big[\Big(|\Delta_{de}|^{2}+\|\Delta_{nu}\|^{2}_{\mathcal{H}}\Big)^{3/2}\;\Big|\;\xi,L(\mathbf{x})\Big]ds+1/2K_{op}\Big(\mathbb{E}[\Delta^{2}_{de}|L(\mathbf{x})]+\mathbb{E}[\|\Delta_{nu}\|_{\mathcal{H}}^{2}\mid\xi,L(\mathbf{x})]\Big)$
+$\leq$ $\frac{\sqrt{2}}{2}\;L_{lip}\Big(\mathbb{E}\left[|\Delta_{de}|^{3}\;\Big|\;\xi,L(\mathbf{x})\right]+\mathbb{E}\left[\|\Delta_{nu}\|^{3}_{\mathcal{H}}\;\Big|\;\xi,L(\mathbf{x})\right]\Big)$
+$+1/2K_{op}\Big(\mathbb{E}[|\Delta|_{de}^{2}\mid\xi,L(\mathbf{x})]+\mathbb{E}[\|\Delta_{nu}\|_{\mathcal{H}}^{2}\mid\xi,L(\mathbf{x})]\Big).$
+As $\mathbb{E}[|\Delta_{de}|^{2}\mid L(\mathbf{x})]\leq(\mathbb{E}[|\Delta_{de}|^{3}\mid L(\mathbf{x})])^{2/3}$ and similarly for $\Delta_{nu}$, the parts regarding the second moment can be ignored at the cost of increasing the $\sqrt{2}/2L_{lip}$ constant.
+Take the outer expectation w.r.t.
+$L(\mathbf{x})$ and $\xi$, we have that
+$\mathbb{E}_{\xi,L(\mathbf{x})}\left[\Big\|\mathbb{E}[R_{2}\mid\xi,L(\mathbf{x})]\Big\|_{\mathcal{H}}\right]\lesssim\mathbb{E}_{L(\mathbf{x})}\Bigg\{\mathbb{E}\left[\big|\Delta_{de}\big|^{3}\mid L(\mathbf{x})\right]+\mathbb{E}\left[\big|\Delta_{nu}\big|^{3}\mid L(\mathbf{x})\right]\Bigg\}.$
+Inside the expectation w.r.t. $L(\mathbf{x})$, the two terms are both $o_{p}(1)$ as
+$\mathbb{E}_{\xi}\big[\big[\mathbb{E}_{n^{*}}\big[\big|\frac{n^{*}_{i}}{\pi_{i}}\big|^{3}\mid\xi\big]\big]\leq\mathbb{E}_{\xi}\Big[\Big(\mathbb{E}_{n^{*}}\big[\big|\frac{n^{*}_{i}}{\pi_{i}}\big|^{4}\mid\xi\big]\Big)^{3/4}\Big]\leq\mathbb{E}_{\xi}\big[\frac{15^{3/4}\xi_{i}}{\pi^{3}_{i}}\big]\lesssim\mathbb{E}[\big(\frac{N}{n_{s,N}}\big)^{2}]=O_{p}(1),\text{ and }$
+$\mathbb{E}\left[|\frac{\xi_{i}}{\pi_{i}}|^{3}\right]=\frac{1}{\pi^{2}_{i}}\leq\frac{1}{\underline{\lambda}^{2}}\mathbb{E}\big[\frac{N}{n_{s,N}}\big]=O_{p}(1),$
+[…]
+$\mathbb{E}\left[\big|\Delta_{de}\big|^{3}\mid L(\mathbf{x})\right]=\left(\big\|\sum_{i=1}^{N}\delta_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\big\|_{L^{3}(P(n^{*},\xi\mid L(\mathbf{x})))}\mid L(\mathbf{x})\right)^{3}$
+$\leq$ $\left(\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\Big\|\frac{n^{*}_{i}}{\pi_{i}}-\frac{\xi_{i}}{\pi_{i}}\Big\|_{L^{3}(P(n^{*},\xi\mid L(\mathbf{x})))}\mid L(\mathbf{x})\right)^{3}$
+$\leq$  $\;4^{1/3}\left(\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\Big(\mathbb{E}_{\xi}\big[\mathbb{E}_{n^{*}}\big[\big|\frac{n^{*}_{i}}{\pi_{i}}\big|^{3}\mid\xi\big]\big]+\mathbb{E}_{\xi}\big[|\frac{\xi_{i}}{\pi_{i}}|^{3}\big]\Big)^{1/3}\right)^{3}=o_{p}(1).$
+For $\Delta_{nu}=I_{b}\sum_{i}\delta_{i}I_{i}k_{i}+(1-I_{b})T_{L,nu}$, by the inequality $\|a+b\|^{3}\leq(\|a\|+\|b\|)^{3}\leq 4(\|a\|^{3}+\|b\|^{3})$ and boundedness of $I_{b}$, an argument similar to what was used for $\Delta_{de}$ can be used for the $I_{b}\sum_{i}\delta_{i}I_{i}k_{i}$ part.
+Then, regularity condition (A4) and bounded kernel implies boundedness of $T_{L,nu}$ in $\|\cdot\|_{\mathcal{H}}$, and Lemma S10 carries $\mathbb{E}[(1-I_{b})\|T_{L,nu}\|^{3}\mid L(\mathbf{x})]=o_{p}(1)$.
+Part D
+Now we bound the distance introduced by $\xi$ in both the numerator and denominator (leaf size).
+$\begin{split}D_{N}(\mathbf{x})=&\Bigg\|\frac{\sum_{i}\xi_{i}/\pi_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))k_{(}Y_{i},\cdot)}{\sum_{i}\xi_{i}/\pi_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))}-\frac{\sum_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))k_{(}Y_{i},\cdot)}{\sum_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))}\Bigg\|_{\mathcal{H}}\\ =&\Big\|\sum_{i}\tilde{\delta}_{i}\Big\{\frac{\mathbbm{1}(X_{i}\in L(\mathbf{x}))}{\sum_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))}(k_{i}-T_{L}(\mathbf{x}))\Big\}+\widetilde{R}_{2}\Big\|_{\mathcal{H}}\\ \leq&\Big\|\sum_{i}\tilde{\delta}_{i}\Big\{\frac{\mathbbm{1}(X_{i}\in L(\mathbf{x}))}{\sum_{i}\mathbbm{1}(X_{i}\in L(\mathbf{x}))}(k_{i}-T_{L}(\mathbf{x}))\Big\}\Big\|_{\mathcal{H}}+\Big\|\widetilde{R}_{2}\Big\|_{\mathcal{H}},\end{split}$ (S41)
+where the first equality uses again the Taylor expansion in the previous argument with $\tilde{\delta}_{i}=(\frac{\xi_{i}}{\pi_{i}}-1),$ $\widetilde{\Delta}_{nu}=\sum_{i}\tilde{\delta}_{i}I_{i}k_{i},$ $\widetilde{\Delta}_{de}=\sum_{i}\tilde{\delta}_{i}I_{i}.$ We proceed to show both parts $o_{p}(1)$ w.r.t.
+$P_{\mathcal{S}^{N}|\omega}$.
+For the first term, in expectation,
+$\begin{split}&\mathbb{E}_{\xi,L(\mathbf{x})}\!\left[\left\|\sum_{i}\tilde{\delta}_{i}\,\frac{I\!\big(X_{i}\in L(\mathbf{x})\big)}{N_{L(\mathbf{x})}}\,\big(k_{i}-T_{L}(\mathbf{x})\big)\right\|_{\mathcal{H}}\right]\\ \leq&\left(\mathbb{E}_{\xi,L(\mathbf{x})}\!\left[\left\|\sum_{i}\tilde{\delta}_{i}\,\frac{I\!\big(X_{i}\in L(\mathbf{x})\big)}{N_{L(\mathbf{x})}}\,\big(k_{i}-T_{L}(\mathbf{x})\big)\right\|_{\mathcal{H}}^{2}\right]\right)^{1/2}\stackrel{{\scriptstyle denote}}{{=}}T_{N}^{1/2},\end{split}$ (S42)
+where the inner part satisfies for any fixed $L(\mathbf{x})$,
+$T_{N}$ $=\mathrm{Var}_{\xi}\!\left[\left.\sum_{i}\tilde{\delta}_{i}\,\frac{I\!\big(X_{i}\in L(\mathbf{x})\big)}{N_{L(\mathbf{x})}}\,\big(k_{i}-T_{L}(\mathbf{x})\big)\ \right|\ L(\mathbf{x})\right]$
+$=\frac{1}{N_{L(\mathbf{x})}^{2}}\sum_{i}\sum_{j}\mathbb{E}_{\xi}\!\left[\tilde{\delta}_{i}\tilde{\delta}_{j}\right]\,I\!\big(X_{i}\in L(\mathbf{x})\big)\,I\!\big(X_{j}\in L(\mathbf{x})\big)\,\big\langle k_{i}-T_{L},\;k_{j}-T_{L}\big\rangle_{\mathcal{H}}$
+$\leq\frac{4K^{2}}{N_{L(\mathbf{x})}^{2}}\left[\sum_{i}\frac{1}{\pi_{i}^{2}}I_{i}+\sum_{i\neq j}\frac{C_{1}}{n_{s,N}\pi_{i}\pi_{j}}\,I_{i}I_{j}\right]\leq\frac{4K^{2}}{N_{L(\mathbf{x})}^{2}}\left[\frac{N^{2}}{\lambda^{2}n_{s,N}^{2}}\,N_{L(\mathbf{x})}+\frac{C_{1}N^{2}}{\lambda^{2}n_{s,N}^{3}}\,N_{L(\mathbf{x})}^{2}\right]$
+$=O_{p}\!\left(\frac{N^{2}}{n_{s,N}^{2}}\cdot\frac{1}{N_{L(\mathbf{x})}}\right)=O_{p}\!\left(\frac{1}{N_{L(\mathbf{x})}}\right).$
+The first inequality holds by (D4) and boundedness of $\mathbf{k}$ (K1); the second inequality by (D3); the next equality by domination from the first term and $N_{L(\mathbf{x})}=\Theta_{p}(Nk^{p}_{N})$ from Lemma S7; the following equality holds by (D2) implying $n_{s,N}=\Theta_{p_{d}}(N)$.
+Thus plugging it back to the upper bound in (S42) yields
+$\mathbb{E}_{\xi,L(\mathbf{x})}\!\left[\left\|\sum_{i}\tilde{\delta}_{i}\,\frac{I\!\big(X_{i}\in L(\mathbf{x})\big)}{N_{L(\mathbf{x})}}\,\big(k_{i}-T_{L}(\mathbf{x})\big)\right\|_{\mathcal{H}}\right]=\sqrt{\mathbb{E}[\frac{C}{N_{L(\mathbf{x})}}]}$
+$=$ $O\!\left(\frac{1}{\sqrt{N_{L(\mathbf{x})}}}\right)=O\!\left((Nk_{N}^{p})^{-1/2}\right)=o(1),$
+where the second equality by Lemma S7 again.
+For $\mathbb{E}[\|\widetilde{R}_{2}\|]$, an argument similar to Part B can be used:
+$\mathbb{E}\left[\|\widetilde{R}_{2}\|_{\mathcal{H}}\right]\leq\mathbb{E}\left[\int_{0}^{1}(1-s)\|\widetilde{M}_{s}\|_{\mathcal{H}}ds\right]\lesssim\mathbb{E}\left[\mathbb{E}\left[|\widetilde{\Delta}_{de}|^{3}\mid L(\mathbf{x})\right]\right]+\mathbb{E}\left[\mathbb{E}\left[\|\widetilde{\Delta}_{nu}\|_{\mathcal{H}}^{3}\mid L(\mathbf{x})\right]\right],\text{ where }$
+$\mathbb{E}\left[|\widetilde{\Delta}_{de}|^{3}\mid L(\mathbf{x})\right]\leq\left(\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\Big\|\frac{\xi_{i}}{\pi_{i}}-1\Big\|_{L^{3}(P(n^{*},\xi\mid L(\mathbf{x})))}\mid L(\mathbf{x})\right)^{3}=o_{p}(1),\text{ as }$
+$\sup_{i\in[N]}\mathbb{E}\left[(\frac{\xi_{i}}{\pi_{i}}-1)^{3}\mid L(\mathbf{x})\right]=\sup_{i\in[N]}\Big(\pi_{i}(\hat{w}_{i}-1)^{3}+(1-\pi_{i})\Big)\leq\sup_{i\in[N]}\frac{1}{\pi_{i}^{2}}+1=O_{p}(1)=o_{p}(1/N_{L(\mathbf{x})}).$
+Similarly, $\mathbb{E}_{\xi}\left[|\widetilde{\Delta}_{nu}|_{nu}^{3}\mid L(\mathbf{x})\right]$ by the kernel be uniformly bounded, implying that $\mathbb{E}_{\xi,L(\mathbf{x})}\left[\|\widetilde{R}_{2}\|_{\mathcal{H}}\right]\lesssim\mathbb{E}_{L(\mathbf{x})}[o_{p}(1)]=o_{p}(1).$
+[…]
+Part E
+Finally, We bound the approximation error from using the leaf-based conditional distribution instead of the point-based target.
+Consider $E_{N}(\mathbf{x})=\|\;T_{L,b}(\mathbf{x})-\mu_{N}(\mathbf{x})\;\|_{\mathcal{H}}.$ Denote the kernel mean embedding for the true super–population conditional $\mu_{\mathcal{P}}(\mathbf{x}):=\mathbb{E}_{P}[\mathbf{k}(Y,\cdot)\mid X=\mathbf{x}]\in\mathcal{H}.$ We first decompose $T_{L,b}(\mathbf{x})$ using smoothness.
+For any $i$ with $X_{i}\in L(\mathbf{x})$, write
+$\mathbf{k}(Y_{i},\cdot)=\mathbb{E}[\mathbf{k}(Y,\cdot)\mid X_{i}]+\varepsilon_{i},$
+where $\varepsilon_{i}:=\mathbf{k}(Y_{i},\cdot)-\mathbb{E}[\mathbf{k}(Y,\cdot)\mid X_{i}]$ satisfies $\mathbb{E}[\varepsilon_{i}\mid X_{i}]=0$ and $\|\varepsilon_{i}\|_{\mathcal{H}}\leq 2K$ a.s.
+by (K1).
+Then
+$\begin{split}T_{L,b}(\mathbf{x})&=\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\left[\mathbb{E}[\mathbf{k}(Y,\cdot)\mid X_{i}]+\varepsilon_{i}\right]\\ &=\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbb{E}[\mathbf{k}(Y,\cdot)\mid X_{i}]+\underbrace{\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\varepsilon_{i}}_{=:R_{N}(\mathbf{x})}.\end{split}$ (S43)
+Since $\{\varepsilon_{i}\}$ are centered with $\|\varepsilon_{i}\|_{\mathcal{H}}\leq 2K$ and $\mathbb{E}[\|\varepsilon_{i}\|_{\mathcal{H}}^{2}\mid X_{i}]\leq\sigma_{\mathbf{k}}^{2}<\infty$,
+$\mathbb{E}_{L}[\|R_{N}(\mathbf{x})\|_{\mathcal{H}}]\leq\sqrt{\mathbb{E}_{L}[\|R_{N}(\mathbf{x})\|_{\mathcal{H}}^{2}]}=\Big(\frac{1}{N_{L(\mathbf{x})}^{2}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbb{E}[\|\varepsilon_{i}\|_{\mathcal{H}}^{2}]\Big)^{1/2}$
+$\leq$ $(\frac{\sigma_{\mathbf{k}}^{2}}{N_{L(\mathbf{x})}})^{1/2}=O\left((Nk_{N}^{p})^{-1/2}\right)=o(1),$
+where the order are obtained by $N_{L(\mathbf{x})}=\Theta_{p}(Nk_{N}^{p})$ from Lemma S7 and $Nk_{N}^{p}\to\infty$ from (B1). For any $X_{i}\in L(\mathbf{x})$, b the Lipschitz assumption,
+[…]
+$\left\|\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))\mathbb{E}[\mathbf{k}(Y,\cdot)\mid X_{i}]-\mu_{\mathcal{P}}(\mathbf{x})\right\|_{\mathcal{H}}$ $\leq\frac{1}{N_{L(\mathbf{x})}}\sum_{i=1}^{N}\mathbbm{1}(X_{i}\in L(\mathbf{x}))L_{lip}\cdot\text{diam}(L(\mathbf{x}))$
+$=L_{lip}\cdot\text{diam}(L(\mathbf{x})).$
+Taking expectations over $L(\mathbf{x})$ and using (B1):
+$\mathbb{E}_{L}[\text{diam}(L(\mathbf{x}))]$ $=\mathbb{E}_{L}[\text{diam}(L(\mathbf{x}))\cdot\mathbbm{1}(\text{diam}(L(\mathbf{x}))<k_{N})]$
+$\quad+\mathbb{E}_{L}[\text{diam}(L(\mathbf{x}))\cdot\mathbbm{1}(\text{diam}(L(\mathbf{x}))\geq k_{N})]$
+$\leq k_{N}+\text{diam}(\mathcal{X})\cdot P_{L}(\text{diam}(L(\mathbf{x}))\geq k_{N})$
+$\leq k_{N}+\text{diam}(\mathcal{X})\cdot t_{N}=k_{N}+o(1),$
+where $\text{diam}(\mathcal{X})<\infty$ since $\mathcal{X}$ is bounded, and $t_{N}\to 0$ by (B1). Combining previous results into (S43), we have
+$\mathbb{E}_{L}[\|T_{L,b}(\mathbf{x})-\mu_{\mathcal{P}}(\mathbf{x})\|_{\mathcal{H}}]\leq L_{lip}(k_{N}+o(1))+o(1)=O(k_{N}).$
+For the finite-population target, on $N_{\mathbf{x}}:=\sum_{i}\mathbbm{1}(X_{i}=\mathbf{x})\geq 1$ with high probability, $\mu_{N}(\mathbf{x})=\frac{1}{N_{\mathbf{x}}}\sum_{i:X_{i}=\mathbf{x}}\mathbf{k}(Y_{i},\cdot)$ is the sample average of $N_{\mathbf{x}}$ i.i.d.
+draws from $P(Y\mid X=\mathbf{x})$.
+By the law of large numbers, $\mu_{N}(\mathbf{x})\to\mu_{\mathcal{P}}(\mathbf{x})$ almost surely as $N_{\mathbf{x}}\to\infty$.
+By the triangle inequality,
+$E_{N}(\mathbf{x})\leq\|T_{L,b}(\mathbf{x})-\mu_{\mathcal{P}}(\mathbf{x})\|_{\mathcal{H}}+\|\mu_{\mathcal{P}}(\mathbf{x})-\mu_{N}(\mathbf{x})\|_{\mathcal{H}}=O_{p_{d}}(k_{N})+o_{p}(1).$
+By assumption (B1), $k_{N}\to 0$ as $N\to\infty$. Therefore, $E_{N}(\mathbf{x})=o_{p_{d}}(1)$ with respect to $P_{\mathcal{S}^{N}\mid\omega}$ and algorithmic randomness. This completes Part E. ∎
+S8 Consistency of Functional Estimate
+Beginning this section, we focus on the goal of consistency for the functional of interest $\Psi$. We first show a sufficient condition for this to be true.
+Lemma S11 (Theorem 3, [1958_Varadarajan]).
+Suppose $\mathcal{Y}$ locally compact Hausdorff and second-countable (hence metrizable and separable).
+Let $Y_{1},Y_{2},\dots,$ i.i.d.
+Borel-measurable random mappings of $\Omega$ into $\mathcal{Y}$ and let ${P}_{N}$ be the induced (empirical) measure.
+Then
+$\Pr\!\bigl({P}_{N}\rightsquigarrow P\bigr)\;=\;1.$
+Proof of Corollary 3.5.
+Denote ${P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the design space and $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ is an estimate for it.
+Denote $P_{\Omega}(\mathbf{x})$ as the true conditional measure of $Y|X=\mathbf{x}$ in the super population space.
+Part A, Design Consistency
+For any fixed $\omega\in\Omega$, the hypothesis $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),P_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\xrightarrow{p_{d}}0$ as $N\to\infty$ implies, by the subsequence criterion, that for any subsequence $\{N_{k}\}$, there exists a further subsequence $\{N_{k_{\ell}}\}$ such that
+$d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}),P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}))\to 0\quad P_{\mathcal{S}^{N}|\omega}\text{-a.s.}$
+By metrization from (K2), this implies $\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\rightsquigarrow P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})$ a.s.
+By the Portmanteau theorem, this yields:
+(I) $\int f\,d\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\to\int f\,dP_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\quad\text{for all }f\in C_{b}(\mathcal{Y}),$
+(II) $\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}((-\infty,t];\mathbf{x})\to P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}((-\infty,t];\mathbf{x})\quad\text{for all continuity points }t\in\mathbb{R}^{d}\text{ of }P_{\mathcal{S}^{N}|\omega}(\mathbf{x}).$
+By the sub-sequence being arbitrary, using the other side of the if and only if statement of the subsequence criterion, we have that (I) and (II) apply also to the overall sequence:
+(I) $\int f\,d\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})\xrightarrow{p_{d}}\int f\,dP_{\mathcal{S}^{N}|\omega}(\mathbf{x})\quad\text{for all }f\in C_{b}(\mathcal{Y}),$
+(II) $\widehat{P}_{\mathcal{S}^{N}|\omega}((-\infty,t];\mathbf{x})\xrightarrow{p_{d}}P_{\mathcal{S}^{N}|\omega}((-\infty,t];\mathbf{x})\quad\text{for all continuity points }t\text{ of }P_{\mathcal{S}^{N}|\omega}(\mathbf{x}).$
+Since $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),d_{\mathbf{k}})\to(\mathcal{V},d)$ is continuous at $P_{\mathcal{S}^{N}|\omega}(\mathbf{x})$, the continuous mapping theorem yields
+$\Psi(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\xrightarrow{p_{d}}\Psi(P_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\quad\text{as }N\to\infty.$
+This establishes design consistency.
+Part B, Model Consistency
+Since $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{N}$ are i.i.d.
+from $P_{\Omega}$, the joint super population measure, for each fixed $\mathbf{x}\in\mathcal{X}$ and $\omega\in\Omega$, the conditional observations $\{Y_{i}:X_{i}(\omega)=\mathbf{x}\}$ form an i.i.d.
+sample from $P_{\Omega}(\mathbf{x})=P_{\Omega}(\cdot\mid X=\mathbf{x})$.
+Define $N_{\mathbf{x}}(\omega)=\sum_{i=1}^{N}\mathbb{I}(X_{i}(\omega)=\mathbf{x})$ and the empirical measure $P_{\mathcal{S}^{N}|\omega}(\cdot\;;\;\mathbf{x})=\frac{1}{N_{\mathbf{x}}(\omega)}\sum_{i:X_{i}(\omega)=\mathbf{x}}\delta_{Y_{i}(\omega)}(\cdot).$ By Lemma S11,
+$P_{\mathcal{S}^{N}|\omega}(\cdot\;;\;\mathbf{x})\rightsquigarrow P_{\Omega}(\cdot\;;\;\mathbf{x})\quad\text{for }P_{\Omega}\text{-a.e.\ }\omega\in E\subset\Omega\quad\text{as }N\to\infty,$ (S44)
+with $P_{\Omega}(E)=1$. As $d_{\mathbf{k}}$ metrizes weak convergence by assumption (K2), (S44) is equivalent to
+$d_{\mathbf{k}}(P_{\mathcal{S}^{N}|\omega}(\mathbf{x}),P_{\Omega}(\mathbf{x}))\to 0\quad\text{for }P_{\Omega}\text{-a.e.\ }\omega\in E.$ (S45)
+For any fixed $\omega\in\Omega$, $d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}),P_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\xrightarrow{p_{d}}0$ as $N\to\infty$ by hypothesis implies, by the subsequence criterion, that for any subsequence $\{N_{k}\}$, there exists a further subsequence $\{N_{k_{\ell}}\}$ such that
+$d_{\mathbf{k}}\left(\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}),P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\right)\to 0\quad P_{\mathcal{S}^{N}|\omega}\text{-a.s.}$
+on $s\in E_{\omega}\subset\mathcal{S}^{N}$, with $P_{\mathcal{S}^{N}|\omega}(E_{\omega})=1$.
+By metrization from (K2), this implies $\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\rightsquigarrow{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})$ a.s.
+on $E_{\omega}$.
+Recall the joint probability space $(\Omega\times\mathcal{S}^{N},\mathcal{A}\otimes\sigma(\mathcal{S}^{N}),P_{\Omega}\otimes P_{\mathcal{S}^{N}|\omega})$.
+Define $E^{*}=\{(\omega,s):\omega\in E,s\in E_{\omega}\}$.
+By Fubini, $\int_{\Omega}P_{\mathcal{S}^{N}|\omega}(E_{\omega})P_{\Omega}(d\omega)=1$, implying that w.r.t.
+the joint measure, $(P_{\Omega}\otimes P_{\mathcal{S}^{N}|\omega})(E^{*})=1$.
+Thus, for any fixed path indexed by $(\omega,s)\in E^{*}$,
+$\begin{split}d_{\mathbf{k}}(\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}),\;P_{\Omega}(\mathbf{x}))&\leq d_{{\mathbf{k}}}\left(\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}),\;P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\right)\\ &\quad+d_{{\mathbf{k}}}\left(P_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x}),\,P_{\Omega}(\mathbf{x})\right)\to 0\end{split}$ (S46)
+as $\ell\to\infty$, since both terms on the right-hand side vanish: the first by the design convergence on $E_{\omega}$ and the second by (S45) on $E$.
+By metrization of $d_{{\mathbf{k}}}$, (S46) implies $\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\rightsquigarrow P_{\Omega}(\mathbf{x})$ on $E^{*}$.
+Thus by the Portmanteau theorem on the joint space, we have
+(I) $\int f\,d\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}(\mathbf{x})\to\int f\,dP_{\Omega}(\mathbf{x})\quad\text{for all }f\in C_{b}(\mathcal{Y})$
+(II) $\widehat{P}_{\mathcal{S}^{N_{k_{\ell}}}|\omega}((-\infty,t];\mathbf{x})\to P_{\Omega}((-\infty,{t}];\mathbf{x})\quad\text{for all continuity points }{t}\in\mathbb{R}^{d}\text{ of }P_{\Omega}(\mathbf{x}).$
+By the sub-sequence being arbitrary, using the other side of the if and only if statement of the subsequence criterion, we have that (I) and (II) apply also to the overall sequence:
+(I) $\int f\,d\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})\xrightarrow{p_{\otimes}}\int f\,dP_{\Omega}(\mathbf{x})\quad\text{for all }f\in C_{b}(\mathcal{Y})$
+(II) $\widehat{P}_{\mathcal{S}^{N}|\omega}((-\infty,t];\mathbf{x})\xrightarrow{p_{\otimes}}P_{\Omega}((-\infty,t];\mathbf{x})\quad\text{for all continuity points }{t}\in\mathbb{R}^{d}\text{ of }P_{\Omega}(\mathbf{x}).$
+Finally, since $\Psi:(\mathcal{M}_{b}(\mathcal{Y}),d_{\mathbf{k}})\to(\mathcal{V},d)$ is continuous at $P_{\Omega}(\mathbf{x})$ for $P_{\Omega}$-a.e.
+$\omega$, the continuous mapping theorem applied to (I) yields
+$\Psi(\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x}))\xrightarrow{p_{\otimes}}\Psi(P_{\Omega}(\mathbf{x})),$
+establishing model consistency and completes the proof. ∎
+Remark.
+Corollary 3.5 suggests that two types of consistency of the functional estimand, once i.i.d.
+of the finite population has been assumed, requires no more than convergence in $p_{d}$ of the MMD between $\widehat{P}_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ and $P_{\mathcal{S}^{N}|\omega}(\mathbf{x})$ in the embedded RKHS.
+S9 SDRF (a variation)
+Algorithm 2 provides a variation that helps spend PSU more gradually as depth increases, as mentioned in Section 3.1.
+It draws fresh Bernoulli indicators at each depth, incrementally allocating PSUs to $D^{*}_{\mathrm{split}}$ as the tree grows (Algorithm 1).
+This depth-adaptive strategy addresses two concerns.
+First, it “spends” PSUs more conservatively: trees that stop early use fewer PSUs for splitting, preserving a larger estimation sample, while deeper trees gradually recruit additional PSUs as needed.
+Second, this approach more closely aligns with the endogenous post-stratification (EPS) framework studied by Breidt and Opsomer (2008) and Dahlke et al.
+(2011), who established asymptotic equivalence between stratification based on sample-fitted models and traditional post-stratification with known strata.
+However, RF involves recursive splitting, and formal extension of these results to our setting remains an open question.
+The cost is computational: tracking time-varying inclusion probabilities $q_{d,i}=\prod_{j=1}^{d}(\tilde{q}I_{k(i),j}+(1-\tilde{q})(1-I_{k(i),j}))$ and maintaining depth-indexed data partitions.
+Algorithm 2 Survey Distributional Random Forest (SDRF)
+[…]
+0: Random forest ensemble $\mathcal{F}=\{T_{1},\dots,T_{B}\}$
+[…]
+2:for $b=1$ to $B$ do
+3:  Draw bootstrap sample $D^{*}_{b}=\{(X_{i},\;Y,\;1/\pi_{i}\times n^{*}_{b,i})\}_{i=1}^{n_{s}}$ mimicking the design.
+4:  Train decision tree $T_{b}$ on $D^{*}_{b}$ by:
+5:  for $d=1,\dots$ do
+6:   Draw $\forall k(i)$, $I_{k(i)}:=\mathbbm{1}(k(i)\in D^{*}_{b,split})\sim Bernoulli(q)$ for all PSUs.
+7:   Update $D^{*}_{\mathrm{split},d}\leftarrow D^{*}_{\mathrm{split},d-1}\bigcup\{D^{*}_{i}:I_{k(i)}=1\}$. $D^{*}_{\mathrm{est},d}\leftarrow D_{b}^{*}\setminus D^{*}_{\mathrm{split},d}$.
+8:   Update ${q}_{d,i}={q}_{d-1,i}\times(\tilde{q}\;I_{k(i),d}+(1-\tilde{q})(1-I_{k(i),d})\}$
+9:   Randomly select $mtry\leq p$ features of $X$.
+10:   Choose the split that maximizes $M^{*}(\mathrm{split};D^{*}_{\mathrm{split},d})=\frac{N^{*}_{L}N^{*}_{R}}{N^{*2}_{P_{a}}}\|\mu_{\mathbf{k}}(\hat{P}^{*}_{L})-\mu_{\mathbf{k}}(\hat{P}^{*}_{R})\|^{2}_{\mathcal{H}}$ until reaching maximum depth, minimum node size, or maximum weight ratio.
+11:   $\mathcal{T}\leftarrow\mathcal{T}\;\bigcup\;\{T_{b}\}$.
+12:  end for
+13:end for
+14:return For all $b$ with $i\in\mathcal{I}(D^{*}_{b,est})$,
+$\widehat{\omega}_{i}(\cdot)=\frac{1}{B}\sum_{b=1}^{B}\frac{n^{*}_{bi}/(\pi_{i}(1-q_{d,i}))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}{\sum_{i\in\mathcal{I}(D^{*}_{b,est})}n^{*}_{bi}/(\pi_{i}(1-q_{d,i}))\,\mathbbm{1}(X_{i}\in L^{*}_{b}(\mathbf{x}))}$
+15: For $\Psi(\cdot)$, obtain plug-in estimator $\Psi(\widehat{P}_{Y\mid X=\mathbf{x}})=\Psi(\sum_{i=1}^{n_{s}}\widehat{\omega}_{i}(\mathbf{x})\;\mathbf{k}(Y_{i},\,\cdot))$.
+S10 Kernel Assumption Details
+Remark.
+(S1) ensures a standard weak-convergence framework on $\mathcal{Y}$ and established results for $C_{0}$-kernels to go through: since $\mathcal{Y}$ is locally compact Polish, it is in particular locally compact Hausdorff (LCH) and separable and complete.
+On such spaces every Borel probability measure is Radon.
+(K1) ensures KME well–defined and measurable.
+Uniformly bounded means $\sup_{y,y^{\prime}}|\mathbf{k}(y,y^{\prime})|=K<\infty$, and together with $P\in\mathcal{M}_{b}(\mathcal{Y})$ ensures $\int\mathbf{k}(y,\,\cdot)dP(y)<\infty$ a well-defined element in $\mathcal{H}$.
+Positive definiteness (PD) means for any $n\in\mathbb{N}$, $\{y_{1},\dots,y_{n}\}\subset\mathcal{Y}$, $\{\alpha_{1},\dots,\alpha_{n}\}\subset\mathbb{R}$, $\sum_{i,j=1}^{n}\alpha_{i}\alpha_{j}\mathbf{k}(y_{i},y_{j})\geq 0$.
+By the Moore–Aronszajn theorem, PD of $\mathbf{k}$ is equivalent to the existence of a unique RKHS $\mathcal{H}$ of real-valued functions on $\mathcal{Y}$ for which $\mathbf{k}$ is the reproducing kernel ([Aronszajn1950]).
+$C_{0}$-continuous implies that $\mathbf{k}$ is continuous and hence Borel measurable on $\mathcal{Y}\times\mathcal{Y}$ with respect to the product sigma algebra $\mathcal{B}(\mathcal{Y})\otimes\mathcal{B}(\mathcal{Y})$, so the KME is also measurable.
+When (S1) and (K1) hold, $\mathbf{k}$ satisfies (K2) if and only if $\mathbf{k}$ is characteristic to the set of probability measures contained in $\mathcal{M}_{b}(\mathcal{Y})$ (Theorem 6, [Sriperumbudur2010-xr]).
+Also, $\mathbf{k}$ is $c_{0}$-universal if and only if $\int\int_{\mathcal{Y}}\mathbf{k}(y_{1},\;y_{2})dP(y_{1})dP(y_{2})>0$, $\forall 0\neq P\in\mathcal{M}_{b}(\mathcal{Y})$ (Proposition 8, [Sriperumbudur2010-xr]), providing a practical condition to check for whether (K2) holds.
+(K3) is not required for the statistical properties of the algorithm, but allows for easier computation.
+Translation-invariant means $\mathbf{k}(y_{1},\;y_{2})=\psi(y_{1}-y_{2})$ for $\psi$ a PD function.
+This condition, combined with (K1), allows the Fourier representation $\mathbf{k}(y_{1},y_{2})=\int e^{\mathbf{i}\mkern 1.0mu\mathbf{\omega}^{T}(y_{1}-y_{2})}d\nu(\mathbf{\omega})=\langle\varphi_{y_{1}},\varphi_{y_{2}}\rangle_{\mathcal{H}}$, $\varphi_{y_{i}}=e^{\mathbf{i}\mkern 1.0mu\mathbf{\omega}^{T}y_{i}}\in\mathbb{C}$ by the Bochner’s theorem ([rudin2011-fag-ebook]).
+We append further details the examples of $\mathcal{Y}$ and $\mathbf{k}$ mentioned in Section 3.1:
+  • Euclidean vectors.
+For $\mathcal{Y}=\mathbb{R}^{d}$, $d\geq 1$, a standard choice is the Gaussian (RBF) kernel $\mathbf{k}_{\mathrm{RBF}}(y_{1},y_{2})=\exp\!\bigl(-\|y_{1}-y_{2}\|_{2}^{2}/(2\sigma^{2})\bigr)$, $\sigma>0$, or the Laplacian kernel $\mathbf{k}_{\mathrm{Lap}}(y_{1},y_{2})=\exp\!\bigl(-\|y_{1}-y_{2}\|_{2}/\ell\bigr)$, $\ell>0$.
+Both are bounded, continuous and positive definite, satisfying (K1)-(K3).
+This covers, for example, $Y$ equal to a vector of continuous biomarkers, low-dimensional embeddings of wearable-device trajectories, or latent factor scores from a longitudinal mixed model.
+  • Edge-weighted Graphs with a fixed labeled vertex set.
+Fix $m$ labeled vertices; then $m^{\prime}=m(m-1)/2$ is the number of possible undirected edges.
+Encode each graph by its edge-weight vector $y=(w_{ij})_{1\leq i<j\leq m}\in[0,1]^{m^{\prime}}$.
+Consider $\mathcal{Y}=[0,1]^{m^{\prime}}$ with the subspace topology inherited from $\mathbb{R}^{m^{\prime}}$.
+A special case is $\mathcal{Y}=\{0,1\}^{m^{\prime}}$, corresponding to simple (unweighted) graphs.
+Any RBF-type kernel on $\mathbb{R}^{m^{\prime}}$, such as $\mathbf{k}(y_{1},y_{2})=\exp\!\bigl(-\|y_{1}-y_{2}\|_{2}^{2}/(2\sigma^{2})\bigr)$, is bounded, continuous, positive definite, and $c_{0}$-universal on the compact set $0,1]^{m^{\prime}}$, so (K1)–-(K2) hold.
+This setting includes examples where vertices are study participants and $w_{ij}$ measures the strength of contact or similarity between individuals.
+  • Laplacian matrix embeddings with a fixed labeled vertex set.
+Alternatively, represent a graph on $m$ labeled vertices by its combinatorial Laplacian $L\in\mathbb{R}^{m\times m}$.
+Define $\mathcal{Y}=\{L\in\mathbb{R}^{m\times m}_{\mathrm{sym}}:L\mathbbm{1}=0,\;L_{ij}\leq 0\ (i\neq j),\;L_{ii}=-\sum_{j\neq i}L_{ij}\}$.
+This is a closed subset of the finite-dimensional Euclidean space $\mathbb{R}^{m\times m}_{\mathrm{sym}}$ and hence LCH and Polish; it is compact if the edge weights are uniformly bounded.
+Kernels can be built from the Frobenius norm, for example $\mathbf{k}(L_{1},L_{2})=\exp\!\bigl(-\|L_{1}-L_{2}\|_{F}^{2}/(2\sigma^{2})\bigr)$, $\sigma>0$.
+This representation is natural for structural brain connectivity networks or graphs of spatial neighborhoods in disease mapping, when the analysis is formulated in terms of smoothing or diffusion over the graph.
+  • Periodic, torus-valued data.
+Consider the $d$-dimensional torus $\mathcal{Y}=\mathbb{T}^{d}=(\mathbb{R}/2\pi\mathbb{Z})^{d}$.
+A convenient construction is to embed $\varphi:\mathbb{T}^{d}\to\mathbb{R}^{2d}$, $\varphi(\theta)=(\cos\theta_{1},\sin\theta_{1},\dots,\cos\theta_{d},\sin\theta_{d})^{\top}$, and define $\mathbf{k}(\theta,\theta^{\prime})=\exp\!\bigl(-\|\varphi(\theta)-\varphi(\theta^{\prime})\|_{2}^{2}/(2\sigma^{2})\bigr)$, $\sigma>0$.
+This yields a bounded, continuous, positive definite kernel on $\mathbb{T}^{d}$ that is universal on this compact manifold.
+Such a setup arises when $Y$ is a vector of phases or circadian times.
+For example, the phase difference between heart and respiratory cycles, or circadian timing of hormone peaks.
+  • Histograms/probability simplices.
+Consider the (p1)(p-1)–dimensional probability simplex $\mathcal{Y}=\Delta^{p-1}:=\{w=(w_{1},\ldots,w_{p})^{\top}\in[0,1]^{p}:\sum_{j=1}^{p}w_{j}=1\}$ with $p\geq 2$.
+$\Delta^{p-1}$ is compact.
+Any continuous universal kernel on $\mathbb{R}^{p}$, such as the Gaussian RBF kernel $\mathbf{k}(w,w^{\prime})=\exp\!\bigl(-\|w-w^{\prime}\|_{2}^{2}/(2\sigma^{2})\bigr)$, remains $c_{0}$-universal when restricted to $\Delta^{p-1}$, so (K1)-–(K2) hold.
+Typical elements includes the normalized histograms over $p$ bins or empirical distributions of a categorical variable.